### v1

# Asymptotic Inference in a Stationary Quantum Time Series

Abstract
We consider a statistical model of a n-mode quantum Gaussian state which is shift invariant and also gauge invariant. Such models can be considered analogs of classical Gaussian stationary time series, parametrized by their spectral density. Defining an appropriate quantum spectral density as the parameter, we establish that the quantum Gaussian time series model is asymptotically equivalent to a classical nonlinear regression model given as a collection of independent geometric random variables. The asymptotic equivalence is established in the sense of the quantum Le Cam distance between statistical models (experiments). The geometric regression model has a further classical approximation as a certain Gaussian white noise model with a transformed quantum spectral density as signal. In this sense, the result is a quantum analog of the asymptotic equivalence of classical spectral density estimation and Gaussian white noise, which is known for Gaussian stationary time series. In a forthcoming version of this preprint, we will also identify a quantum analog of the periodogram and provide optimal parametric and nonparametric estimates of the quantum spectral density.
1 Main Results
1.1 Introduction
Quantum stationary time series models have arisen in the context of quantum system identification and control theory [GY16], [LGN18]. For some context, we will first describe some basic asymptotic inference results for classical time series models in statistics.
Local asymptotic normality (LAN, Le Cam [LC86]) is a fundamental property of a sequence of statistical experiments, which essentially reduces inference for large sample size to the case of a normal location model. Let $\left(P_{n,\theta},\theta\in\Theta\right)$ be a sequence of families of p.m.’s on $\left(\Omega_{n},\mathcal{A}_{n}\right)$ where $\Theta\subset\mathbb{R}^{k}$; assume that for given $n$, all $P_{n,\theta}$ are mutually absolutely continuous. The sequence is LAN at $\theta\in\mathrm{int}\left(\Theta\right)$ if there exists a positive $k\times k$ matrix $J_{\theta}$ and random $k$-vectors $\Delta_{n,\theta}$ on $\Omega_{n}$ such that $\mathcal{L}\left(\Delta_{n}|P_{n,\theta}\right)\Rightarrow N\left(0,J_{\theta}\right),$ and for $h\in\mathbb{R}^{k}$ one has
$\log\frac{dP_{n,\theta+h/\sqrt{n}}}{dP_{n,\theta}}=h^{\prime}\Delta_{n,\theta}-\frac{1}{2}h^{\prime}J_{\theta}h+o_{P}\left(1\right)\text{ as }n\rightarrow\infty,$ (1.1)
with probability convergence taking place under the $P_{n,\theta}$ law, uniformly over compacts in $h$. The underlying idea here is that the log-likelihood ratio asymptotically, and locally in neighborhoods of $\theta$, takes the form associated to a Gaussian shift experiment
$\left(N_{k}\left(h,J_{\theta}^{-1}\right),\;h\in\mathbb{R}^{k}\right).$ (1.2)
The latter model then serves as a benchmark for optimal inference in the original model $\left(P_{n,\theta},\theta\in\Theta\right)$, typically giving risk bounds in terms of the Fisher information matrix $J_{\theta}$. One of the earliest results establishing the LAN property, beyond the basic i.i.d. case, has been Davies [Dav73] for a stationary Gaussian time series with spectral density depending on a parameter $\theta$. Later developments and extensions within the framework of parametric statistical inference for time series are summarized in the monographs [Dzh86] and [TK00].
When parameters are infinite dimensional, defining a framework of nonparametric inference, the proper analog of LAN to describe risk benchmarks for procedures is asymptotic equivalence in the sense of Le Cam’s $\Delta$-distance. To define it, assume all measurable sample spaces are Polish (complete separable) metric spaces equipped with their Borel sigma algebra. For measures $P,$ $Q$ on the same sample space, let $\left\|P-Q\right\|_{1}$ be $L_{1}$-distance. For the general case where $P,$ $Q$ are not necessarily on the same sample space, suppose $K$ is a Markov kernel such that $KP$ is a measure on the same sample space as $Q$. In that case, $\left\|Q-KP\right\|_{1}$ is defined and will be used to measure the distance between $Q$ and a Markov kernel randomization of $P$.
Consider now experiments (families of measures) $\mathcal{F}=\left(Q_{\theta},\;\theta\in\Theta\right)$ and $\mathcal{E}=\left(P_{\theta},\;\theta\in\Theta\right)$, on possibly different sample spaces, but with the same parameter space $\Theta$ (of arbitrary nature). All experiments here are assumed dominated by a sigma-finite measure on their respective sample space. The deficiency of $\mathcal{E}$ with respect to $\mathcal{F}$ is defined as
$\delta\left(\mathcal{E},\mathcal{F}\right)=\inf_{K}\sup_{\theta\in\Theta}\left\|Q_{\theta}-KP_{\theta}\right\|_{1}$
where $\inf$ extends over all appropriate Markov kernels. Le Cam’s pseudodistance $\Delta\left(\mathcal{\cdot},\cdot\right)$ between $\mathcal{E}$ and $\mathcal{F}$ then is
$\Delta\left(\mathcal{E},\mathcal{F}\right)=\max\left(\delta\left(\mathcal{E},\mathcal{F}\right),\delta\left(\mathcal{F},\mathcal{E}\right)\right).$ (1.3)
It is well known that for two experiments $\mathcal{E}$ and $\mathcal{F}$ having the same parameter space, $\Delta(\mathcal{E},\mathcal{F})<\varepsilon$ implies that for any decision problem with loss bounded by $1$ and any statistical procedure in the experiment $\mathcal{F}$ there is a (randomized) procedure in $\mathcal{E}$, the risk of which evaluated in $\mathcal{E}$ nearly matches (within $\varepsilon$) the risk of the original procedure evaluated in $\mathcal{F}$. In this statement the roles of $\mathcal{E}$ and $\mathcal{F}$ can also be reversed. Two sequences $\mathcal{E}_{n},\mathcal{F}_{n}$ are said to be asymptotically equivalent if $\Delta(\mathcal{E}_{n},\mathcal{F}_{n})\rightarrow 0$.
A result on approximation in $\Delta$-distance of a classical Gaussian stationary time series model has been obtained in [GNZ10]. Assume a sample $y^{(n)}=\left(y(1),\ldots,y(n)\right)^{\prime}$ from a real Gaussian stationary sequence $y(t)$ with zero mean, autocovariance function $\gamma_{j}=\mathrm{E}y(t)y(t+j)$ and real spectral density $f$ on $\pi,\pi]$ such that $f\left(\omega\right)=f\left(-\omega\right)$ and
$\gamma_{j}=\int_{-\pi}^{\pi}\exp\left(-ij\omega\right)f\left(\omega\right)d\omega.$ (1.4)
Define a nonparametric set $\Sigma_{\alpha,M}=B^{\alpha}(M)\cap\mathcal{F}_{M},$ where $B^{\alpha}(M)$ is a (Besov) smoothness class of spectral densities with smoothness coefficient $\alpha$ and $\mathcal{F}_{M}$ is the set of real even positive functions $f$ on $\pi,\pi]$ such that $\left|\log f\right|\leq M$. Then it is shown that observations $y^{(n)}$ with spectral density $f$ are asymptotically equivalent to a white noise model
$dZ_{\omega}=\log f(\omega)d\omega+2\pi^{1/2}n^{-1/2}dW_{\omega}\text{, }\omega\in[-\pi,\pi]$ (1.5)
if the parameter space is given by $f\in$ $\Sigma_{\alpha,M}$ for some $M>0$ and $\alpha>1/2$. This represents the nonparametric (asymptotic equivalence) version of the classical LAN property for parametric models $\left(f_{\theta},\theta\in\Theta\right)$ of spectral densities [Dav73] [Dzh86] [TK00]. Here the Gaussian white noise model (1.5) represents an analog of the basic Gaussian location model (1.2), with the approximation valid globally (over all spectral densities $f\in$ $\Sigma_{\alpha,M}$). Also established were local approximations (via the connection to [GN98]) around a fixed spectral density $f_{0}$ like
$dZ_{\omega}=f(\omega)d\omega+2\pi^{1/2}n^{-1/2}f_{0}(\omega)dW_{\omega}\text{, }\omega\in[-\pi,\pi]$ (1.6)
which are more suitable for obtaining risk bounds for estimation on $f$ itself, rather than for $\log f$. Here the log-transformation plays the role of a variance stabilizing transformation, removing the factor $f_{0}$ from the noise term and allowing to proceed from the local asymptotic equivalence (1.6) (valid for $f\approx f_{0}$) to the global variant (1.5) (cf. [GN98] for details).
The analog of the $\Delta$-distance for quantum statistical models has been introduced and studied by several authors. In [GK06], [KG09] it was used to define a (strong) quantum analog of the LAN property (1.1) for tensor product models of qubits and finite dimensional states. Alternative approaches to quantum LAN were pursued by [GJ07] and [YFG13], via two different definitions of a quantum likelihood ratio. Recently in [BGN18] the quantum Le Cam distance was used to establish asymptotic equivalence of a tensor product model of infinite dimensional pure states to a quantum Gaussian white noise model. Although the approximation is local, valid in a neighborhood of a fixed pure state (and thus is an analog of (1.6)), it allows to establish some explicit results for nonparametric inference on pure states (estimation and testing).
The object of the present paper is to investigate, with regard to asymptotic equivalence, a quantum Gaussian model studied earlier by Mosonyi [Mos09]. We will consider an $n$-mode quantum Gaussian system to define a quantum Gaussian time series of ”length” $n$.
A one mode quantum system is given by the Hilbert space $L_{2}\left(\mathbb{R}\right)$ and self-adjoint operators acting on appropriately defined domains as
$\left(Qf\right)\left(x\right)=xf\left(x\right),\;\;\;\;\;\left(Pf\right)\left(x\right)=-i\frac{df\left(x\right)}{dx}$
which satisfy the commutation relations
$\left[Q,P\right]=QP-PQ=i\mathbf{1}.$
The Hilbert space of an $n$-mode system is $L_{2}^{\otimes n}\left(\mathbb{R}\right)\cong L_{2}\left(\mathbb{R}^{n}\right)$ on which ”canonical pairs” $\left(Q_{j},P_{j}\right)$ are defined acting on the $j$th tensor factor as above, and as identity on the other tensor factors. Thus the commutation relations on $L_{2}\left(\mathbb{R}^{n}\right)$ are
$\left[Q_{j},Q_{k}\right]=\left[P_{j},P_{k}\right]=0,\left[Q_{j},P_{k}\right]=i\delta_{jk}\mathbf{1.}$ (1.7)
Write the vector of observables as $\mathbf{R}:=\left(Q_{1},\ldots,Q_{n},P_{1},\ldots,P_{n}\right)$ and for $x\in\mathbb{R}^{2n}$ introduce the Weyl unitaries as
$W\left(x\right)=\exp\left(i\mathbf{R}x\right).$ (1.8)
For $x,y\in$ $\mathbb{R}^{2n}$ define a bilinear, antisymmetric (symplectic) form as
$D\left(x,y\right)=\sum_{j=1}^{n}\left(x_{j}y_{j+n}-x_{j+n}y_{j}\right).$
The operators $W\left(x\right)$, $x\in\mathbb{R}^{2n}$ satisfy $W\left(x\right)^{\ast}=W\left(-x\right)$ and
$W\left(x\right)W\left(y\right)=W\left(x+y\right)\exp\left(-\frac{i}{2}D\left(x,y\right)\right),\;x,y\in\mathbb{R}^{2n},$ (1.9)
i.e. the Weyl canonical commutations relations, or CCR. The $C^{\ast}$-algebra generated by $\left\{W\left(x\right),x\in\mathbb{R}^{2n}\right\}$ is the Schrödinger representation of $CCR\left(\mathbb{R}^{2n},D\right)$ ([BR97], 5.2.16). The von Neumann algebra generated by $\left\{W\left(x\right),x\in\mathbb{R}^{2n}\right\}$ is the full algebra $\mathcal{L}(L_{2}\left(\mathbb{R}^{n}\right))$ of bounded operators on $L_{2}\left(\mathbb{R}^{n}\right)$.
1.2 Gaussian states
A state $\varphi$ on a von Neumann algebra $\mathcal{A}$ is a positive normal linear functional $\varphi:$ $\mathcal{A}\rightarrow\mathbb{C}$ which takes value $1$ on the unit of $\mathcal{A}$; cf. Section A.1 for a short overview of the relevant concepts. In the case of $\mathcal{A=L}(L_{2}\left(\mathbb{R}^{n}\right))$, a state is entirely determined by its values on the Weyl unitaries, which allows to define the characteristic function of $\varphi$ at argument $x\in\mathbb{R}^{2n}$ as
$\hat{W}\left[\varphi\right]\left(x\right):=\varphi\left(W\left(x\right)\right).$ (1.10)
Consider a real positive definite symmetric $2n\times 2n$ matrix $\Sigma$ satisfying
$\frac{1}{4}\left(D\left(x,y\right)\right)^{2}\leq\left\langle x,\Sigma x\right\rangle\left\langle y,\Sigma y\right\rangle\text{, }x,y\in\mathbb{R}^{2n}.$ (1.11)
Then there exists a unique state $\varphi\left(0,\Sigma\right)$ on $\mathcal{L}(L_{2}\left(\mathbb{R}^{n}\right))$ with characteristic function
$\hat{W}\left[\varphi\left(0,\Sigma\right)\right]\left(x\right)=\exp\left(-\frac{1}{2}\left\langle x,\Sigma x\right\rangle\right)\text{, }x\in\mathbb{R}^{2n}$ (1.12)
([Pet90], Theorem 3.4). Such states are called centered Gaussian (or quasifree) with covariance matrix $\Sigma$. The inequality (1.11) is required by Heisenberg’s uncertainty relation ([Hol11], Theorem 5.5.1).
1.3 Shift invariant states
In a centered Gaussian state $\varphi\left(0,\Sigma\right)$, every observable $R\left(x\right)=\mathbf{R}x$ has a normal distribution
$R\left(x\right)\sim N\left(0,\left\langle x,\Sigma x\right\rangle\right).$ (1.13)
Define the two vectors $\mathbf{R}_{s}:=\left(Q_{1+s},\ldots,Q_{n-1+s},P_{1+s},\ldots,P_{n-1+s}\right)$, $s=0,1$. The state $\varphi\left(0,\Sigma\right)$ is shift invariant if for every $t\in\mathbb{R}^{2(n-1)}$ the observables
$R_{s}\left(t\right)=\mathbf{R}_{s}t\text{, }s=0,1$
have the same distribution. It is easily seen that this implies shift invariance for the one mode subsystem $\left(Q_{1,}P_{1}\right)$, and also shift invariance for any $r$-mode subsystem $\left(Q_{1,},\ldots,Q_{r,}P_{1},\ldots,P_{r}\right)$, $1\leq r<n$. It follows that the covariance matrix $\Sigma$ is such that all four $r\times r$ submatrices in
$\Sigma=\left(\begin{array}[c]{cc}\Sigma_{11}&\Sigma_{12}\\ \Sigma_{12}^{\prime}&\Sigma_{22}\end{array}\right)$
are Toeplitz. Equivalently, if $\check{\Sigma}$ is the permutation of $\Sigma\$such that for $\mathbf{\check{R}}:=(Q_{1},P_{1},\ldots,Q_{n},P_{n},)$ we have
$\mathbf{\check{R}}x\sim N\left(0,\left\langle x,\check{\Sigma}x\right\rangle\right)\text{, }x\in\mathbb{R}^{2n}$ (1.14)
then $\check{\Sigma}$ is block Toeplitz, i.e. it is of form $\Sigma=\left(\Sigma_{j-k}^{0}\right)_{j,k=1}^{n}$ where $\left\{\Sigma_{k}^{0}\right\}_{k=0}^{n-1}$ is a sequence of $2\times 2$ matrices. The block Toeplitz structure is familiar in the statistical theory for classical multivariate time series.
1.4 Gauge invariant states
The Weyl unitaries $W\left(x\right)$, $x\in\mathbb{R}^{2n}$ can equivalently be indexed by complex $u\in\mathbb{C}^{n}$ such that $V\left(u\right):=W\left(\underline{u}\right)$ where $\underline{u}:=\left(-\operatorname{Im}u\right)\oplus\operatorname{Re}u$, whereupon the CCR relation (1.9) writes as
$V\left(u\right)V\left(v\right)=V\left(u+v\right)\exp\left(-\frac{i}{2}\operatorname{Im}\left\langle u,v\right\rangle\right),\;u,v\in\mathbb{C}^{n}.$ (1.15)
A state $\rho$ is gauge invariant if for every $z\in\mathbb{C}$, $\left|z\right|=1$ one has $\rho\left(V\left(zu\right)\right)=\rho\left(V\left(u\right)\right),$ $u\in\mathbb{C}^{n}$. A quasifree state $\rho\left(0,\Sigma\right)$ is gauge invariant if and only if
$\left\langle\underline{zu},\Sigma\underline{zu}\right\rangle=\left\langle\underline{u},\Sigma\underline{u}\right\rangle,u\in\mathbb{C}^{n},z\in\mathbb{C},\left|z\right|=1$
or equivalently, if there exists a self-adjoint positive operator $A$ on $\mathbb{C}^{n}$ such that
$\left\langle\underline{u},\Sigma\underline{u}\right\rangle=\frac{1}{2}\left\langle u,Au\right\rangle,\;u\in\mathbb{C}^{n}.$
The matrix $A$ is called the symbol of $\varphi\left(0,\Sigma\right)$; it is related to the covariance matrix $\Sigma$ by
$\Sigma=\Sigma\left(A\right):=\frac{1}{2}\left(\begin{array}[c]{cc}\operatorname{Re}A&-\operatorname{Im}A\\ \operatorname{Im}A&\operatorname{Re}A\end{array}\right)$ (1.16)
where $\operatorname{Re}A$ is symmetric and $\operatorname{Im}A$ is antisymmetric ($\left(\operatorname{Im}A\right)^{\prime}=-\operatorname{Im}A$). Relation (1.11) then can be written
$\left(\operatorname{Im}\left\langle u,v\right\rangle\right)^{2}\leq\left\langle u,Au\right\rangle\left\langle v,Av\right\rangle\text{, }u,v\in\mathbb{C}^{n}.$ (1.17)
Upon setting $v=iu$, this implies $A\geq I$, and conversely every Hermitian $A\geq I$ satisfies (1.17) and thus is the symbol of an $n$-mode gauge invariant centered Gaussian state. For the gauge invariant centered Gaussian state $\rho=\varphi\left(0,\Sigma\right)$ with symbol $A$, covariance matrix $\Sigma\left(A\right)$ and characteristic function
$\hat{W}\left[\varphi\left(0,\Sigma\right)\right]\left(\underline{u}\right)=\rho\left(V\left(u\right)\right)=\exp\left(-\frac{1}{4}\left\langle u,Au\right\rangle\right)\text{, }u\in\mathbb{C}^{n}$ (1.18)
we write
$\rho=\varphi\left(0,\Sigma\left(A\right)\right)=\mathfrak{N}_{n}\left(0,A\right).$ (1.19)
With this notation we suggest an analogy to the $n$-variate centered normal distribution with covariance matrix $M$, usually written $N_{n}\left(0,M\right)$. Note that for one mode ($n=1$), a gauge invariant centered Gaussian state has symbol $a\in\mathbb{R}$, $a\geq 1$ and covariance matrix $\Sigma=aI_{2}/2$. Thus
$\mathfrak{N}_{1}\left(0,a\right)=\varphi\left(0,aI_{2}/2\right)$ (1.20)
is the vacuum state for $a=1$ and a thermal state for $a>1$. If $A$ is diagonal $A=\mathrm{diag}\left(a_{1},\ldots,a_{n}\right)>0$ then $\mathfrak{N}_{n}\left(0,A\right)$ is the $n$-fold tensor product of thermal states $\mathfrak{N}_{1}\left(0,a_{j}\right)$.
1.5 The asymptotic setup
The quantum statistical model for fixed $n$ is now given by a family of gauge invariant and shift invariant centered Gaussian states $\left\{\mathfrak{N}_{n}\left(0,A\right),A\in\mathfrak{A}_{n}\right\}$ where $\mathfrak{A}_{n}$ is a set of $n\times n$ complex Hermitian Toeplitz matrices with $A\geq I$. In accordance wí th the usage in classical statistics, the model might be described as a stationary quantum Gaussian time series. For asymptotic inference in that model, we assume that the $n\times n$ symbols $A=\left(a_{j,k}\right)_{j,k=1}^{n}$ are related to a given positive bounded measurable function $a:\left[-\pi,\pi\right]\rightarrow\mathbb{R}$ as follows:
$a_{jk}=a_{k-j},\;a_{k}=\frac{1}{2\pi}\int_{-\pi}^{\pi}\exp\left(-ik\omega\right)a\left(\omega\right)d\omega,\;j,k\in\mathbb{Z}.$ (1.21)
such that
$a\left(\omega\right)=\sum_{k=-\infty}^{\infty}a_{k}\phi_{k}\left(\omega\right)\text{ where }\phi_{k}\left(\omega\right)=\exp\left(ik\omega\right)\text{, }\omega\in\mathbb{R}\text{.}$ (1.22)
Here $a_{j}$ are analogs of the autocovariances of a classical stationary complex valued time series, fulfilling $\overline{a_{k}}=a_{-k}$. Accordingly the function $a\left(\omega\right)$ may be described as the quantum spectral density. We assume $a$ to be real and fulfilling $a\geq 1$, and we write $A_{n}\left(a\right)$ for the Hermitian Toeplitz matrix generated by (1.21) for given $a$. We then have $A_{n}\left(a\right)\geq I$ (cf. Lemma 2.10 below); our quantum statistical model is now a family of states
$\left\{\mathfrak{N}_{n}\left(0,A_{n}\left(a\right)\right),a\in\Theta\right\}$ (1.23)
where $\Theta$ is a family of quantum spectral densities on $\left[-\pi,\pi\right]$ fulfilling $a\geq 1$. Note that if $f$ is a real function with $f\geq 0$ on $\left[-\pi,\pi\right]$ which is even (i.e. symmetric, $f\left(\omega\right)=f\left(-\omega\right)$) then the matrix $A_{n}\left(2\pi f\right)$ is real symmetric nonnegative definite, i.e. it is the covariance matrix of a real random vector. As $A_{n}\left(2\pi f\right)$ is also a sequence of Toeplitz matrices, it would describe the standard setup for asymptotic inference in classical stationary real valued time series [BD91], [Dzh86], [GNZ10]. Indeed comparing (1.21) with (1.4), we see that $a_{j}=\gamma_{j}$ if in (1.21) we set $a\left(\omega\right)=2\pi f\left(\omega\right),$ $\omega\in\left[-\pi,\pi\right]$. Our asymptotic model (1.23), where the spectral density $a$ is the parameter, is a quantum analog of a classical time series, involving the symbol matrices as analogs of covariance matrices. The model has first been treated in [Mos09] in the problem of discrimination between two spectral densities $a_{1},a_{2}$. There the quantum Chernoff bound has been computed for the specified quantum Gaussian models, based on the general form of the quantum Chernoff bound as previously found in [NS09] and [ANSV08].
1.6 Quantum Le Cam distance
We follow [GJ07] for defining the quantum analog of the $\Delta$-distance (1.3). So far the quantum Gaussian states $\mathfrak{N}_{n}\left(0,A\right)$ have been defined on the von Neumann algebra $\mathcal{L}\left(L_{2}\left(\mathbb{R}^{n}\right)\right)$, but in order to incorporate classical families of probability distributions into this framework, one needs to consider commutative von Neumann algebras defined by spaces $L^{\infty}\left(\mu\right)$ of functions on a $\sigma$-finite measure space $\left(X,\Omega,\mu\right)$. In our appendix section A.1 we clarify how states on a von Neumann algebra $\mathcal{A}$ can be understood as elements of the predual $\mathcal{A}_{\ast}$ of $\mathcal{A}$. The predual $\mathcal{A}_{\ast}$ is a Banach space with norm $\left\|\cdot\right\|_{1}$ such that $\mathcal{A}$ is its dual Banach space, and the states $\varphi$ are positive elements of $\mathcal{A}_{\ast}$ which fulfill $\left\|\varphi\right\|_{1}=1$. In the case $\mathcal{A}=\mathcal{L}\left(L_{2}\left(\mathbb{R}^{n}\right)\right)$, it is well known that a state $\varphi$ has a density operator $\rho_{\varphi}$ (a positive operator on $L_{2}\left(\mathbb{R}^{n}\right)$ with unit trace) such that
$\varphi\left(V\left(x\right)\right)=\mathrm{Tr\;}\rho_{\varphi}V\left(x\right),x\in\mathbb{C}^{n}.$
In that case $\left\|\varphi\right\|_{1}=\mathrm{Tr\;}\rho_{\varphi}=1$ and for states $\varphi,\sigma$, the distance
$\left\|\varphi-\sigma\right\|_{1}=\mathrm{Tr\;}\left|\rho_{\varphi}-\rho_{\sigma}\right|$
is the usual trace distance. In the case $\mathcal{A}=L^{\infty}\left(\mu\right)$, states are positive elements $f$ of $L^{1}\left(\mu\right)$ fulfilling $\left\|f\right\|_{1}=\int fd\mu=1$, i.e. probability density functions, and for states $f,g$ on $L^{\infty}\left(\mu\right)$, the distance
$\left\|f-g\right\|_{1}=\int\left|f-g\right|d\mu$
is the usual $L^{1}$-distance.
A quantum statistical experiment $\mathcal{E}=\left\{\mathcal{A},\rho_{\theta},\theta\in\Theta\right\}$ is given by a family of states $\rho_{\theta},\theta\in\Theta$ on a von Neumann algebra $\mathcal{A}$ where $\rho_{\theta}\in\mathcal{A}_{\ast}$. As a regularity condition, it is assumed that experiments are homogeneous and in reduced form (cf. Subsection A.1.11). Let $\mathcal{F}:=\left\{\mathcal{B},\sigma_{\theta},\theta\in\Theta\right\}$ be another quantum statistical experiment, indexed by the same parameter $\theta$. The deficiency $\delta\left(\mathcal{E},\mathcal{F}\right)$ is defined as
$\delta\left(\mathcal{E},\mathcal{F}\right):=\inf_{\alpha}\sup_{\theta}\left\|\sigma_{\theta}\circ\alpha-\rho_{\theta}\right\|_{1}$ (1.24)
where the infimum is taken over all quantum channels $\alpha:\mathcal{B}\rightarrow\mathcal{A}$ (see Appendix, A.1 for the definition of channels). The channels $\alpha$ are certain linear and (completely) positive maps between the von Neumann algebras; they give rise to quantum state transitions (TP-CP maps) $T:\mathcal{A}_{\ast}\rightarrow\mathcal{B}_{\ast}$ via the duality (A.10). If $\mathcal{A}$ and $\mathcal{B}$ are of type $L^{\infty}\left(\mu_{i}\right)$, $i=1,2$ then the TP-CP maps are transitions in the sense of Le Cam between dominated families of probability measures, which under regularity conditions are given by Markov kernels (cf. (A.13)). In the mixed case where $\mathcal{B=}L^{\infty}\left(\mu\right)$ and $\mathcal{A=L}\left(\mathcal{H}\right),$ the channel $\alpha$ is an observation channel (measurement) which arises from a POVM (positive operator valued measure), cf. Subsection A.1.9.
The Le Cam distance between $\mathcal{E}$ and $\mathcal{F}$ is
$\Delta\left(\mathcal{E},\mathcal{F}\right)=\max\left(\delta\left(\mathcal{E},\mathcal{F}\right),\delta\left(\mathcal{F},\mathcal{E}\right)\right).$ (1.25)
We say that $\mathcal{E}$ is more informative than $\mathcal{F}$ if $\delta\left(\mathcal{E},\mathcal{F}\right)=0$; if the reverse also holds (i.e. $\Delta\left(\mathcal{E},\mathcal{F}\right)=0$) then $\mathcal{E},\mathcal{F}$ are said to be statistically equivalent.
Consider now sequences of experiments, where the algebras and states depend on $n$, but the parameter space $\Theta$ remains fixed. A sequence $\mathcal{E}_{n}=\left\{\mathcal{A}_{n},\rho_{n,\theta},\theta\in\Theta\right\}$ is said to be asymptotically more informative than $\mathcal{F}_{n}=\left\{\mathcal{B}_{n},\sigma_{n,\theta},\theta\in\Theta\right\}$ if
$\delta\left(\mathcal{E}_{n},\mathcal{F}_{n}\right)\rightarrow 0\text{ as }n\rightarrow\infty.$
We write $\mathcal{E}_{n}\succsim\mathcal{F}_{n}$ in this case. If the reverse also holds, i.e. if
$\Delta\left(\mathcal{E}_{n},\mathcal{F}_{n}\right)\rightarrow 0\text{ as }n\rightarrow\infty.$
then $\mathcal{E}_{n}$ and $\mathcal{F}_{n}$ are said to be asymptotically equivalent, written $\mathcal{E}_{n}\approx\mathcal{F}_{n}$ .
As to the statistical meaning of the relation $\mathcal{E}_{n}\succsim\mathcal{F}_{n}$, assume that $\mathcal{E}_{n}=\left\{\mathcal{A}_{n},\rho_{n,\theta},\theta\in\Theta\right\}$ and $\mathcal{F}_{n}=\left\{\mathcal{B}_{n},\sigma_{n,\theta},\theta\in\Theta\right\}$. Then there is a sequence of TP-CP maps (transitions) between preduals $T_{n}:\mathcal{A}_{n\ast}\rightarrow\mathcal{B}_{n\ast}$ such that
$\sup_{\theta}\left\|\sigma_{n,\theta}-T_{n}\left(\rho_{n,\theta}\right)\right\|_{1}\rightarrow 0.$ (1.26)
Assume that statistical decisions are to be made in the experiment $\mathcal{F}_{n}$. Let $M_{n}$ be an observation channel (measurement) to be applied in $\mathcal{F}_{n}$, such that $M_{n}:\mathcal{B}_{n\ast}\rightarrow L^{1}\left(\nu\right)$ where $\nu$ is a sigma-finite measure on $\left(X,\Omega\right)$. Then $p_{n,\theta}:=M_{n}\left(\sigma_{n,\theta}\right)$ is a $\nu$-probability density on $\left(X,\Omega\right)$, and combining the transitions $M_{n}$ and $T_{n}$, we obtain a $\nu$-probability density $p_{n,\theta}^{\prime}:=M_{n}\left(T_{n}\left(\rho_{n,\theta}\right)\right)$. Then by the contraction property (A.9) of the state transition $M_{n}$
$\sup_{\theta}\left\|p_{n,\theta}-p_{n,\theta}^{\prime}\right\|_{1}\leq\sup_{\theta}\left\|\sigma_{n,\theta}-T_{n}\left(\rho_{n,\theta}\right)\right\|_{1}\rightarrow 0.$ (1.27)
Let a set of $\Omega$-measurable loss functions $W_{n,\theta}:X\rightarrow\left[0,1\right]$, $\theta\in\Theta$ be given. Then a measurement $M_{n}$ as above can be interpreted as a (randomized) decision rule in experiment $\mathcal{F}_{n},$ where the aim is to make $\int W_{n,\theta}p_{n,\theta}d\nu$ small for every $\theta$ (or small in a worst case sense). Then (1.27) implies
$\sup_{\theta}\left|\int W_{n,\theta}p_{n,\theta}d\nu-\int W_{n,\theta}p_{n,\theta}^{\prime}d\nu\right|\rightarrow 0.$
In other words, if the sequence $\mathcal{E}_{n}$ is asymptotically more informative than $\mathcal{F}_{n}$ ($\mathcal{E}_{n}\succsim\mathcal{F}_{n}$) then for every randomized decision rule in $\mathcal{F}_{n}$ there exists one in $\mathcal{E}_{n}$ which is asymptotically as good, uniformly in $\theta\in\Theta$.
In applications, when a parameter $\theta$ is to be estimated by $\hat{\theta}_{n}$ , the loss functions $W_{n,\theta}$ are typically of the form $W_{n,\theta}\left(\hat{\theta}_{n}\right)=\ell\left(\eta_{n}\left\|\hat{\theta}_{n}-\theta\right\|\right)$ where $\ell:\left[0,\infty\right)\rightarrow\left[0,1\right]$ is a monotone function and $\eta_{n}\rightarrow\infty$ is a norming sequence. Then the relation $\mathcal{E}_{n}\succsim\mathcal{F}_{n}$ means that estimators in $\mathcal{F}_{n}$ cannot be asymptotically better than those in $\mathcal{E}_{n}$, i. e. the relation provides lower asymptotic risk bounds. If also the converse $\mathcal{F}_{n}\succsim\mathcal{E}_{n}$ can be shown then risk bounds attainable in $\mathcal{E}_{n}$ can also be attained in $\mathcal{F}_{n}$. Applications to optimal estimation of the quantum spectral density will be discussed in a forthcoming version of this preprint.
1.7 Main theorems
For any set $\Theta$ of quantum spectral densities, i.e. real functions $a$ on $\left[-\pi,\pi\right]$ such that $a\left(\omega\right)\geq 1$, $\omega\in\left[-\pi,\pi\right]$ consider the quantum statistical experiment
$\mathcal{E}_{n}\left(\Theta\right):=\left\{\mathfrak{N}_{n}\left(0,A_{n}\left(a\right)\right),a\in\Theta\right\}$ (1.28)
where $A_{n}\left(a\right)$ is the $n\times n$ symbol matrix pertaining to $a$. Define also a corresponding classical geometric regression experiment $\mathcal{G}_{n}\left(\Theta\right)$ as follows. For any function $a\in\Theta$ define a set of functionals (local averages on $\left[-\pi,\pi\right]$) as
$J_{j,n}\left(a\right)=n\int_{\left(j-1\right)/n}^{j/n}a\left(2\pi\left(x-1/2\right)\right)dx\text{.}$ (1.29)
Also consider the geometric distribution $\mathrm{Geo}\left(p\right)$ with probabilities $\left(1-p\right)p^{j}$, $j=0,1,\ldotswhere$ the parameter $p\in\left(0,1\right)$ depends on some $\lambda>1$ via $p\left(\lambda\right)=\left(\lambda-1\right)/\left(\lambda+1\right)$. Define
$\mathcal{F}_{n}\left(\Theta\right):=\left\{{\displaystyle\bigotimes\limits_{j=1}^{n}}\mathrm{Geo}\left(p\left(J_{j,n}\left(a\right)\right)\right),a\in\Theta\right\}$ (1.30)
Consider the set of quantum spectral densities $a$ defined as follows: the set of real functions on $\left[-\pi,\pi\right]$, such that for some $\alpha>0$, $M>1$
$\Theta_{1}\left(\alpha,M\right)$ $:=\left\{a:\text{ }\left|a_{0}\right|^{2}+\sum_{j=-\infty}^{\infty}j^{2\alpha}\left|a_{j}\right|^{2}\leq M\right\}\cap\mathcal{L}_{M},$ (1.31)
$\mathcal{L}_{M}$ $:=\left\{a:a\left(\omega\right)\geq 1+M^{-1}\text{, }\omega\in\left[-\pi,\pi\right]\right\},$ (1.32)
where $a_{j}$ are defined by (1.21).
Theorem 1.1
If $\Theta=\Theta_{1}\left(\alpha,M\right)$ for some $\alpha>1/2,$ $M>1$ then
$\delta\left(\mathcal{E}_{n}\left(\Theta\right),\mathcal{F}_{n}\left(\Theta\right)\right)\rightarrow 0\text{ as }n\rightarrow\infty,$
i.e. $\mathcal{F}_{n}\left(\Theta\right)$ is asymptotically more informative than $\mathcal{E}_{n}\left(\Theta\right)$: $\mathcal{E}_{n}\left(\Theta\right)\lesssim\mathcal{F}_{n}\left(\Theta\right)$.
Let us further introduce an experiment of the type ”signal in Gaussian white noise” on the interval $\left[-\pi,\pi\right]$. Consider the function
$\mathrm{arc\cosh}\left(x\right)=\log\left(x+\sqrt{x^{2}-1}\right),x>1$
and let $Q_{n}\left(a\right)$ be the distribution of the stochastic process $Y_{\omega},\omega\in\left[-\pi,\pi\right]$ given by the stochastic differential equation
$dY_{\omega}=\mathrm{arc\cosh}\left(a\left(\omega\right)\right)d\omega+\left(2\pi/n\right)^{1/2}dW_{\omega}\text{, }\omega\in\left[-\pi,\pi\right].$ (1.33)
and where $dW_{\omega},$ $\omega\in\left[-\pi,\pi\right]$ is Gaussian white noise and $Y_{\omega}=\int_{-\pi}^{\omega}dY_{\omega}$. Here $Q_{n}\left(a\right)$ is a distribution on the measurable space $\left(C_{\left[-\pi,\pi\right]},\mathcal{B}\left(C_{\left[-\pi,\pi\right]}\right)\right)$. For $\Theta=\Theta_{1}\left(\alpha,M\right)$ consider the experiment $\mathcal{G}_{n}\left(\Theta\right)=\left\{Q_{n}\left(a\right),\;a\in\Theta\right\}$.
Theorem 1.2
If $\Theta=\Theta_{1}\left(\alpha,M\right)$ for some $\alpha>1,$ $M>1$ then
$\Delta\left(\mathcal{F}_{n}\left(\Theta\right),\mathcal{G}_{n}\left(\Theta\right)\right)\rightarrow 0\text{ as }n\rightarrow\infty$
i.e. $\mathcal{F}_{n}\left(\Theta\right)$ and $\mathcal{G}_{n}\left(\Theta\right)$ are asymptotically equivalent: $\mathcal{F}_{n}\left(\Theta\right)\approx\mathcal{G}_{n}\left(\Theta\right)$
This claim essentially follows from the results of [GN98]. It implies that for $\alpha>1$, for the quantum time series, the white noise model $\mathcal{G}_{n}^{\prime}\left(\Theta\right)$ is an upper information bound as well. Note that the function $\mathrm{arc\cosh}$ is the analog of the log-transformation of the spectral density in (1.5).
Converse results can be established if the parameter space is restricted to be finite dimensional. For a nonnegative integer $d$ and some $M>1$ define
$\Theta_{2}\left(d,M\right):=\left\{a:\text{ }\sum_{j=-d}^{d}\left|a_{j}\right|^{2}\leq M,\;a_{j}=0\text{, }\left|j\right|>d\right\}\cap\mathcal{L}_{M}.$ (1.34)
Then the symbol matrices $A_{n}\left(a\right)$ are banded Toeplitz and the quantum states $\mathfrak{N}_{n}\left(0,A_{n}\left(a\right)\right)$ form a $d$-dependent quantum time series.
Theorem 1.3
If $\Theta=\Theta_{2}\left(d,M\right)$ for an integer $d\geq 0$ and some $M>1$ then
$\delta\left(\mathcal{E}_{n}\left(\Theta\right),\mathcal{G}_{n}\left(\Theta\right)\right)\rightarrow 0\text{ as }n\rightarrow\infty,$
i.e. $\mathcal{G}_{n}\left(\Theta\right)$ is asymptotically less informative than $\mathcal{E}_{n}\left(\Theta\right)$: $\mathcal{G}_{n}\left(\Theta\right)\lesssim\mathcal{E}_{n}\left(\Theta\right)$.
Clearly $\Theta_{2}\left(d,M\right)\subset\Theta_{1}\left(\alpha,M^{\prime}\right)$ for $\alpha=1$ and some $M^{\prime}>M$ (cf. Lemma 3.20 below). This implies
Corollary 1.4
If $\Theta=\Theta_{2}\left(d,M\right)$ for $d\geq 0$ and $M>1$ then
$\Delta\left(\mathcal{E}_{n}\left(\Theta\right),\mathcal{G}_{n}\left(\Theta\right)\right)\rightarrow 0\text{ as }n\rightarrow\infty,$
i.e. $\mathcal{G}_{n}\left(\Theta\right)$ and $\mathcal{E}_{n}\left(\Theta\right)$ are asymptotically equivalent: $\mathcal{E}_{n}\left(\Theta\right)\approx\mathcal{G}_{n}\left(\Theta\right)$.
The proofs of Theorems 1.1, 1.2 and 1.3 are in Subsections 2.6, 2.7 and 3.7, respectively.
In a forthcoming version of this preprint, we will also identify a quantum analog of the periodogram and provide optimal parametric and nonparametric estimates of the quantum spectral density.
Further notation.
Consider quantum statistical experiments $\mathcal{E}=\left\{\mathcal{A},\rho_{\theta},\theta\in\Theta\right\}$ and $\mathcal{F}:=\left\{\mathcal{B},\sigma_{\theta},\theta\in\Theta\right\}$ having the same parameter space. For the special case that $\mathcal{A}=\mathcal{B}$ define their trace norm distance
$\Delta_{0}\left(\mathcal{E},\mathcal{F}\right)=\sup_{\theta}\left\|\rho_{\theta}-\sigma_{\theta}\right\|_{1}.$
In general we will use the following notation involving quantum experiments $\mathcal{E}$ and $\mathcal{F}$.
$\mathcal{E}$ $\mathbf{\preceq}$ $\mathcal{F}$ ($\mathcal{F}$ more informative than $\mathcal{E}$): $\delta\left(\mathcal{F},\mathcal{E}\right)=0$
$\mathcal{E}$ $\mathcal{\sim}$ $\mathcal{F}$ (equivalent): $\Delta\left(\mathcal{E},\mathcal{F}\right)=0$
$\mathcal{E}_{n}$ $\simeq$ $\mathcal{F}_{n}$ (asymptotically trace norm equivalent): $\Delta_{0}\left(\mathcal{F}_{n},\mathcal{E}_{n}\right)\rightarrow 0$
[…]
$\mathcal{E}_{n}$ $\approx$ $\mathcal{F}_{n}$ (asymptotically equivalent): $\Delta\left(\mathcal{F}_{n},\mathcal{E}_{n}\right)\rightarrow 0$
Note that ”more informative” above is used in the sense of a semi-ordering, i.e. its actual meaning is ”at least as informative”. If $\mathcal{E},\mathcal{F}$ are classical experiments, where the trace norm distance is a multiple of the total variation distance between probability measures, the relation $\mathcal{E}_{n}\simeq\mathcal{F}_{n}\mathcal{\ }$will also be described as asymptotic equivalence in total variation.
2 Upper informativity bound
2.1 Gaussian states on symmetric Fock space
Let $\mathcal{H}$ be a complex separable Hilbert space. Let $\vee^{m}\mathcal{H}$ denote the $m$-fold symmetric tensor power, that is, the subspace of $\mathcal{H}^{\otimes m}$ consisting of vectors which are symmetric under permutations of the tensors, with $\vee^{0}\mathcal{H}:=\mathbb{C}$. The Fock space over $\mathcal{H}$ is the Hilbert space
$\mathcal{F}\left(\mathcal{H}\right):={\displaystyle\bigoplus_{m\geq 0}}\vee^{m}\mathcal{H}.$
For each $x\in\mathcal{H}$ let
$x_{F}:={\displaystyle\bigoplus_{m\geq 0}}\frac{1}{\sqrt{m!}}x^{\otimes m}$ (2.1)
denote the corresponding exponential vector (or Fock vector). The exponential vectors are linearly independent and their linear span is dense in $\mathcal{F}\left(\mathcal{H}\right)$. The Weyl unitaries $V\left(x\right)$, $x\in\mathcal{H}$ are defined by their action on exponential vectors as
$V\left(x\right)y_{F}:=\left(y+2^{-1/2}x\right)_{F}\exp\left(-\frac{1}{4}\left\|x\right\|^{2}-2^{-1/2}\left\langle x,y\right\rangle\right)\text{, }y\in\mathcal{H}.$ (2.2)
These can be seen to satisfy the relation
$V\left(x\right)V\left(y\right)=V\left(x+y\right)\exp\left(-\frac{i}{2}\operatorname{Im}\left\langle x,y\right\rangle\right)$ (2.3)
and for $\mathcal{H}=\mathbb{C}^{n}$ this coincides with the CCR (1.15) stated in the Schrödinger representation. Denote by $\left\{V_{j}\left(x\right),x\in\mathbb{C}^{n}\right\}$, $j=1,2$ these two versions of the Weyl unitaries ($i=2$ corresponding to (2.2) ); since both sets of operators are irreducible, there is a linear isometric map $U$ $L_{2}\left(\mathbb{R}^{n}\right)\mapsto\mathcal{F}\left(\mathbb{C}^{n}\right)$ such that
$V_{1}\left(x\right)=U^{\ast}V_{2}\left(x\right)U\text{, }x\in\mathbb{C}^{n}\text{.}$
The corresponding generated C*-algebras are hence *-isomorphic and are denoted by $CCR\left(\mathbb{C}^{n}\right)$; henceforth in this section we will work with the Fock representation $V\left(x\right)=V_{2}\left(x\right)$ of (2.2). A state $\varphi$ on $CCR\left(\mathcal{H}\right)$ is a positive linear functional $\varphi:CCR\left(\mathcal{H}\right)\mapsto\mathbb{C}$ that takes the value $1$ on the unit on $CCR\left(\mathcal{H}\right)$.
Let $B\in\mathcal{B}\left(\mathcal{H}\right)$ be a bounded operator on $\mathcal{H}$, and let $\vee^{m}B$ be the restriction of $B^{\otimes m}$ to $\vee^{m}\mathcal{H}$. The Fock operator $B_{F}$ corresponding to $B$ is
$B_{F}:={\displaystyle\bigoplus_{m\geq 0}}\vee^{m}B$
with an appropriate domain $\mathcal{D}\left(B_{F}\right)$ (cf. [Mos09], Appendix for more details). Then $\left(Bx\right)_{F}=B_{F}x_{F}$ holds for exponential vectors $x_{F}$, and for $A\in\mathcal{B}\left(\mathcal{H}\right)$, the relation
$A_{F}B_{F}=\left(AB\right)_{F}$ (2.4)
holds on a dense subset of $\mathcal{F}\left(\mathcal{H}\right)$. Then, for a gauge invariant centered Gaussian state with symbol matrix $A$, the density operator on $\mathcal{F}\left(\mathbb{C}^{n}\right)$ can be described as follows (cp. [Mos09], A5):
$\mathfrak{N}_{n}\left(0,A\right)=\frac{2^{n}}{\det\left(A+I\right)}\left(\frac{A-I}{A+I}\right)_{F}.$ (2.5)
A proof is given in subsection A.2.2 below.
2.2 Distance of states in terms of symbols
Our model is the quantum statistical experiment $\mathcal{E}_{n}\left(\Theta_{1}\left(\alpha,M\right)\right)$ described in Theorem 1.1. To characterize the parameter space $\Theta_{1}\left(\alpha,M\right)$ for $\alpha>1/2$, define any real valued $a\in L_{2}(-\pi,\pi)$ and its Fourier coefficients (1.21)
$\left|a\right|_{2,\alpha}^{2}:=\sum_{k=-\infty}^{\infty}\left|k\right|^{2\alpha}\left|a_{k}\right|^{2},\;\left\|a\right\|_{2,\alpha}^{2}:=a_{0}^{2}+\left|a\right|_{2,\alpha}^{2}$ (2.6)
provided the r.h.s. is finite. The set of real functions
$W^{\alpha}(M)=\left\{a\in L_{2}(-\pi,\pi):\left\|a\right\|_{2,\alpha}^{2}\leq M\right\}.$ (2.7)
then describes a ball in the scale of periodic fractional Sobolev spaces with smoothness coefficient $\alpha$. Note that for $\alpha>1/2$, by an embedding theorem ([GNZ09], Lemma 5.6) , functions in $W^{\alpha}(M)$ are also uniformly bounded. For $M>0$, define a set of real valued functions on $\pi,\pi]$
$\mathcal{F}_{M}=\left\{a:M^{-1}\leq a(\omega)-1,\omega\in[-\pi,\pi]\right\}.$ (2.8)
Then for the parameter space $\Theta_{1}\left(\alpha,M\right)$ of Theorem 1.1 we have
$\Theta_{1}\left(\alpha,M\right)=W^{\alpha}(M)\cap\mathcal{F}_{M}.$ (2.9)
Therefore we can assume there exists $C=C_{M,\alpha}>0$ such that
$1+C^{-1}\leq a\left(\omega\right)\leq C,\text{ }\omega\in\left[-\pi,\pi\right]$ (2.10)
holds for all $a\in\Theta_{1}\left(\alpha,M\right)$. Introducing notation
$Q:=\left(A-I\right)/2\text{, }R:=\frac{Q}{Q+I}$ (2.11)
we obtain from (2.5)
$\mathfrak{N}_{n}\left(0,A\right)$ $=\frac{1}{\det\left(I+Q\right)}\left(\frac{Q}{I+Q}\right)_{F}$ (2.12)
$=\frac{1}{\det\left(I+Q\right)}R_{F}.$ (2.13)
In the sequel we will approximate a state $\mathfrak{N}_{n}\left(0,A_{1}\right)$, given by symbol $A_{1}$ by the corresponding state for a symbol $A_{2}$. Specifically, $A_{1}$ will be taken as the Hermitian Toeplitz matrix $A_{n}\left(a\right)$ and $A_{2}$ will be a (truncated) circulant matrix. We assume that $A_{i}$, $i=1,2$ are Hermitian $n\times n$ such that there exists $c>0,$ independent of $n$, such that
$\lambda_{\min}\left(A_{i}-I\right)\geq c.$
This assumption will be justified later for the cases at hand, on the basis of (2.10). In the Fock representation (2.5) it then follows from Lemma A.2 that
$\lambda_{\min}\left(\frac{A_{i}-I}{I+A_{i}}\right)_{F}>0$
(cp. (A.35) below), hence the states $\mathfrak{N}_{n}\left(0,A_{i}\right)$ are faithful.
We begin with a bound for the trace norm in terms of relative entropy. The trace norm between states $\rho,\sigma$ is defined as
$\left\|\rho-\sigma\right\|_{1}:=\mathrm{Tr}\;\left|\rho-\sigma\right|.$
For finite dimensional states $\rho$ and $\sigma$, the relative entropy is
$S\left(\rho||\sigma\right)=\left\{\begin{tabular}[c]{l}$\mathrm{Tr}\;\rho\left(\log\rho-\log\sigma\right)$ if{ supp}$\;\sigma\supseteq\;${supp}$\;\rho$\\ $\infty$ otherwise.\end{tabular}\ \ \ \ \ \right$. (2.14)
This formula extends to faithful Gaussian states with density operators $\rho,$ $\sigma$, (2.14), in the sense of agreeing with the definition of relative entropy for normal states on a von Neumann algebra ([Pet08], sec 3.4). As we argued above, both our states $\rho,\sigma$ are assumed faithful, so $\sigma\supseteq\;supp$$\rho$ holds and $K\left(\rho,\sigma\right)$ can be computed from the first line of (2.14). Then a quantum analog of Pinsker’s inequality holds ([OP93], Theorem 5.5): for the trace norm distance between $\rho$ and $\sigma$ one has
$\left\|\rho-\sigma\right\|_{1}^{2}\leq 2S\left(\rho||\sigma\right).$ (2.15)
Consider symbols $A_{j}$, $j=1,2$ and let $\rho_{j}=\mathfrak{N}_{n}\left(0,A_{j}\right)$, $j=1,2$ be the corresponding Gaussian states. Our purpose in this section is to obtain an upper bound on the trace norm distance in terms of the symbols, by using (2.15) and an appropriate upper bound on $S\left(\rho||\sigma\right)$.
For general Gaussian states, explicit expression for $S\left(\rho||\sigma\right)$ in terms of the first two moments have been obtained ([PLOB17] and references therein). Below we give a special formula which focuses on the zero mean gauge invariant case, and writes out $S\left(\rho||\sigma\right)$ directly in terms of the symbols rather than the covariance matrices.
Consider the relative entropy between two Bernoulli laws $\left(1-p_{j},p_{j}\right)$ with $p_{j}\in\left(0,1\right)$, $j=1,2$:
$S_{2}\left(p_{1}||p_{2}\right)=p_{1}\log\frac{p_{1}}{p_{2}}+\left(1-p_{1}\right)\log\frac{1-p_{1}}{1-p_{2}}.$
An analog for $n\times n$ Hermitian $R_{j}$ satisfying $0<R_{i}<I$ is
$S_{2}\left(R_{1}||R_{2}\right):=R_{1}\left(\log R_{1}-\log R_{2}\right)+\left(I-R_{1}\right)\left(\log\left(I-R_{1}\right)-\log\left(I-R_{2}\right)\right).$ (2.16)
Proposition 2.1
Let $A_{j}$, $i=1,2$ be Hermitian $n\times n$ such that $\lambda_{\min}\left(A_{j}-I\right)>0$, and let
$\rho_{j}=\mathfrak{N}_{n}\left(0,A_{j}\right)=\frac{2^{n}}{\det\left(I+A_{j}\right)}\left(\frac{A_{j}-I}{A_{j}+I}\right)_{F}.$
be the corresponding Gaussian states. Let $Q_{j}$ and $R_{j}$ be defined by
$Q_{j}:=\left(A_{j}-I\right)/2\text{, }R_{j}:=\frac{Q_{j}}{Q_{j}+I}=\frac{A_{j}-I}{A_{j}+I}\text{, }j=1,2.$
Then for the relative entropy one has
$S\left(\rho_{1}||\rho_{2}\right)=\mathrm{Tr}\;\left(I+Q_{1}\right)S_{2}\left(R_{1}||R_{2}\right)$ (2.17)
where $S_{2}\left(\cdot||\cdot\right)$ is defined by (2.16).
Proof. Assume a Gaussian state is given by $\rho=\frac{1}{\det\left(I+Q\right)}R_{F}$ according to (2.13). Then
$\log\rho$ $=-\log\det\left(I+Q\right)I_{F}+\log R_{F}$
$=-\log\det\left(I+Q\right)I_{F}+\oplus_{m=0}^{\infty}\log\vee^{m}R.$ (2.18)
Using Lemma A.4 we find
$\log\rho$ $=-\log\det\left(I+Q\right)I_{F}+\oplus_{m=0}^{\infty}\Gamma_{m}\left(\log R\right)$
$=-\log\det\left(I+Q\right)I_{F}+\Gamma\left(\log R\right)$
with the definition of $\Gamma\left(\log R\right)$ given in Lemma A.3. Setting $\rho=\rho_{2}$ and applying this lemma for the case $A=R_{1}$, $B=\log R_{2}$, we obtain
$\mathrm{Tr}\;\rho_{1}\log\rho_{2}$ $=\frac{1}{\det\left(I+Q_{1}\right)}\mathrm{Tr}\;\left(R_{1}\right)_{F}\left(-\log\det\left(I+Q_{2}\right)I_{F}+\Gamma\left(\log R_{2}\right)\right)$
$=-\log\det\left(I+Q_{2}\right)+\frac{1}{\det\left(I+Q_{1}\right)}\mathrm{Tr}\;\left(R_{1}\right)_{F}\Gamma\left(\log R_{2}\right)$
$=-\log\det\left(I+Q_{2}\right)+\frac{1}{\det\left(I+Q_{1}\right)}\frac{1}{\det\left(I-R_{1}\right)}\mathrm{Tr}\;\frac{R_{1}}{I-R_{1}}\log R_{2}.$ (2.19)
In view of (2.11) we have
$I-R_{1}$ $=I-\frac{Q_{1}}{I+Q_{1}}=\frac{I}{I+Q_{1}},$
$\det\left(I-R_{1}\right)$ $=1/\det\left(I+Q_{1}\right),$
$-\log\det\left(I+Q_{2}\right)$ $=\log\det\left(I-R_{2}\right)=\mathrm{Tr}\;\log\left(I-R_{2}\right).$
Applied to (2.19) this implies
$\mathrm{Tr}\;\rho_{1}\log\rho_{2}$ $=\mathrm{Tr}\;\log\left(I-R_{2}\right)+\mathrm{Tr}\;\frac{R_{1}}{I-R_{1}}\log R_{2}$ (2.20)
$=\mathrm{Tr}\;\log\left(I-R_{2}\right)+\mathrm{Tr}\left(I+Q_{1}\right)\;R_{1}\log R_{2}$
$=\mathrm{Tr}\;\left(I+Q_{1}\right)\left[\left(I-R_{1}\right)\log\left(I-R_{2}\right)+\;R_{1}\log R_{2}\right].$ (2.21)
For the case $\rho_{1}=\rho_{2}$ we obtain
$\mathrm{Tr}\;\rho_{1}\log\rho_{1}=\mathrm{Tr}\;\left(I+Q_{1}\right)\left[\left(I-R_{1}\right)\log\left(I-R_{1}\right)+R_{1}\log R_{1}\right]$ (2.22)
From (2.14), (2.21) and (2.22) we finally obtain
$S\left(\rho_{1}||\rho_{2}\right)=\mathrm{Tr}\;\rho_{1}\left(\log\rho_{1}-\log\rho_{2}\right)$
$=\mathrm{Tr}\;\left(I+Q_{1}\right)\left[R_{1}\left(\log R_{1}-\log R_{2}\right)+\left(I-R_{1}\right)\left(\log\left(I-R_{1}\right)-\log\left(I-R_{2}\right)\right)\right]$
$=\mathrm{Tr}\;\left(I+Q_{1}\right)S_{2}\left(R_{1}||R_{2}\right).$
Since $Q_{i}$ are positive definite $n\times n$ Hermitian, the matrices $R_{i}$ and $I-R_{i}=I/\left(Q_{i}+I\right)$ also have these properties; in particular
$0<R_{i}<I\text{, }i=1,2.$ (2.23)
Hence $S_{2}\left(R_{1}||R_{2}\right)$ defined by (2.17) is finite, and thus $S\left(\rho_{1}||\rho_{2}\right)$ is also finite. In order to achieve uniformity of estimates over the $R_{i}$ considered, we assume a strengthened version of (2.23): there exists $\lambda\in\left(1/2,1\right)$ such that
$\left(1-\lambda\right)I<R_{i}<\lambda I\text{, }\;i=1,2.$ (2.24)
It is immediate, in view of (2.11), that this condition is equivalent to each of the following two:
$\frac{1-\lambda}{\lambda}I$ $<Q_{i}<\frac{\lambda}{1-\lambda}I,$ (2.25)
$\left(\frac{2}{\lambda}-1\right)I$ $<A_{i}<\frac{1+\lambda}{1-\lambda}I.$ (2.26)
Also, (2.25) implies
$I+Q_{i}<\frac{1}{1-\lambda}I\text{, }\;i=1,2.$ (2.27)
Our next task is to estimate (2.17) in terms of the difference $H=R_{1}-R_{2}$. To that end we use an expansion
$\log R_{2}=\log\left(I-\left(I-R_{2}\right)\right)=-\sum_{k=1}^{\infty}\frac{1}{k}\left(I-R_{2}\right)^{k}.$
That is valid if $I-R_{2}$ has all eigenvalues contained in $\left(-1,1\right)$, which holds due to (2.23). Similarly we expand $\log R_{1}$ and obtain
$\log R_{1}-\log R_{2}$ $=\sum_{k=1}^{\infty}\frac{1}{k}\left[\left(I-R_{2}\right)^{k}-\left(I-R_{1}\right)^{k}\right]$ (2.28)
$=\sum_{k=1}^{\infty}\frac{1}{k}\left[\left(I-R_{1}+H\right)^{k}-\left(I-R_{1}\right)^{k}\right].$ (2.29)
Furthermore we obtain
$\left(I-R_{1}+H\right)^{k}-\left(I-R_{1}\right)^{k}=$
$=H\left(I-R_{1}\right)^{k-1}+\left(I-R_{1}\right)H\left(I-R_{1}\right)^{k-2}+\ldots+\left(I-R_{1}\right)^{k-1}H$ (2.30)
$+H^{2}\left(I-R_{1}\right)^{k-2}+H\left(I-R_{1}\right)H\left(I-R_{1}\right)^{k-3}+\ldots+\left(I-R_{1}\right)^{k-2}H^{2}$
$\ldots$
$+H^{k-1}\left(I-R_{1}\right)+H^{k-2}\left(I-R_{1}\right)H+\ldots+\left(I-R_{1}\right)H^{k-1}$
$+H^{k}.$
A similar expansion holds for the log terms in the second summand of (2.16): writing G=HG=-H, we have $R_{2}=R_{1}+G$ and
$\log\left(I-R_{1}\right)-\log\left(I-R_{2}\right)=\sum_{k=1}^{\infty}\frac{1}{k}\left[\left(R_{1}+G\right)^{k}-R_{1}^{k}\right],$ (2.31)
$\left(R_{1}+G\right)^{k}-R_{1}^{k}=$
$GR_{1}^{k-1}+R_{1}GR_{1}^{k-2}+\ldots+R_{1}^{k-1}G$ (2.32)
$+G^{2}R_{1}^{k-2}+GR_{1}GR_{1}^{k-3}+\ldots+R_{1}^{k-2}G^{2}$
$\ldots$
$+G^{k-1}R_{1}+G^{k-2}R_{1}G+\ldots+R_{1}G^{k-1}$
$+G^{k}.$
We also denote
$M_{1}$ $=\left(I+Q_{1}\right)R_{1},$
$M_{2}$ $=\left(I+Q_{1}\right)\left(I-R_{1}\right).$
Furthermore, for matrices $A$ we write the operator norm $\left|A\right|=\lambda_{\max}^{1/2}\left(A^{\ast}A\right)$, so that for Hermitian positive $A$ we have $\left|A\right|=\lambda_{\max}\left(A\right)$. The Hilbert-Schmidt norm is written $\left\|A\right\|_{2}=\left(\mathrm{Tr}\;A^{\ast}A\right)^{1/2}$. We then have
$\left\|AB\right\|_{2}$ $\leq\left|A\right|\left\|B\right\|_{2},$ (2.33)
$\left|AB\right|$ $\leq\left|A\right|\left|B\right|.$ (2.34)
Consider now the series expression for $S\left(\rho_{1}||\rho_{2}\right)$ given by (2.17) and the expansions (2.30), (2.32), i.e. the series obtained for $\mathrm{Tr}\;\left(I+Q_{1}\right)\;S_{2}\left(R_{1}||R_{2}\right)$. Consider first the question whether it converges absolutely.
To that end we denote the generic term in the expansion (2.30) by $T_{k,j,l}$, in such a way that
  • $k$ is as indicated, i.e it pertains to a term in the expansion of $\left(I-R_{1}+H\right)^{k}-\left(I-R_{1}\right)^{k}$, where $1\leq k<\infty$
  • $j$ is the order in $H$, i.e. the total number of factors $H$ (such that kjk-j is the total number of factors $I-R_{1}$), and $1\leq j\leq k$
[…]
In a similar way, we denote the generic term in the expansion (2.32) by $U_{k,j,l}$, in such a way that
  • $k$ is as indicated, i.e it pertains to a term in the expansion of $\left(R_{1}+G\right)^{k}-R_{1}^{k}$, where $1\leq k<\infty$
  • $j$ is the order in $G$, i.e. the total number of factors $G$ (such that kjk-j is the total number of factors $R_{1}$), and $1\leq j\leq k$
[…]
Lemma 2.2
For $\left\|H\right\|_{2}<1-\lambda$, with $\lambda$ from (2.24), the series
$\mathrm{Tr}\;\left(I+Q_{1}\right)\;S_{2}\left(R_{1}||R_{2}\right)=$
[…]
converges absolutely.
Proof. Consider the first series and all terms with $j=1$. Since $\left(I+Q_{1}\right)$ and $R_{1}$ are commuting and positive, we have by Cauchy-Schwartz, for $1\leq l\leq k$
$\left|\mathrm{Tr}\;M_{1}\;T_{k,1,l}\right|$ $=\left|\mathrm{Tr}\;\left(I+Q_{1}\right)R_{1}\left(I-R_{1}\right)^{k-1}H\right|$
$\leq\left\|\left(I+Q_{1}\right)R_{1}\left(I-R_{1}\right)^{k-1}\right\|_{2}\left\|H\right\|_{2}$
$\leq\lambda^{k}\left\|\left(I+Q_{1}\right)\right\|_{2}\left\|H\right\|_{2}$
where we used (2.33) and $\left|R_{1}\right|<\lambda$, $\left|I-R_{1}\right|<\lambda$ due to (2.24). Consequently
$\sum_{k=1}^{\infty}\sum_{l=1}^{k}\frac{1}{k}\left|\mathrm{Tr}\;M_{1}\;T_{k,1,l}\right|$ $\leq\sum_{k=1}^{\infty}\lambda^{k}\left\|\left(I+Q_{1}\right)\right\|_{2}\left\|H\right\|_{2}$
$=\frac{\lambda}{1-\lambda}\left\|\left(I+Q_{1}\right)\right\|_{2}\left\|H\right\|_{2}<\infty.$ (2.36)
Next consider all quadratic terms in $H$, i.e. the case $j=2$. The general form of such a term, with $k\geq 2$, is
$\mathrm{Tr}\;M_{1}\;T_{k,2,l}=\mathrm{Tr}\;\left(I+Q_{1}\right)R_{1}\left(I-R_{1}\right)^{a}H\left(I-R_{1}\right)^{b}H$
where $a$ and $b$ depend on $k$ and $l$, with a+b=k2a+b=k-2, $a,b\geq 0$. By Cauchy-Schwartz we obtain
$\left|\mathrm{Tr}\;M_{1}\;T_{k,2,l}\right|\leq$
$\left\|\left(I+Q_{1}\right)R_{1}\left(I-R_{1}\right)^{a}H\right\|_{2}\cdot\left\|\left(I-R_{1}\right)^{b}H\right\|_{2}.$ (2.37)
Setting $\beta:=1/\left(1-\lambda\right)$ and using the bound (2.27), the first factor above can be upper bounded as $\beta\lambda^{a+1}\left\|H\right\|_{2}$. Similarly the second factor in (2.37) can be bounded by $\lambda^{b}\left\|H\right\|_{2}$. As a result we get
$\left|\mathrm{Tr}\;M_{1}\;T_{k,2,l}\right|\leq\beta\lambda^{k-1}\left\|H\right\|_{2}^{2}.$ (2.38)
Thus for the totality of second order terms we have
[…]
$=\beta\left\|H\right\|_{2}^{2}\sum_{k=2}^{\infty}\frac{k-1}{2}\lambda^{k-1}=\frac{\beta}{2}\left\|H\right\|_{2}^{2}\frac{\lambda}{\left(1-\lambda\right)^{2}}$ (2.39)
using relation (A.42). Next consider all terms with $j\geq 3$, i.e. with higher order than $2$ in $H$. The general form of such a term, with $k\geq j$, is
$\mathrm{Tr}\;M_{1}\;T_{k,j,l}=\mathrm{Tr}\;\left(I+Q_{1}\right)R_{1}\left(I-R_{1}\right)^{a}H\Pi_{k,j,l}H\;$
where $a$ depends on $k$ and $l$, with $a\leq k-j$, $a\geq 0$ and $\Pi_{k,j,l}$ is a matrix monomial containing $b$ factors $I-R_{1}$ and j2j-2 factors $H$ (recall that $I-R_{1}$ and $H$ do not commute). Here $b$ depends on $k$ and $l$ and fulfills a+b=kja+b=k-j with $a,b\geq 0$. Again we estimate, analogously to (2.37),
$\left|\mathrm{Tr}\;M_{1}\;T_{k,j,l}\right|$ $\leq\beta\lambda^{a+1}\left\|H\right\|_{2}\left\|\Pi_{k,j,l}H\right\|_{2}$
$\leq\beta\lambda^{a+1}\left|\Pi_{k,j,l}\right|\left\|H\right\|_{2}^{2}.$
Successive application of the inequality (2.34) gives
$\left|\Pi_{k,j,l}\right|\leq\lambda^{b}\left|H\right|^{j-2}.$ (2.40)
As an illustration consider the simple case $\Pi_{k,j,l}=\left(I-R_{1}\right)H\left(I-R_{1}\right)$ where $k=6,j=3,a=1,b=2$ . Then
$\left|\Pi_{k,j,l}\right|$ $=\left|\left(I-R_{1}\right)H\left(I-R_{1}\right)\right|\leq\lambda\left|H\left(I-R_{1}\right)\right|$
$\leq\lambda\left|H\right|\left|\left(I-R_{1}\right)\right|\leq\lambda^{2}\left|H\right|.$
Since (2.40) holds generally, applying the bound $\left|H\right|\leq\left\|H\right\|_{2}$ we obtain for $j\geq 3$
$\left|\mathrm{Tr}\;M_{1}\;T_{k,j,l}\right|\leq\beta\lambda^{k-j+1}\left\|H\right\|_{2}^{j}.$ (2.41)
From (2.41) we obtain for the totality of terms of third order or higher
[…]
To show that the expression in $\left\{\cdot\right\}$ is finite, set h:=k2h:=k-2, m:=j2m:=j-2. Then the expression in $\left\{\cdot\right\}$ is
$\sum_{h=1}^{\infty}\sum_{m=1}^{h}\frac{\left(h+1\right)!}{\left(m+2\right)!\left(h-m\right)!}\lambda^{h-m+1}\left\|H\right\|_{2}^{m}$
[…]
$\leq\frac{\lambda}{4}\sum_{h=1}^{\infty}\left(h+1\right)\left(\lambda+\left\|H\right\|_{2}\right)^{h}.$
Denote $\gamma=\lambda+\left\|H\right\|_{2}$ and note that $\gamma<1$ due to $\left\|H\right\|_{2}<1-\lambda$. Using relation (A.42) again, we find
$\sum_{h=1}^{\infty}\left(h+1\right)\gamma^{h}=\frac{\gamma}{\left(1-\gamma\right)^{2}}+\frac{\gamma}{1-\gamma}\leq\frac{2}{\left(1-\gamma\right)^{2}}.$
From (2.42) we find for the totality of terms of third order or higher
[…]
The argument for the terms involving $U_{k,j,l}$ is analogous, which proves the lemma.
Lemma 2.3
For given $\lambda$ from (2.24) there exists $\delta>0$, not depending on dimension $n$, such that $\left\|H\right\|<\delta$ implies
$\mathrm{Tr}\;\left(I+Q_{1}\right)\;S_{2}\left(R_{1}||R_{2}\right)\leq\delta^{-1}\left\|H\right\|^{2}.$
Here $\delta$ can be chosen as
$\delta=\min\left(\left(1-\lambda\right)/2,\left(1-\lambda\right)^{3}/8\lambda\right).$
Proof. In the series (2.35) we can now rearrange terms; consider the series given by all linear (in $H$) terms. This is found as
$\sum_{k=1}^{\infty}\sum_{l=1}^{k}\frac{1}{k}\mathrm{Tr}\;M_{1}\;T_{k,1,l}+\sum_{k=1}^{\infty}\sum_{l=1}^{k}\frac{1}{k}\mathrm{Tr}\;M_{2}\;U_{k,1,l}$
$=\sum_{k=1}^{\infty}\mathrm{Tr}\;\left(I+Q_{1}\right)R_{1}\left(I-R_{1}\right)^{k-1}H+\sum_{k=1}^{\infty}\mathrm{Tr}\;\left(I+Q_{1}\right)\left(I-R_{1}\right)R_{1}^{k-1}G$
$=\mathrm{Tr}\;H\left(I+Q_{1}\right)R_{1}\left(\sum_{k=1}^{\infty}\left(I-R_{1}\right)^{k-1}\right)+\sum_{k=1}^{\infty}\mathrm{Tr}\;G\left(I+Q_{1}\right)\left(I-R_{1}\right)\left(\sum_{k=1}^{\infty}R_{1}^{k-1}\right).$ (2.45)
We note that
$\sum_{k=1}^{\infty}\left(I-R_{1}\right)^{k-1}$ $=\sum_{k=0}^{\infty}\left(I-R_{1}\right)^{k}$
$=\left(I-\left(I-R_{1}\right)\right)^{-1}=R_{1}^{-1},$
$\sum_{k=1}^{\infty}R_{1}^{k-1}$ $=\left(I-R_{1}\right)^{-1}.$
Thus, in view of G=HG=-H, (2.45) equals
$\mathrm{Tr}\;H\left(I+Q_{1}\right)-\mathrm{Tr}\;H\left(I+Q_{1}\right)=0.$
In the series (2.35) there now remain only terms of quadratic and higher order in $H$. By (2.39), (2.44) and the analogous bounds for terms involving $U_{k,j,l}$ with $j\geq 2$, recalling $\left\|H\right\|=\left\|G\right\|$, we have
$\mathrm{Tr}\;\left(I+Q_{1}\right)\;S_{2}\left(R_{1}||R_{2}\right)\leq$
$\beta\left\|H\right\|^{2}\frac{\lambda}{\left(1-\lambda\right)^{2}}+\frac{\beta\lambda\left\|H\right\|^{2}}{\left(1-\lambda-\left\|H\right\|\right)^{2}}\leq\frac{2\beta\lambda\left\|H\right\|^{2}}{\left(1-\lambda-\left\|H\right\|\right)^{2}}.$
Set $\delta_{0}:=\left(1-\lambda\right)/2$; then for $\left\|H\right\|<\delta_{0}$
$\mathrm{Tr}\;\left(I+Q_{1}\right)\;S_{2}\left(R_{1}||R_{2}\right)\leq\frac{8\beta\lambda}{\left(1-\lambda\right)^{2}}\left\|H\right\|^{2}$
Now with $\delta:=\min\left(\delta_{0},\left(1-\lambda\right)^{2}/8\beta\lambda\right)$ and $\beta=\left(1-\lambda\right)^{-1}$ we obtain the assertion.
Having bounded the relative entropy $S\left(\rho_{1}||\rho_{2}\right)$ from (2.17) in terms of the difference $H=R_{1}-R_{2}$, in the next step we have to estimate $H$ in terms of the difference $A_{1}-A_{2}$. Recalling 2.11, we have
$H=\frac{Q_{1}}{Q_{1}+I}-\frac{Q_{2}}{Q_{2}+I}.$
We will estimate $H$ in terms of $D:=Q_{2}-Q_{1}$, and in view of the relation $Q:=\left(A-I\right)/2$, we have $D=\left(A_{2}-A_{1}\right)/2$.
Lemma 2.4
Under condition (2.24) we have
$\left\|H\right\|^{2}=\left\|R_{1}-R_{2}\right\|^{2}\leq\frac{1}{\left(1-\lambda\right)^{2}}\left\|A_{1}-A_{2}\right\|^{2}.$
Proof. We have
$\left\|H\right\|$ $=\left\|\frac{Q_{1}}{Q_{1}+I}-\left(Q_{2}+I\right)^{-1}\left(Q_{1}+D\right)\right\|$
$\leq\left\|\left(\left(Q_{1}+I\right)^{-1}-\left(Q_{2}+I\right)^{-1}\right)Q_{1}\right\|+\left\|\left(Q_{2}+I\right)^{-1}D\right\|$ (2.46)
The first term in (2.46) equals
$\left\|\left(Q_{1}+I\right)^{-1}\left(\left(Q_{2}+I\right)-\left(Q_{1}+I\right)\right)\left(Q_{2}+I\right)^{-1}Q_{1}\right\|$
$=\left\|\left(Q_{1}+I\right)^{-1}D\left(Q_{2}+I\right)^{-1}Q_{1}\right\|$
$\leq\left|\left(Q_{1}+I\right)^{-1}\right|\left\|D\left(Q_{2}+I\right)^{-1}Q_{1}\right\|.$
Here $Q_{i}>0$ so that $\left(Q_{1}+I\right)^{-1}<I,$ and the above is bounded by
$\left\|D\left(Q_{2}+I\right)^{-1}Q_{1}\right\|$ $=\left\|Q_{1}\left(Q_{2}+I\right)^{-1}D\right\|$
$\leq\left|Q_{1}\right|\cdot\left|\left(Q_{2}+I\right)^{-1}\right|\cdot\left\|D\right\|$
$\leq\frac{\lambda}{1-\lambda}\left\|D\right\|$ (2.47)
in view of (2.25). The second term in (2.46) can be bounded by $\left\|D\right\|$. In conjunction with (2.47) this gives
$\left\|H\right\|^{2}$ $\leq\left(\frac{\lambda}{1-\lambda}+1\right)^{2}\left\|D\right\|^{2}$
$=\frac{1}{\left(1-\lambda\right)^{2}}\left\|A_{1}-A_{2}\right\|^{2}.$
We can can summarize the results of this subsection as follows.
Proposition 2.5
Let $A_{i}$, $i=1,2$ be Hermitian $n\times n$ symbols fulfilling for some $\mu\in\left(0,1\right)$
$\left(1+\mu\right)I\leq A_{i}\leq\mu^{-1}I\text{, }i=1,2.$
Let the Gaussian states $\rho_{i}$, $i=1,2$ be defined as in Proposition 2.1, and let $S\left(\rho_{1}||\rho_{2}\right)$ be the relative entropy. Then there exists $\delta>0$, depending on $\mu$ but not on $n$, such that $\left\|A_{1}-A_{2}\right\|<\delta$ implies
$S\left(\rho_{1}||\rho_{2}\right)\leq\delta^{-1}\left\|A_{1}-A_{2}\right\|^{2}.$
Proof. Given $\mu\in\left(0,1\right)$, we can find $\lambda\in\left(1/2,1\right)$ such that
$\frac{2}{\lambda}-1\leq\mu<\mu^{-1}\leq\frac{1+\lambda}{1-\lambda}.$
Then (2.26) and hence (2.24) is fulfilled. The previous two lemmas then prove the claim.
2.3 Approximation of Toeplitz matrices
We follow [Nik20], 5.5 to collect some basic facts about Toeplitz and circulant matrices. Assume $m$ is an odd numer, let $\mathbf{c}=\mathbf{c}_{0}=\left(c_{0},\ldots,c_{m-1}\right)^{\prime}$ be a column vector of complex elements, let $\mathbf{c}_{1}=\left(c_{m-1},c_{0},\ldots,c_{m-2}\right)^{\prime}$ be a cyclic shift, and let $\mathbf{c}_{k}$ be the $k$-th cyclic shift such that $\mathbf{c}_{m}=\left(c_{1},c_{2},\ldots,c_{m-1},c_{0}\right)^{\prime}$. Then the $m\times m$ circulant pertaining to $\mathbf{c}$ is
$T_{m}=\left(\begin{array}[c]{ccc}\mathbf{c}_{0}&\ldots&\mathbf{c}_{m-1}\end{array}\right).$ (2.48)
Then $\mathbf{c}$ is the representing vector and the representing polynomial is $p\left(z\right)=\sum_{k=0}^{m-1}c_{k}z^{k}$; we write $T_{m}=T_{m}\left(\mathbf{c}\right)=T_{m}\left(p\right)$. Clearly $T_{m}\left(\mathbf{c}\right)$ is a Toeplitz matrix. To describe the spectral properties, define
$\epsilon_{k}=\exp\left(2\pi ik/m\right)\text{, }\mathbf{u}_{k}:=\left(1,\epsilon_{k},\epsilon_{k}^{2},\ldots,\epsilon_{k}^{m-1}\right)^{\prime}m^{-1/2}\text{, }k\in\mathbb{Z}$ (2.49)
and the discrete Fourier transform $\mathcal{F}_{d,m}:\mathbb{C}^{m}\rightarrow$ $\mathbb{C}^{m}$ by its matrix
$\mathcal{F}_{d,m}=\left(\mathbf{u}_{0},\ldots,\mathbf{u}_{m-1}\right).$ (2.50)
Then $\mathcal{F}_{d,m}$ is unitary, and diagonalizes every circulant $T_{m}\left(p\right)$ in the sense that
$\mathcal{F}_{d,m}^{\ast}T_{m}\left(p\right)\mathcal{F}_{d,m}=\mathrm{diag}\left(p\left(1\right),p\left(\bar{\epsilon}_{1}\right),\ldots,p\left(\bar{\epsilon}_{m-1}\right)\right)$ (2.51)
(cf. [Nik20], 5.5.4).
We give an alternative description of the spectral properties as follows. Define
$\phi_{k}\left(\omega\right)$ $=\exp\left(ik\omega\right)\text{, }\omega\in\mathbb{R}\text{, }k\in\mathbb{Z},$ (2.52)
$\omega_{j,m}$ $=\frac{2\pi j}{m}\text{, }j\in\mathbb{Z}.$ (2.53)
Lemma 2.6
Assume that $m$ is odd; define $c_{-k}=c_{m-k}$, $k=1,\ldots,\left(m-1\right)/2$ and a function
$g_{m}\left(\mathbf{c},\omega\right)=\sum_{k=-\left(m-1\right)/2}^{\left(m-1\right)/2}c_{k}\phi_{-k}\left(\omega\right)\text{, }\omega\in\mathbb{R}\text{.}$
Then (2.51) can be written
$\mathcal{F}_{d,m}^{\ast}T_{m}\left(\mathbf{c}\right)\mathcal{F}_{d,m}=\mathrm{diag}\left(g_{m}\left(\mathbf{c},\omega_{0,m}\right),\ldots,g_{m}\left(\mathbf{c},\omega_{m-1,m}\right)\right).$ (2.54)
Furthermore define a unitary $m\times m$ matrix, with $\mathbf{u}_{k}$ from (2.49)
$U_{m}=\left(\mathbf{u}_{-(m-1)/2},\ldots,\mathbf{u}_{0},\ldots\mathbf{u}_{(m-1)/2}\right).$ (2.55)
Then (2.54) is equivalent to
$U_{m}^{\ast}T_{m}\left(\mathbf{c}\right)U_{m}=\mathrm{diag}\left(g_{m}\left(\mathbf{c},\omega_{-\left(m-1\right)/2,m}\right),\ldots,g_{m}\left(\mathbf{c},\omega_{\left(m-1\right)/2,m}\right)\right).$ (2.56)
Proof. First note that the eigenvalues $p\left(\bar{\epsilon}_{j}\right)$ in (2.51) can be written as
$p\left(\bar{\epsilon}_{j}\right)$ $=\sum_{k=0}^{m-1}c_{k}\bar{\epsilon}_{j}^{k}=\sum_{k=0}^{m-1}c_{k}\exp\left(-2\pi ijk/m\right)$
$=\sum_{k=0}^{m-1}c_{k}\phi_{-k}\left(\omega_{j.m}\right).\text{ }j=0,\ldots,m-1.$
By periodicity we have
$\phi_{-\left(m-k\right)}\left(\omega_{j,m}\right)$ $=\exp\left(-i\left(m-k\right)\omega_{j,m}\right)$
$=\exp\left(ik\omega_{j,m}\right)\exp\left(-im\frac{2\pi j}{m}\right)=\phi_{k}\left(\omega_{j,m}\right)$
for $k=1,\ldots,m-1$. Hence
$p\left(\bar{\epsilon}_{j}\right)$ $=\sum_{k=0}^{\left(m-1\right)/2}c_{k}\phi_{-k}\left(\omega_{j,m}\right)+\sum_{k=\left(m+1\right)/2}^{m-1}c_{k}\phi_{-k}\left(\omega_{j,m}\right)$
$=\sum_{k=0}^{\left(m-1\right)/2}c_{k}\phi_{-k}\left(\omega_{j,m}\right)+\sum_{k=1}^{\left(m-1\right)/2}c_{m-k}\phi_{-\left(m-k\right)}\left(\omega_{j,m}\right)$
$=\sum_{k=0}^{\left(m-1\right)/2}c_{k}\phi_{-k}\left(\omega_{j,m}\right)+\sum_{k=1}^{\left(m-1\right)/2}c_{-k}\phi_{k}\left(\omega_{j,m}\right)=g_{m}\left(\mathbf{c},\omega_{j,m}\right)\text{, }j=0,\ldots,m-1$
which implies (2.54). This relation is equivalent to
$T_{m}\left(\mathbf{c}\right)$ $=\mathcal{F}_{d,m}\mathrm{diag}\left(g_{m}\left(\mathbf{c},\omega_{0,m}\right),\ldots,g_{m}\left(\mathbf{c},\omega_{m-1,m}\right)\right)\mathcal{F}_{d,m}^{\ast}$
$=\sum_{k=0}^{m-1}\mathbf{u}_{k}\mathbf{u}_{k}^{\ast}g_{m}\left(\mathbf{c},\omega_{k,m}\right).$
By periodicity of the function $g_{m}$ in $\omega$ we have $g_{m}\left(\mathbf{c},\omega_{m-k,m}\right)=g_{m}\left(\mathbf{c},\omega_{-k,m}\right)$, $k=1,\ldots,\left(m-1\right)/2$, and we also have $\epsilon_{m-k}=\epsilon_{-k}$ and hence $\mathbf{u}_{m-k}=\mathbf{u}_{-k}$. Thus we obtain
$T_{m}\left(\mathbf{c}\right)$ $=\sum_{k=0}^{m-1}\mathbf{u}_{k}\mathbf{u}_{k}^{\ast}g_{m}\left(\mathbf{c},\omega_{k,m}\right)=\sum_{k=0}^{\left(m-1\right)/2}\mathbf{u}_{k}\mathbf{u}_{k}^{\ast}g_{m}\left(\mathbf{c},\omega_{k,m}\right)+\sum_{k=1}^{\left(m-1\right)/2}\mathbf{u}_{m-k}\mathbf{u}_{m-k}^{\ast}g_{m}\left(\mathbf{c},\omega_{m-k,m}\right)$
$=\sum_{k=-\left(m-1\right)/2}^{\left(m-1\right)/2}\mathbf{u}_{k}\mathbf{u}_{k}^{\ast}g_{m}\left(\mathbf{c},\omega_{k,m}\right),$
implying (2.56)
Note that the matrix $U_{m}$ is a permutation of $\mathcal{F}_{d,m}$, thus it can be considered a version of the discrete Fourier transform.
Since we will use the circulants to approximate the Hermitian Toeplitz symbol matrices $A_{m}\left(a\right)$, we will also assume that $T_{m}\left(\mathbf{c}\right)$ is Hermitian. From (2.48) it can be seen that in terms of $\mathbf{c}$ this means
$c_{0}=\bar{c}_{0}\ \text{and }\bar{c}_{k}=c_{m-k},\ k=1,\ldots,m-1.$ (2.57)
Then $c_{-k}=\bar{c}_{k}$ and consequently the function $g_{m}\left(\mathbf{c},\omega\right)$ is real, thus also the eigenvalues of $T_{m}\left(\mathbf{c}\right)$ are real.
For a symbol matrix $A_{m}\left(a\right)$ defined by (1.21) pertaining to spectral density $a$ we will define a circular approximant $\tilde{A}_{m}\left(a\right)$ as
$\tilde{A}_{m}\left(a\right):=T_{m}\left(\mathbf{c}\right)$ (2.58)
for a representing vector
$\mathbf{c}=\left(a_{0},a_{-1},\ldots,a_{-\left(m-1\right)/2},a_{\left(m-1\right)/2},\ldots,a_{1}\right)^{\prime}.$ (2.59)
In view of $\bar{a}_{k}=a_{-k}$ it can be checked that (2.57) is fulfilled and thus $\tilde{A}_{m}\left(a\right)$ is Hermitian. One then checks that $g_{m}\left(\mathbf{c},\omega\right)$ takes the form
$g_{m}\left(\mathbf{c},\omega\right)$ $=\tilde{a}_{m}\left(\omega\right)\text{ where}$
$\tilde{a}_{m}\left(\omega\right)$ $=\sum_{k=-\left(m-1\right)/2}^{\left(m-1\right)/2}a_{k}\phi_{k}\left(\omega\right)\text{, }\omega\in\mathbb{R}\text{.}$ (2.60)
According to (2.56), the eigenvalues of $\tilde{A}_{m}\left(a\right)$ are then $\tilde{a}_{m}\left(\omega_{j,m}\right)$, $j=-\left(m-1\right)/2,\ldots,\left(m-1\right)/2$. Now $\tilde{a}_{m}$ is a Fourier series approximation to $a$; indeed it follows from (1.21), if $a$ is square integrable on $\left(-\pi,\pi\right)$, that
$a\left(\omega\right)=\sum_{k=-\infty}^{\infty}a_{k}\phi_{k}\left(\omega\right)\text{, }\omega\in\mathbb{R}\text{.}$
For later reference we state the following simple approximation result.
Lemma 2.7
Assume $a\in\Theta_{1}\left(\alpha,M\right)$ for $\alpha>1/2$. Then as $m\rightarrow\infty$
$\sup_{\omega\in\left(-\pi.\pi\right)}\left|a\left(\omega\right)-\tilde{a}_{m}\left(\omega\right)\right|=o\left(1\right)$
Proof. We have
$\sup_{\omega\in\left(-\pi.\pi\right)}\left|a\left(\omega\right)-\tilde{a}_{m}\left(\omega\right)\right|^{2}=\sup_{\omega\in\left(-\pi.\pi\right)}\left|\sum_{\left|k\right|>\left(m-1\right)/2}a_{k}\phi_{k}\left(\omega\right)\right|^{2}$
$\leq\left(\sum_{\left|k\right|>\left(m-1\right)/2}\left|a_{k}\right|^{2}k^{2\alpha}\right)\left(\sum_{\left|k\right|>\left(m-1\right)/2}k^{-2\alpha}\right)\leq M\;C_{\alpha}\left(m-3\right)^{1-2\alpha}=o\left(1\right)$
as $m\rightarrow\infty$, where the constant $C_{\alpha}$ depends only on $\alpha$.
We summarize the above facts about circulants as follows.
Lemma 2.8
For a real valued function $a\in L^{2}\left(-\pi,\pi\right)$ and odd $m$, consider the $m\times m$ circulant matrix $\tilde{A}_{m}\left(a\right)$ given by (2.58), (2.59). Also define a diagonal matrix
$\tilde{\Lambda}_{m}\left(a\right)=\mathrm{diag}\left(\tilde{a}_{m}\left(\omega_{-\left(m-1\right)/2,m}\right),\ldots,\tilde{a}_{m}\left(\omega_{\left(m-1\right)/2,m}\right)\right)$ (2.61)
where $\tilde{a}_{m}$ is the Fourier series approximation to $a$ given in (2.60) and $\omega_{j,m}=2\pi j/m$, $j\in\mathbb{Z}$. Then, for the unitary $U_{m}$ defined in (2.55) we have
$U_{m}^{\ast}\tilde{A}_{m}\left(a\right)U_{m}=\tilde{\Lambda}_{m}\left(a\right).$ (2.62)
Recall that the quantum statistical experiment considered in Theorem 1.1 is
$\mathcal{E}_{n}\left(\Theta_{1}\left(\alpha,M\right)\right):=\left\{\mathfrak{N}_{n}\left(0,A_{n}\left(a\right)\right),a\in\Theta_{1}\left(\alpha,M\right)\right\}.$
For some odd $m>n$, let $\tilde{A}_{m}\left(a\right)$ be the circulant approximation (2.58) to $A_{m}\left(a\right)$, and consider the $m$-mode state (or quantum time series) $\mathfrak{N}_{m}\left(0,\tilde{A}_{m}\left(a\right)\right)$ given by symbol $\tilde{A}_{m}\left(a\right)$. Furthermore consider the subsystem of the latter given by symbol $\tilde{A}_{n,m}\left(a\right)$, where $\tilde{A}_{n,m}\left(a\right)$ is the upper left $n\times n$ submatrix of $\tilde{A}_{m}\left(a\right)$. The following result on approximation of symbols in Hilbert-Schmidt norm $\left\|A\right\|_{2}=\left(\mathrm{Tr}\;A^{\ast}A\right)^{1/2}$ is key for an approximation of the corresponding states.
Lemma 2.9
Assume $m$ is odd, $n<m<2\left(n-1\right)$. Then for $a\in W^{\alpha}(M)$, $\alpha>1/2$ (cp. (2.7)) we have
$\left\|A_{n}\left(a\right)-\tilde{A}_{n,m}\left(a\right)\right\|_{2}^{2}\leq 4\left(m-n+1\right)^{1-2\alpha}M.$
Proof. The restriction on $m$ implies $\left(m+1\right)/2\leq n-1$. From the definitions of $A_{n}\left(a\right)$ and $\tilde{A}_{n,m}\left(a\right)$ we immediately obtain
$\left\|A_{n}\left(a\right)-\tilde{A}_{n,m}\left(a\right)\right\|_{2}^{2}$
$=2\sum_{k=\left(m+1\right)/2}^{n-1}\left(n-k\right)\left|a_{k}-\bar{a}_{m-k}\right|^{2}$ (2.63)
$\leq 4\sum_{k=\left(m+1\right)/2}^{n-1}(n-k)\left(\left|a_{k}\right|^{2}+\left|a_{m-k}\right|^{2}\right).$ (2.64)
Note that for $m>n$, the relation $\left(m+1\right)/2\leq k\leq n-1$ implies $k>\left(n+1\right)/2$ and therefore nk<kn-k<k, and note also nk<mkn-k<m-k. We obtain an upper bound for (2.64)
$4\sum_{k=\left(m+1\right)/2}^{n-1}k\left|a_{k}\right|^{2}+4\sum_{k=\left(m+1\right)/2}^{n-1}(m-k)\left|a_{m-k}\right|^{2}$
$=4\sum_{k=\left(m+1\right)/2}^{n-1}k\left|a_{k}\right|^{2}+4\sum_{k=m-n+1}^{\left(m-1\right)/2}k\left|a_{k}\right|^{2}=4\sum_{k=m-n+1}^{n-1}k\left|a_{k}\right|^{2}$
$\leq 4(m-n+1)^{1-2\alpha}\sum_{k=m-n+1}^{n-1}k^{2\alpha}\left|a_{k}\right|^{2}\leq 4(m-n+1)^{1-2\alpha}\left|a\right|_{2,\alpha}^{2}$
where $\left|\cdot\right|_{2,\alpha}^{2}$ is defined in (2.6), and $\alpha>1/2$. Now $\left|a\right|_{2,\alpha}^{2}\leq M$ for $a\in W^{\alpha}(M)$ proves the claim.
2.4 Upper information bound via approximation of symbols
To apply Lemma 2.9 on approximation of symbols $A$ to the corresponding states $\mathfrak{N}_{n}\left(0,A\right)$ via Proposition 2.5, we need uniform bounds on the eigenvalues of the symbols involved.
Lemma 2.10
Suppose $a\in\Theta_{1}\left(\alpha,M\right)$ for $\alpha>1/2$, $M>1$. Then there exists $C=C_{M,\alpha}>1$ such that for $n\geq 1$
$\left(1+C^{-1}\right)\;I\leq A_{n}\left(a\right)\leq C\;I.$ (2.65)
Furthermore, there exist $C_{1}>1$ and $m_{0}$ such that for odd $m\geq m_{0}$ and all $n<m$
$\left(1+C_{1}\right)\;I\leq\tilde{A}_{n,m}\left(a\right)\leq C_{1}\;I\text{.}$ (2.66)
Proof. Consider $x\in\mathbb{C}^{n}$ with $\left\|x\right\|=1$; then in view of (1.21)
$\left\langle x,A_{n}\left(a\right)x\right\rangle=\frac{1}{2\pi}\int_{-\pi}^{\pi}\left|\sum_{j=1}^{n}x_{j}\exp\left(ij\omega\right)\right|^{2}a\left(\omega\right)d\omega.$
Applying the second inequality in (2.10) we obtain
$\left\langle x,A_{n}\left(a\right)x\right\rangle$ $\leq\frac{\mu^{-1}}{2\pi}\int_{-\pi}^{\pi}\left|\sum_{j=1}^{n}x_{j}\exp\left(ij\omega\right)\right|^{2}d\omega$
$=\mu^{-1}\sum_{j=1}^{n}\left|x_{j}\right|^{2}=\mu^{-1}.$
Analogously we obtain from the first inequality in (2.10) $\left\langle x,A_{n}\left(a\right)x\right\rangle\geq\left(1+\mu\right)$, so that (2.65) is shown. To establish (2.66), note first that since $\tilde{A}_{n,m}\left(a\right)$ is a central submatrix of $\tilde{A}_{m}\left(a\right)$, we have
$\lambda_{\min}\left(\tilde{A}_{m}\left(a\right)\right)\leq\lambda_{\min}\left(\tilde{A}_{n,m}\left(a\right)\right)\text{, }\lambda_{\max}\left(\tilde{A}_{n,m}\left(a\right)\right)\leq\lambda_{\max}\left(\tilde{A}_{m}\left(a\right)\right)$
so we need to deal only with $\tilde{A}_{m}\left(a\right)$. Lemma 2.8 describes the eigenvalues of this matrix as certain function values $\tilde{a}_{m}\left(\omega_{j,m}\right)$. Now according to Lemma 2.7 $\tilde{a}_{m}$ approximates $a$ uniformly if $a\in\Theta_{1}\left(\alpha,M\right)$ for $\alpha>1/2$. In conjunction with (2.10) this proves the second claim.
In this section and the next, the parameter space for the quantum statistical experiments to be considered will always be the set $\Theta_{1}\left(\alpha,M\right)$ considered in Theorem 1.1, and will often be omitted from notation.
Proposition 2.11
Consider the experiment $\mathcal{E}_{n}=\mathcal{E}_{n}\left(\Theta_{1}\left(\alpha,M\right)\right)$ defined in (1.28) and define also for odd $m$
[…]
where $\tilde{A}_{m}\left(a\right)$ is the circulant matrix defined in (2.58) such that $n<m<2\left(n-1\right)$. Assume $m$ is chosen such that $m-n\rightarrow\infty$; then
[…]
Proof. Consider the submatrix $\tilde{A}_{n,m}\left(a\right)$ of $\tilde{A}_{m}\left(a\right)$ occurring in Lemma 2.9. If $m-n\rightarrow\infty$ then Lemma 2.9 in conjunction with Proposition 2.5 and Lemma 2.10 implies existence of a constant $\delta>0$ such that for the relative entropy $S\left(\cdot||\cdot\right)$
$S\left(\mathfrak{N}_{n}\left(0,A_{n}\left(a\right)\right)||\mathfrak{N}_{n}\left(0,\tilde{A}_{n,m}\left(a\right)\right)\right)$ $\leq\delta^{-1}\left\|A_{n}\left(a\right)-\tilde{A}_{n,m}\left(a\right)\right\|^{2}$
$\leq\delta^{-1}4\left(m-n+1\right)^{1-2\alpha}M=o\left(1\right)$
since $\alpha>1/2$. By inequality (2.15) we then also have
$\sup_{a\in\Theta_{1}\left(\alpha,M\right)}\left\|\mathfrak{N}_{n}\left(0,A_{n}\left(a\right)\right)-\mathfrak{N}_{n}\left(0,\tilde{A}_{n,m}\left(a\right)\right)\right\|_{1}^{2}\rightarrow 0.$ (2.67)
Obviously there is a quantum channel which maps the $m$-mode state $\mathfrak{N}_{m}\left(0,\tilde{A}_{m}\left(a\right)\right)$ into the $n$-mode state $\mathfrak{N}_{n}\left(0,\tilde{A}_{n,m}\left(a\right)\right)$, as the quantum equivalent of ”omitting observations”, i.e. the partial trace. Formally this channel $\alpha$ is described in terms of a map between the respective algebras in the Appendix, Subsection A.2.1; we then have
$\mathfrak{N}_{m}\left(0,\tilde{A}_{m}\left(a\right)\right)\circ\alpha=\mathfrak{N}_{n}\left(0,\tilde{A}_{n,m}\left(a\right)\right).$
From (2.67) we then obtain
$\sup_{a\in\Theta_{1}\left(\alpha,M\right)}\left\|\mathfrak{N}_{m}\left(0,\tilde{A}_{m}\left(a\right)\right)\circ\alpha-\mathfrak{N}_{n}\left(0,A_{n}\left(a\right)\right)\right\|_{1}^{2}\rightarrow 0$
which implies the claim.
2.5 The geometric regression model
The spectral decomposition of the circulant matrix $\tilde{A}_{m}\left(a\right)$ is described in Lemma 2.8. Since
$\mathfrak{N}_{m}\left(0,\tilde{A}_{m}\left(a\right)\right)=\frac{2^{m}}{\det\left(\tilde{A}_{m}\left(a\right)+I\right)}\left(\frac{\tilde{A}_{m}\left(a\right)-I}{\tilde{A}_{m}\left(a\right)+I}\right)_{F}$
by (2.5), we can use the property of Fock operators (2.4) to diagonalize the state. For the diagonal symbol matrix $\tilde{\Lambda}_{m}\left(a\right)$ defined in (2.61), consider an experiment
[…]
Lemma 2.12
For all odd $m\geq 3$, we have statistical equivalence
[…]
Proof. From (2.62) and (2.4) it follows that
$\left(U_{m}^{\ast}\right)_{F}\mathfrak{N}_{m}\left(0,\tilde{A}_{m}\left(a\right)\right)\left(U_{m}\right)_{F}$ $=\frac{2^{m}}{\det\left(\tilde{\Lambda}_{m}\left(a\right)+I\right)}\left(\frac{\tilde{\Lambda}_{m}\left(a\right)-I}{\tilde{\Lambda}_{m}\left(a\right)+I}\right)_{F}$
$=\mathfrak{N}_{m}\left(0,\tilde{\Lambda}_{m}\left(a\right)\right).$
Since $\left(U_{m}\right)_{F}$ is unitary, the above mapping of $\mathfrak{N}_{m}\left(0,\tilde{A}_{m}\left(a\right)\right)$ to $\mathfrak{N}_{m}\left(0,\tilde{\Lambda}_{m}\left(a\right)\right)$ represents an invertible state transition (or dual channel, cf. Subsection A.1). This implies the equivalence claim by definition of $\Delta\left(\cdot,\cdot\right)$.
[…]
$\lambda_{j,m}\left(a\right)=\tilde{a}_{m}\left(\omega_{j-\left(m+1\right)/2,m}\right)\text{, }j=1,\ldots,m$
and define (for odd $m$) a set of probability measures (products of geometric distributions)
$\mathcal{\tilde{G}}_{m}=\left\{{\displaystyle\bigotimes\limits_{J=1}^{m}}\mathrm{Geo}\left(p\left(\lambda_{j,m}\left(a\right)\right)\right),a\in\Theta_{1}\left(\alpha,M\right)\right\}.$ (2.68)
where $p\left(x\right)=\left(x-1\right)/\left(x+1\right)$.
Proposition 2.13
For all odd $m\geq 3$, we have statistical equivalence
[…]
Proof. Consider the covariance matrix of $\mathfrak{N}_{m}\left(0,\Lambda_{m}\left(a\right)\right)$, which according to (1.16) is
$\Sigma=\frac{1}{2}\left(\begin{array}[c]{cc}\Lambda_{m}\left(a\right)&0\\ 0&\Lambda_{m}\left(a\right)\end{array}\right).$
This corresponds to a vector of canonical observables $\mathbf{R}=\left(Q_{1},\ldots,Q_{m},P_{1},\ldots,P_{m}\right)$. with a rearrangement as $\mathbf{\check{R}}:=(Q_{1},P_{1},\ldots,Q_{m},P_{m})$ as in (1.14) the covariance matrix becomes block diagonal
$\check{\Sigma}=\frac{1}{2}\left(\begin{array}[c]{ccc}\lambda_{1,m}\left(a\right)I_{2}&&0\\ &\ldots&\\ 0&&\lambda_{m,m}\left(a\right)I_{2}\end{array}\right).$
The centered $m$-mode Gaussian state is clearly the tensor product of $m$ one-mode Gaussian states with covariance matrix $\frac{1}{2}\lambda_{j,m}\left(a\right)I_{2},$ $j=1,\ldots,m$. A centered Gaussian state with covariance matrix $\frac{1}{2}\lambda I_{2}$, $\lambda>1$ has a representation in Fock space $\mathcal{F}(\mathbb{C})$ (according to (2.5))
$\mathfrak{N}_{1}\left(0,\lambda\right)=\frac{2}{\lambda+1}{\displaystyle\bigoplus\limits_{k\geq 0}}\left(\frac{\lambda-1}{\lambda+1}\right)^{k}$
and setting $p\left(\lambda\right)=\left(\lambda-1\right)/\left(\lambda+1\right)$, we obtain
$\mathfrak{N}_{1}\left(0,\lambda\right)=\left(1-p\left(\lambda\right)\right){\displaystyle\bigoplus\limits_{k\geq 0}}p\left(\lambda\right)^{k}.$ (2.69)
which corresponds to a one mode thermal state with covariance matrix $\frac{1}{2}\lambda I_{2}$, $\lambda>1$ [WPGP+12]. We obtain that $\mathcal{E}_{m}^{d}$ is equivalent to
$\left({\displaystyle\bigotimes\limits_{j=1}^{m}}\mathfrak{N}_{1}\left(0,\lambda_{j,m}\left(a\right)\right),a\in\Theta_{1}\left(\alpha,M\right)\right)$
which in turn, by measuring each tensor factor in the coordinate basis, is equivalent to observing $m$ independent r.v.’s $X_{j}$ having geometric distributions (cf. Subsection A.4)
$X_{j}\sim\mathrm{Geo}\left(p\left(\lambda_{j,m}\left(a\right)\right)\right),\;j=1,\ldots,m.$ (2.70)
This establishes the equivalence claimed .
2.6 Comparing geometric regression models
Having obtained an experiment $\mathcal{\tilde{G}}_{m}$ consisting of classical probability measures, further developments will take place in this framework. Consider the Hellinger distance $H\left(P,Q\right)$ between probability measures $P,Q$ on the same sample space, defined as follows: for $\mu=P+Q$, $p=dP/d\mu$, $q=dP/d\mu$,
$H^{2}\left(P,Q\right)=\int\left(p^{1/2}-q^{1/2}\right)^{2}d\mu.$
Note the relationship to $L_{1}$-distance $\left\|P-Q\right\|_{1}$:
$\frac{1}{2}\left\|P-Q\right\|_{1}\leq H\left(P,Q\right)$ (2.71)
([Tsy09], Lemma 2.3). Also, for product measures $\otimes_{j=1}^{n}P_{j}$ and $\otimes_{j=1}^{n}Q_{j}$ we have
$H^{2}\left(\otimes_{j=1}^{n}P_{j},\otimes_{j=1}^{n}Q_{j}\right)\leq 2\sum_{j=1}^{n}H^{2}\left(P_{j},Q_{j}\right).$ (2.72)
as follows from Lemma 2.19 in [Str85].
For general $n$, consider intervals in $\left(-\pi,\pi\right)$ of equal length
$W_{j,n}=2\pi\left(\frac{\left(j-1\right)}{n}-\frac{1}{2},\frac{j}{n}-\frac{1}{2}\right)\text{, }j=1,\ldots,n$ (2.73)
and for any real $f\in L_{2}(-\pi.\pi),$ let $\bar{f}_{n}$ be the $L_{2}$-projection onto the piecewise constant functions, i.e.
$\bar{f}_{n}=\sum_{j=1}^{n}J_{j,n}(f)\mathbf{1}_{W_{j,n}}\text{, where }J_{j,n}(f)=\frac{n}{2\pi}\int_{W_{j,n}}f(x)dx.$ (2.74)
In agreement with (2.60) define the Fourier series approximation to $f$, for odd $n$
$\tilde{f}_{n}\left(\omega\right)$ $=\sum_{k=-\left(n-1\right)/2}^{\left(n-1\right)/2}f_{k}\phi_{k}\left(\omega\right)\text{, }\omega\in[-\pi,\pi]\text{,}$
$f_{k}$ $=\frac{1}{2\pi}\int_{-\pi}^{\pi}\exp\left(-ik\omega\right)f\left(\omega\right)d\omega.$
Recall also the definition of the seminorm $\left|f\right|_{2,\alpha}^{2}$ and the norm $\left\|f\right\|_{2,\alpha}^{2}$ in (2.6).
Lemma 2.14
For $f\in L_{2}(-\pi.\pi)$, assume $\left|f\right|_{2,\alpha}^{2}$ is finite for given $0<\alpha<1$. Then
(i) there is a constant $C_{\alpha}$ such that
$\left\|f-\bar{f}_{n}\right\|_{2}^{2}\leq C_{\alpha}\;n^{-2\alpha}\;\left|f\right|_{2,\alpha}^{2}.$
(ii) Assume that $1/2<\alpha<1$, that $n$ is odd and let $\tilde{\omega}_{j,n}$ be the midpoint of $W_{j,n}$ $j=1,\ldots,n$. Then there is a constant $C_{\alpha}$ such that
$\sum_{j=1}^{n}\left(\tilde{f}_{n}(\tilde{\omega}_{j,n})-J_{j,n}(f)\right)^{2}\leq C_{\alpha}\;n^{1-2\alpha}\;\left\|f\right\|_{2,\alpha}^{2}.$
Proof. (i) A version of the claim for functions $f$ defined on $\left(0,1\right)$ is proved in Lemma 5.3 [GNZ09]; a rescaling to the interval $\left(-\pi,\pi\right)$ yields the present claim. Also, in [GNZ09] the inequality is proved for a seminorm $\left|f\right|_{B_{2,2}^{\alpha}}^{2}$ in place of $\left|f\right|_{2,\alpha}^{2}$, but Lemma 5.5 in [GNZ09] shows that if $\left|f\right|_{2,\alpha}^{2}<\infty$ then $\left|f\right|_{B_{2,2}^{\alpha}}^{2}\leq C_{\alpha}\left|f\right|_{2,\alpha}^{2}$.
(ii) Again, for an interval $\left(0,1\right)$ the claim is proved in Lemma 5.3 of [GNZ09].
Our next task is to compare the geometric regression experiment $\mathcal{\tilde{G}}_{m}$ defined in (2.68) with the basic one of (1.30) involving the local averages $J_{j,n}\left(a\right)$ from (1.29) for $m=n$. We now write the latter as
$\mathcal{G}_{n}=\left\{{\displaystyle\bigotimes\limits_{j=1}^{n}}\mathrm{Geo}\left(p\left(J_{j,n}\left(a\right)\right)\right),a\in\Theta_{1}\left(\alpha,M\right)\right\}.$
Lemma 2.15
We have asymptotic equivalence, along odd $m\rightarrow\infty$
$\mathcal{\tilde{G}}_{m}\approx\mathcal{G}_{m}.$
Proof. In view of inequalities (2.71) and (2.72) it suffices to prove for the Hellinger distance $H\left(\cdot,\cdot\right)$
$\sum_{j=1}^{m}H^{2}\left(\mathrm{Geo}\left(p\left(\lambda_{j,m}\left(a\right)\right)\right),\mathrm{Geo}\left(p\left(J_{j,m}\left(a\right)\right)\right)\right)=o\left(1\right)$
[…]
$H^{2}\left(\mathrm{Geo}\left(p\left(\lambda_{j,m}\left(a\right)\right)\right),\mathrm{Geo}\left(p\left(J_{j,m}\left(a\right)\right)\right)\right)\leq\frac{\left(\lambda_{j,m}\left(a\right)-J_{j,m}\left(a\right)\right)^{2}}{\left(\lambda_{j,m}\left(a\right)-1\right)\left(J_{j,m}\left(a\right)-1\right)}.$ (2.75)
For the numerator on the r.h.s., observe that $a\in\Theta_{1}\left(\alpha,M\right)$ implies $a\left(\omega\right)\geq 1+M^{-1}$, $\omega\in\left[-\pi,\pi\right]$ and hence also $J_{j,m}\left(a\right)-1\geq M^{-1}$, $j=1,\ldots,m$. Furthermore for $\lambda_{j,m}\left(a\right)=\tilde{a}_{m}\left(\omega_{j-\left(m+1\right)/2,m}\right)$ we can use Lemma 2.7 to show that
$\inf_{j=1,\ldots,m}\lambda_{j,m}\left(a\right)-1\geq M^{-1}\left(1+o\left(1\right)\right).$
It follows that
$\sum_{j=1}^{m}H^{2}\left(\mathrm{Geo}\left(p\left(\lambda_{j,m}\left(a\right)\right)\right),\mathrm{Geo}\left(p\left(J_{j,m}\left(a\right)\right)\right)\right)$
$\leq\left(1+o\left(1\right)\right)M^{-2}\sum_{j=1}^{m}\left(\lambda_{j,m}\left(a\right)-J_{j,m}\left(a\right)\right)^{2}.$
Now observe in the setting of Lemma 2.14 (ii), the midpoints $\tilde{\omega}_{j,m}$ of the intervals $W_{j,m}$ coincide with $\omega_{j-\left(m+1\right)/2,m}$, for $j=1,\ldots,m$. Now reference to the latter result establishes the claim.
Our next task is to compare the basic geometric regression models $\mathcal{G}_{n}$ for different sample sizes $n$ and $m$.
Proposition 2.16
If $m=$ $n+r_{n}$, $0\leq r_{n}=o\left(n^{1/2}\right)$ then we have asymptotic equivalence
$\mathcal{G}_{n}\approx\mathcal{G}_{m}\text{ as }n\rightarrow\infty.$
This will follow from Lemmas 2.172.19 below. Abbreviate $\Theta=\Theta_{1}\left(\alpha,M\right)$ and introduce an experiment
$\mathcal{G}_{n,m}=\left\{{\displaystyle\bigotimes\limits_{j=1}^{n}}\mathrm{NB}^{\otimes m}\left(m^{-1},p\left(J_{j,n}\left(a\right)\right)\right),a\in\Theta\right\}$
where $\mathrm{NB}\left(r,p\right)$ denotes the negative binomial distribution (see Subsection A.5) and $\mathrm{NB}^{\otimes m}\left(r,p\right)$ its $m$-fold product.
Lemma 2.17
For any $n,m>0$ we have equivalence
$\mathcal{G}_{n}\sim\mathcal{G}_{n,m}.$ (2.76)
Proof. Consider a parametric model of independent r.v.’s $X_{k}\sim\mathrm{NB}\left(m^{-1},p\right)$, $k=1,\ldots,m$, $p\in\left(0,1\right)$. Then, as argued in connection with (A.76) below, $\sum_{k=1}^{m}X_{k}$ is a sufficient statistic, and $\sum_{j=1}^{n}X_{i}\sim\mathrm{Geo}\left(p\right)$. Consequently
$\left\{\mathrm{Geo}\left(p\right),\;p\in\left(0,1\right)\right\}\sim\left\{\mathrm{NB}^{\otimes m}\left(m^{-1},p\right),\;p\in\left(0,1\right)\right\}.$
This equivalence via sufficiency easily extends to the experiments given by product measures
$\left\{{\displaystyle\bigotimes\limits_{j=1}^{n}}\mathrm{Geo}\left(p_{j}\right),\;\left(p_{1},\ldots,p_{n}\right)\in\left(0,1\right)^{\times n}\right\}\sim\left\{{\displaystyle\bigotimes\limits_{j=1}^{n}}\mathrm{NB}^{\otimes m}\left(m^{-1},p_{j}\right),\;\left(p_{1},\ldots,p_{n}\right)\in\left(0,1\right)^{\times n}\right\}.$
The common parameter space for $\mathcal{G}_{n},\mathcal{G}_{n,m}$ can be construed as subspace of the one above, which implies the claim.
Introduce an intermediate experiment
$\mathcal{G}_{m,n}^{\ast}=\left\{{\displaystyle\bigotimes\limits_{j=1}^{m}}\mathrm{NB}^{\otimes n}\left(m^{-1},p\left(J_{j,m}\left(a\right)\right)\right),a\in\Theta\right\}.$
Lemma 2.18
For $m\geq n$, we have asymptotic total variation equivalence
$\mathcal{G}_{m,n}^{\ast}\simeq\mathcal{G}_{n,m}\text{ as }n\rightarrow\infty.$
Proof. Write the measures in $\mathcal{G}_{n,m}$ as a product of $mn$ components, i.e. as $\otimes_{j=1}^{mn}Q_{1,j}$ where the component measures $Q_{1,j}$ are defined as follows. For every $j=1,\ldots,mn$, let $k(1,j)$ be the unique index $k\in\left\{1,\ldots,n\right\}$ such that there exists $l\in\left\{1,\ldots,m\right\}$ for which j=(k1)m+lj=(k-1)m+l. Then
$Q_{1,j}=\mathrm{NB}\left(m^{-1},p\left(J_{k(1,j),n}\left(a\right)\right)\right),\;j=1,\ldots,mn.\text{ }$
Analogously, let $k(2,j)$ be the unique index $k\in\left\{1,\ldots,m\right\}$ such that there exists $l\in\left\{1,\ldots,n\right\}$ for which j=(k1)n+lj=(k-1)n+l. Then the measures in $\mathcal{G}_{m,n}^{\ast}$ can be written $\otimes_{j=1}^{mn}Q_{2,j}$ where
$Q_{2,j}=\mathrm{NB}\left(m^{-1},p\left(J_{k(2,j),m}\left(a\right)\right)\right).$
The Hellinger distance between measures in $\mathcal{G}_{n,m}$ and $\mathcal{G}_{m,n}^{\ast}$ is, using (2.72) and then Lemma A.9 (i)
$H^{2}\left({\displaystyle\bigotimes\limits_{j=1}^{mn}}Q_{1,j},{\displaystyle\bigotimes\limits_{j=1}^{mn}}Q_{2,j}\right)\leq 2\sum_{j=1}^{mn}H^{2}\left(Q_{1,j},Q_{2,j}\right)$
$\leq\frac{2}{m}\sum_{j=1}^{mn}\frac{\left(J_{k(1,j),n}\left(a\right)-J_{k(2,j),m}\left(a\right)\right)^{2}}{\left(J_{k(1,j),n}\left(a\right)-1\right)\left(J_{k(2,j),m}\left(a\right)-1\right)}.$ (2.77)
Since $a\in\Theta_{1}\left(\alpha,M\right)$, we have $a\left(\omega\right)\geq 1+M^{-1}$, $\omega\in\left[-\pi,\pi\right]$ and hence also
$\inf_{j=1,\ldots,mn}\min\left(J_{k(1,j),n}\left(a\right),J_{k(2,j),m}\left(a\right)\right)\geq 1+M^{-1}.$
This implies that (2.77) can be bounded by
$\leq\frac{2M^{2}}{m}\sum_{j=1}^{mn}\left(J_{k(1,j),n}\left(a\right)-J_{k(2,j),m}\left(a\right)\right)^{2}.$ (2.78)
The expression $J_{k(1,j),n}(a)-J_{k(2,j),m}(a)$ can be described as follows. For any $x\in\left(\left(j-1\right)/mn,j/mn\right)$, $i=1,\ldots,mn$ we have
$J_{k(1,j),n}(a)-J_{k(2,j),m}(a)=\bar{a}_{n}(x)-\bar{a}_{m}(x)$ (2.79)
where $\bar{a}_{n}$ defined by (2.74). Hence
$\frac{1}{mn}\sum_{j=1}^{mn}\left(J_{k(1,j),n}\left(a\right)-J_{k(2,j),m}\left(a\right)\right)^{2}=\left\|\bar{a}_{n}-\bar{a}_{m}\right\|_{2}^{2}.$ (2.80)
Now as a consequence of Lemma 2.14 (i), if $a\in\Theta_{1}\left(\alpha,M\right)$ and $1/2<\alpha<1$
$\left\|a-\bar{a}_{n}\right\|_{2}^{2}\leq C_{\alpha}\;n^{-2\alpha}\;\left|a\right|_{2,\alpha}^{2}\leq C_{\alpha}\;n^{-2\alpha}M.$
If $\alpha\geq 1$ then for any $\beta\in\left(0,1\right)$ we have $\left|a\right|_{2,\beta}^{2}\leq\left|a\right|_{2,\alpha}^{2}$ and so if $a\in\Theta_{1}\left(\alpha,M\right)$ for $\alpha>1/2$ then there exists $\beta>1/2$ such that
$\left\|a-\bar{a}_{n}\right\|_{2}^{2}\leq C_{\beta}\;n^{-2\beta}M.$
Hence generally there exists a constant $C$ such that
$n\left\|\bar{a}_{n}-\bar{a}_{m}\right\|_{2}^{2}$ $\leq 2n\left\|\bar{a}_{n}-a\right\|_{2}^{2}+2n\left\|\bar{a}_{m}-a\right\|_{2}^{2}$
$\leq 2CM\left(n^{1-2\beta}+nm^{-2\beta}\right)\leq 4CMn^{1-2\beta}=o\left(1\right)$
uniformly over $a\in\Theta_{1}\left(\alpha,M\right)$. This relation along with (2.77)-(2.80) proves that
$\sup_{a\in\Theta_{1}\left(\alpha,M\right)}H^{2}\left({\displaystyle\bigotimes\limits_{j=1}^{mn}}Q_{1,j},{\displaystyle\bigotimes\limits_{j=1}^{mn}}Q_{2,j}\right)=o\left(1\right).$
Now (2.71) establishes the claim.
The remaining task is to compare $\mathcal{G}_{m,n}^{\ast}$ to $\mathcal{G}_{m}$.
Lemma 2.19
For $m=n+r_{n}$, $0\leq r_{n}=o\left(n^{-1/2}\right)$ we have asymptotic equivalence
$\mathcal{G}_{m}\approx\mathcal{G}_{m,n}^{\ast}\text{ as }n\rightarrow\infty.$
Proof. The sufficiency argument for the negative binomial applied in Lemma 2.17 can be used to show that
$\mathcal{G}_{m,n}^{\ast}\sim\mathcal{G}_{m}^{\ast}:=\left\{{\displaystyle\bigotimes\limits_{j=1}^{m}}\mathrm{NB}\left(nm^{-1},p\left(J_{j,m}\left(a\right)\right)\right),a\in\Theta\right\}.$
Now it suffices to show asymptotic total variation equivalence $\mathcal{G}_{m}^{\ast}\simeq\mathcal{G}_{m}$. Recall that $\mathrm{Geo}\left(p\right)=\mathrm{NB}\left(1,p\right)$ and note that for the Hellinger distance we have, according to Lemma A.9 (ii)
$H^{2}\left(\mathrm{NB}\left(1,p\left(J_{j,m}\left(a\right)\right)\right),\mathrm{NB}\left(nm^{-1},p\left(J_{j,m}\left(a\right)\right)\right)\right)\leq 1-\frac{\Gamma\left(\left(1+nm^{-1}\right)/2\right)}{\Gamma^{1/2}\left(1\right)\Gamma^{1/2}\left(nm^{-1}\right)}$
$\leq\frac{1}{\Gamma^{1/2}\left(nm^{-1}\right)}\left(\Gamma^{1/2}\left(nm^{-1}\right)-\Gamma\left(\left(1+nm^{-1}\right)/2\right)\right)$
where we used $\Gamma\left(1\right)=1$. Since the Gamma function is infinitely differentiable on $(0,\infty)$ and $nm^{-1}\rightarrow 1$, the first factor above is $1+o\left(1\right)$. Furthermore, write $n/m=1-\delta$ where $\delta=r_{n}/m$; by a Taylor expansion we obtain
$\Gamma\left(\left(1+n/m\right)/2\right)$ $=\Gamma(1-\delta/2)=1-\Gamma^{\prime}(1)\frac{\delta}{2}+O(\delta^{2}),$
$\Gamma^{1/2}\left(n/m\right)$ $=\Gamma^{1/2}(1-\delta)=1-\frac{1}{2}\Gamma^{\prime}(1)\delta+O(\delta^{2}).$
Consequently
$\left(\Gamma^{1/2}\left(nm^{-1}\right)-\Gamma\left(\left(1+nm^{-1}\right)/2\right)\right)$ $=O(\delta^{2})$
$=O\left(r_{n}^{2}/m^{2}\right).$
Applying (2.72) we find that the squared Hellinger distance between the respective product measures in $\mathcal{G}_{m}^{\ast}$ and $\mathcal{G}_{m}$ is of order
$mO\left(r_{n}^{2}/m^{2}\right)\leq O\left(r_{n}^{2}/n\right)=o\left(1\right)$
in view of the condition $r_{n}=o(n^{1/2})$. Applying (2.71) again establishes the claim $\mathcal{G}_{m}^{\ast}\simeq\mathcal{G}_{m}$.
[…]
2.7 Geometric regression and white noise
Consider an variant of the geometric regression model (1.29) where the local averages $J_{j,n}\left(a\right)$ of the spectral density $a$ are replaced by values at points
$t_{j,n}=2\pi\left(\frac{j}{n}-\frac{1}{2}\right)\text{, }j=1,\ldots,n.$ (2.81)
Accordingly define the experiment
$\mathcal{F}_{n}^{\prime}\left(\Theta\right):=\left\{{\displaystyle\bigotimes\limits_{j=1}^{n}}\mathrm{Geo}\left(p\left(a\left(t_{j,n}\right)\right)\right),a\in\Theta\right\}$ (2.82)
where $p\left(x\right)=\left(x-1\right)/\left(x+1\right)$ for $x>1$. To introduce an appropriate class of spectral densities $a$ with this model, define the Hölder norm for functions on $\left[-\pi,\pi\right],$ with $\alpha\in\left(0,1\right]$
$\left\|f\right\|_{C^{\alpha}}:=\left\|f\right\|_{\infty}+\sup_{x\neq y}\frac{\left|f(x)-f(y)\right|}{\left|x-y\right|^{\alpha}}$ (2.83)
and the corresponding Hölder class of functions
$C^{\alpha}(M):=\left\{f:\left[-\pi,\pi\right]\rightarrow\mathbb{R},\;\;\left\|f\right\|_{C^{\alpha}}\leq M\right\}.$ (2.84)
The periodic Sobolev norm $\left\|\cdot\right\|_{2,\alpha}$ for functions $f$ on $\left[-\pi,\pi\right]$ for smoothness index $\alpha>0$ is given by (2.6). A basic embedding theorem ([GNZ09], Lemma 5.6) gives a norm inequality, for $\alpha\in\left(0,1\right]$
$\left\|f\right\|_{C^{\alpha}}\leq C\left\|f\right\|_{2,\alpha+1/2}^{2}$
where $C$ depends only on $\alpha$. Thus, if we consider a set of spectral densities, in analogy to (2.9) and (1.31)
$\Theta_{1,c}\left(\alpha,M\right)$ $:=C^{\alpha}(M)\cap\mathcal{F}_{M},$ (2.85)
$\mathcal{F}_{M}$ $:=\left\{f:\left[-\pi,\pi\right]\rightarrow\mathbb{R},\text{ }\;f\left(\omega\right)\geq 1+M^{-1}\text{, }\omega\in\left[-\pi,\pi\right]\right\}$ (2.86)
then we have the inclusion, for $\alpha\in\left(0,1\right]$
$\Theta_{1}\left(\alpha+1/2,M\right)\subset\Theta_{1,c}\left(\alpha,M^{\prime}\right)$ (2.87)
for some $M^{\prime}>0$.
Lemma 2.20
If $\Theta=\Theta_{1,c}\left(\alpha,M\right)$ for $\alpha\in\left(1/2,1\right],$ $M>0$ then we have asymptotic total variation equivalence
[…]
Proof. As with Lemma 2.15 it suffices to prove for the Hellinger distance $H\left(\cdot,\cdot\right)$
$\sum_{j=1}^{n}H^{2}\left(\mathrm{Geo}\left(p\left(a\left(t_{j,n}\right)\right)\right),\mathrm{Geo}\left(p\left(J_{j,n}\left(a\right)\right)\right)\right)=o\left(1\right)$ (2.88)
uniformly over $a\in\Theta$. According to (2.75) we have
$H^{2}\left(\mathrm{Geo}\left(p\left(a\left(t_{j,n}\right)\right)\right),\mathrm{Geo}\left(p\left(J_{j,m}\left(a\right)\right)\right)\right)\leq\frac{\left(a\left(t_{j,n}\right)-J_{j,m}\left(a\right)\right)^{2}}{\left(a\left(t_{j,n}\right)-1\right)\left(J_{j,m}\left(a\right)-1\right)}.$
Here for $a\in$ $\Theta_{1,H}\left(\alpha,M\right)$ we have $a\left(t_{j,n}\right)-1\geq M^{-1}$, $J_{j,m}\left(a\right)-1\geq M^{-1}$, hence
$H^{2}\left(\mathrm{Geo}\left(p\left(a\left(t_{j,n}\right)\right)\right),\mathrm{Geo}\left(p\left(J_{j,m}\left(a\right)\right)\right)\right)\leq M^{2}\left(a\left(t_{j,n}\right)-J_{j,m}\left(a\right)\right)^{2}.$
Recalling the definition of the intervals $W_{j,n}$ in (2.73) and (2.74), we obtain
$\left|a\left(t_{j,n}\right)-J_{j,n}\left(a\right)\right|$ $=\frac{n}{2\pi}\left|\int_{W_{j,n}}\left(f(x)-a\left(t_{j,n}\right)\right)dx\right|$
$\leq M\left(\frac{2\pi}{n}\right)^{\alpha}$
and hence the l.h.s. of (2.88) is bounded by
$\sum_{j=1}^{n}M^{2}\left(a\left(t_{j,n}\right)-J_{j,m}\left(a\right)\right)^{2}\leq M^{4}\left(2\pi\right)^{2\alpha}\;n^{1-2\alpha}=o\left(1\right).$
Consider again the probability measures $Q_{n,1}\left(a\right)$ given by the white noise model (1.33).
Lemma 2.21
For $\Theta=\Theta_{1,c}\left(\alpha,M\right)$ consider the experiment $\mathcal{G}_{n,1}\left(\Theta\right)=\left\{Q_{n,1}\left(a\right),\;a\in\Theta\right\}$. If $\alpha\in\left(1/2,1\right],$ $M>0$ then
$\Delta\left(\mathcal{F}_{n}^{\prime}\left(\Theta\right),\mathcal{G}_{n,1}\left(\Theta\right)\right)\rightarrow 0\text{ as }n\rightarrow\infty.$
Proof. This follows from the results of [GN98]. Let $\left\{Q\left(\tau\right),\tau\in T\right\}$ be a one parameter exponential family where $\tau$ is the canonical (natural) parameter and $T=\left[t_{1},t_{2}\right]$ is a closed interval in $\mathbb{R}$. It is assumed in [GN98] that $X_{i}$ are independent observations having distributions $Q\left(f\left(u_{i}\right)\right)$ where $f$ is a function $f:\left[0,1\right]\rightarrow T$ and $u_{i}=i/n$, $i=1,\ldots,n$. The following regularity condition is assumed: $T$ is in the interior of the natural parameter space of the exponential family, and there exist $\varepsilon>0$ and constants $C_{1},C_{2}$ such the Fisher information $I\left(\tau\right)$ fulfills
$0<C_{1}\leq I\left(\tau\right)\leq C_{2}<\infty,\;\tau\in\left[t_{1}-\varepsilon,t_{2}+\varepsilon\right].$ (2.89)
According to Subsection A.4, the geometric distributions have densities (with respect to counting measure $\mu$ on $\mathbb{Z}_{+}$) which can be written as those of an exponential family of densities in canonical form, cf. (A.56):
$Q\left(\tau\right)\left(x\right)=\exp\left(\tau x-V\left(\tau\right)\right)\text{, }x\in\mathbb{Z}_{+}$
where $\tau=\log p$ and $P\left(X=x\right)=\left(1-p\right)p^{x}$. In our setting, $\tau$ will be parametrized according to (A.59) as $\tau=\log\left(\left(a-1\right)/\left(a+1\right)\right)$, so if $a\in\left[1+M^{-1},M\right]$ for some $M>2$ then $\tau\in\left[t_{1},t_{2}\right]$ for some $t_{1},t_{2}$ fulfilling $\infty<t_{1}<t_{2}<0$. For the Fisher information $I\left(\tau\right)$ we have according to (A.58)
$I\left(\tau\right)=\frac{\exp\tau}{\left(1-\exp\tau\right)^{2}}$ (2.90)
such that (2.89) is fulfilled for sufficiently small $\varepsilon$.
In [GN98] the function $f$ is defined on $[0,1]$ and assumed to vary in a smoothness class $C_{1}^{\alpha}(M)$, defined as the analog of $C^{\alpha}(M)$ from (2.84) on the interval $[0,1]$. It is easy to see that the experiment $\mathcal{G}_{n,1}\left(\Theta\right)$ can be cast in this form. Indeed define functions
$H\left(a\right)$ $=\log\frac{a-1}{a+1}\text{ for }a\in\left[1+M^{-1},M\right],$
$s\left(x\right)$ $=2\pi\left(x-1/2\right).$
Note that $s\left(u_{j}\right)=s\left(j/n\right)=t_{j,n}$, $j=1,\ldots,n.$ Thus in $\mathcal{G}_{n}^{\prime}\left(\Theta\right)$ observations $X_{j}$ are independent with distribution
$X_{j}\sim Q\left(H\left(a\left(s\left(j/n\right)\right)\right)\right)\text{, }j=1,\ldots,n.$
Setting
$f\left(x\right)=H\left(a\left(s\left(x\right)\right)\right),$
[…]
$M^{\prime}=M\left(2\pi\right)^{\alpha}\sup_{z\in\left[1+M^{-1},M\right]}H^{\prime}\left(z\right).$
Also, since $H$ is strictly increasing on $\left[1+M^{-1},M\right]$ (cf. (A.60)), the function $f$ takes values in $T=\left[t_{1},t_{2}\right]$ with $t_{1}=H\left(1+M^{-1}\right)$, $t_{2}=H\left(M\right)$, $t_{1}<t_{2}<0$. Thus the experiment
$\left\{{\displaystyle\bigotimes\limits_{j=1}^{n}}Q\left(f\left(j/n\right)\right)\text{, }f=H\circ a\circ s\text{, }a\in\Theta_{1,c}\left(\alpha,M\right)\right\}$
can be approximated in $\Delta$-distance by the white noise model
$dZ_{x}=G\left(f\left(x\right)\right)dx+n^{-1/2}dW_{x}\text{, }x\in\left[0,1\right]$ (2.91)
where $f=H\circ a\circ s$, the function $a$ varies in $\Theta_{1,c}\left(\alpha,M\right)$ and $G$ is the variance stabilizing transform pertaining to the exponential family $\left\{Q\left(\tau\right),\tau\in T\right\}$ (cf. Section 3.3 of [GN98] or Remark 3.3 in [GN02]). Here $G$ is unique up to additive constants; $G$ fulfills
$\frac{d}{d\tau}G\left(\tau\right)=\sqrt{I\left(\tau\right)}$
with $I\left(\tau\right)$ given by (2.90). Finding the function $G$ is equivalent to finding the function
$g\left(a\right):=G\left(H\left(a\right)\right)\text{, }a\in\left(1,\infty\right).$
We have
$\frac{d}{da}g\left(a\right)$ $=G^{\prime}\left(H\left(a\right)\right)\;H^{\prime}\left(a\right)$
$=\sqrt{I\left(H\left(a\right)\right)}\;H^{\prime}\left(a\right).$
By (A.58) and (A.61)
$I\left(H\left(a\right)\right)=V^{\prime\prime}\left(H\left(a\right)\right)=\frac{a^{2}-1}{4}$
whereas by (A.60)
$H^{\prime}\left(a\right)=\frac{2}{a^{2}-1}.$
Thus $g$ must fulfill
$\frac{d}{da}g\left(a\right)=\frac{\sqrt{a^{2}-1}}{2}\frac{2}{a^{2}-1}=\frac{1}{\sqrt{a^{2}-1}}.$ (2.92)
It can be checked that the function
$g(x)=\mathrm{arc\cosh}\left(x\right)=\log\left(x+\sqrt{x^{2}-1}\right),x>1$
fulfills (2.92). From (2.91) we obtain that the experiment given by $Z=\left\{Z_{x},x\in\left[0,1\right]\right\}$ with
$dZ_{x}=g\left(a\left(s\left(x\right)\right)\right)dx+n^{-1/2}dW_{x}\text{, }x\in\left[0,1\right]$ (2.93)
and $a\in\Theta=\Theta_{1,c}\left(\alpha,M\right)$ is asymptotically equivalent to $\mathcal{G}_{n}^{\prime}\left(\Theta\right)$. Define the stochastic process $Y=\left\{Y_{\omega},\omega\in\left[-\pi,\pi\right]\right\}$ by $Y_{\omega}=2\pi Z_{s^{-1}\left(\omega\right)}$; then $Y$ satisfies
$dY_{\omega}=g\left(a\left(\omega\right)\right)d\omega+\left(2\pi/n\right)^{1/2}dW_{\omega}\text{, }\omega\in\left[-\pi,\pi\right]$ (2.94)
so that according to (1.33), $Y$ has distribution $Q_{n,1}\left(a\right)$. The claim now follows from the fact that the mapping between the processes $Y$ and $Z$ is one-to-one.
Proof of Theorem 1.2. Consider experiment $\mathcal{F}_{n}\left(\Theta\right)$ for $\Theta=\Theta_{1}\left(\alpha,M\right)$ where $\alpha>1$. By relation (2.87) one has $\Theta_{1}\left(\alpha,M\right)\subset\Theta_{1,c}\left(\alpha-1/2,M^{\prime}\right)$. From Lemma 2.20 it then follows that $\mathcal{F}_{n}\left(\Theta\right)\approx\mathcal{F}_{n}^{\prime}\left(\Theta\right)$, and Lemma 2.21 implies that $\mathcal{F}_{n}^{\prime}\left(\Theta\right)\approx\mathcal{G}_{n}\left(\Theta\right)$. By the transitivity of the equivalence relation $\approx$ for sequences of experiments, one has $\mathcal{F}_{n}\left(\Theta\right)\approx\mathcal{G}_{n}\left(\Theta\right)$ as claimed.
For later reference we note a localized version of the white noise model (1.33), where $a$ itself appears as the drift function rather than the $\mathrm{arc\cosh}$-transformation, but the approximation holds in a neighborhood of a fixed function $a_{\left(0\right)}\in\Theta_{1,c}\left(\alpha,M\right)$. This will be the analog of the localized white noise approximation (1.6) for the classical stationary Gaussian process. Define for some sequence $\gamma_{n}=o\left(1\right)$
$B\left(a_{0},\gamma_{n}\right).=\left\{a:\left[-\pi,\pi\right]\rightarrow\mathbb{R}\text{, }\left\|a-a_{\left(0\right)}\right\|_{\infty}\leq\gamma_{n}\right\}$
and consider restricted function sets
$\Theta_{1,c}\left(\alpha,M\right)\cap B\left(a_{\left(0\right)},\gamma_{n}\right).$ (2.95)
Furthermore let $Q_{n.2}\left(a,a_{\left(0\right)}\right)$ be the distribution of the process $Y=\left\{Y_{\omega},\omega\in\left[-\pi,\pi\right]\right\}$ described by
$dY_{\omega}=a\left(\omega\right)d\omega+\left(2\pi/n\right)^{1/2}\left(a_{\left(0\right)}^{2}\left(\omega\right)-1\right)^{1/2}dW_{\omega}\text{, }\omega\in\left[-\pi,\pi\right]$ (2.96)
and $Y_{\omega}=\int_{-\pi}^{\omega}dY_{\omega}$, and define the experiment
$\mathcal{G}_{n,2}\left(a_{\left(0\right)},\Theta\right):=\left\{Q_{n}\left(a,a_{\left(0\right)}\right),\;a\in\Theta\right\}.$ (2.97)
At this point we use notation $\mathcal{G}_{n,1}\left(\Theta\right):=\mathcal{G}_{n}\left(\Theta\right)$ where $\mathcal{G}_{n}\left(\Theta\right)$ describes the experiment given by (1.33), i.e. by
$dY_{\omega}=\mathrm{arc\cosh}\left(a\left(\omega\right)\right)d\omega+\left(2\pi/n\right)^{1/2}dW_{\omega},\omega\in\left[-\pi,\pi\right]$ (2.98)
with $a\in\Theta$.
Lemma 2.22
Assume $\alpha\in\left(1/2,1\right]$. Then for every sequence $\gamma_{n}=o\left(\left(n/\log n\right)^{-\alpha/(2\alpha+1)}\right)$ and $\Theta_{n}=\Theta_{1,c}\left(\alpha,M\right)\cap B\left(a_{\left(0\right)},\gamma_{n}\right)$ one has
$\sup_{a_{0}\in\Theta_{1,c}\left(\alpha,M\right)}\Delta\left(\mathcal{G}_{n,1}\left(\Theta_{n}\right),\mathcal{G}_{n,2}\left(a_{\left(0\right)},\Theta_{n}\right)\right)\rightarrow 0\text{ as }n\rightarrow\infty.$
Proof. This is essentially Theorem 3.3 in [GN98], specialized to the present exponential family, i.e. the geometric distribution. The white noise model (3.8) in [GN98] corresponds to (2.96), and the variance-stable white noise model (3.15) in [GN98] corresponds to (2.98). The models in [GN98] are defined on the unit interval, but the result carries over $\left[-\pi,\pi\right]$ in the same way as has been noted with processes (2.93) and (2.94).
3 Lower informativity bound
3.1 Constructing the basic observables
In this section we assume $n$ is an odd number. Consider the creation and annihilation operators $\hat{A}_{j}=\frac{1}{\sqrt{2}}\left(Q_{j}+iP_{j}\right)$, $\hat{A}_{j}^{\ast}=\frac{1}{\sqrt{2}}\left(Q_{j}-iP_{j}\right)$. As a consequence of (1.7), these fulfill the commutation relations
$\left[\hat{A}_{j},\hat{A}_{j}^{\ast}\right]$ $=\mathbf{1},\;\;\;\;j=1,\ldots,n$ (3.1)
$\;\left[\hat{A}_{j},\hat{A}_{k}^{\ast}\right]$ $=\left[\hat{A}_{j},\hat{A}_{k}\right]=\left[\hat{A}_{j}^{\ast},\hat{A}_{k}^{\ast}\right]=0,\;\;j,k=1,\ldots,n,j\neq k,$ (3.2)
Furthermore
$\hat{A}_{j}^{\ast}\hat{A}_{j}=\frac{1}{2}\left(Q_{j}^{2}+P_{j}^{2}-\mathbf{1}\right)$ (3.3)
are the number operators. Thus $\hat{A}_{j}^{\ast}\hat{A}_{j}$, $j=1,\ldots,n$ is a commuting set of observables; the following lemma describes the first and second moment properties of this set.
Lemma 3.1
Let $\rho=\mathfrak{N}_{n}\left(0,A\right)$ for a symbol matrix $A=\left(a_{jk}\right)_{j,k=1}^{n}$ fulfilling $A>I$ (not necessarily Toeplitz). Then we have for $j,k=1,\ldots,n$
[…]
Proof. (i) Consider first the case $j=k$. Then $\hat{A}_{j}^{\ast}\hat{A}_{j}$ is the number operator of the $j$-th mode, and its distribution under $\rho$ is the same as under the marginal state of the $j$-th mode, $\rho_{(j)}$ say, i.e. the partial trace of $\rho$ when all other modes are traced out. By a reasoning analogous to Subsection A.2.1, it follows that $\rho_{(j)}=\mathfrak{N}_{1}\left(0,a_{jj}\right)$, which according to (1.16) and (1.19) can also be described as $\varphi\left(0,\Sigma\right)$ for $\Sigma=\frac{1}{2}a_{jj}I_{2}$. Thus $\rho_{(j)}$ is the thermal state with covariance matrix $\frac{1}{2}a_{jj}I_{2}$ (cp. also (2.69)), where the number operator has a geometric distribution:
[…]
The expectation is (cf. Subsection A.4)
$\left\langle\hat{A}_{j}^{\ast}\hat{A}_{j}\right\rangle_{\rho}=\frac{p}{1-p}=\frac{a_{jj}-1}{2}$ (3.5)
which proves the claim for $j=k$. For $j\neq k$
$Tr\left[\hat{A}_{j}^{\ast}\hat{A}_{k}\rho\right]$ $=\frac{1}{2}\left\langle\left(Q_{j}-iP_{j}\right)\left(Q_{k}+iP_{k}\right)\right\rangle_{\rho}$
$=\frac{1}{2}\left(\left\langle Q_{j}Q_{k}\right\rangle_{\rho}+\left\langle P_{j}P_{k}\right\rangle_{\rho}+i\left\langle Q_{j}P_{k}\right\rangle_{\rho}-i\left\langle P_{j}Q_{k}\right\rangle_{\rho}\right).$ (3.6)
Consider the marginal state $\rho_{\left(j,k\right)}$ of $\rho$ where all modes except $j$ and $k$ are traced out. Again, by a reasoning analogous to Subsection A.2.1, it follows that $\rho_{(j,k)}=\mathfrak{N}_{2}\left(0,A_{\left(j,k\right)}\right)$ where $A_{\left(j,k\right)}$ is the submatrix of AA
$A_{\left(j,k\right)}=\left(\begin{array}[c]{cc}a_{jj}&a_{jk}\\ a_{kj}&a_{kk}\end{array}\right).$
According to (1.16), the covariance matrix of $\rho_{(j,k)}$ is
$\Sigma\left(A_{\left(j,k\right)}\right)$ $=\frac{1}{2}\left(\begin{array}[c]{cc}\operatorname{Re}A_{\left(j,k\right)}&-\operatorname{Im}A_{\left(j,k\right)}\\ \operatorname{Im}A_{\left(j,k\right)}&\operatorname{Re}A_{\left(j,k\right)}\end{array}\right)$ (3.9)
$=\frac{1}{2}\left(\begin{array}[c]{cc}\begin{array}[c]{cc}a_{jj}&\operatorname{Re}a_{jk}\\ \operatorname{Re}a_{jk}&a_{kk}\end{array}&\begin{array}[c]{cc}0&-\operatorname{Im}a_{jk}\\ \operatorname{Im}a_{jk}&0\end{array}\\ \begin{array}[c]{cc}0&\operatorname{Im}a_{jk}\\ -\operatorname{Im}a_{jk}&0\end{array}&\begin{array}[c]{cc}a_{jj}&\operatorname{Re}a_{jk}\\ \operatorname{Re}a_{jk}&a_{kk}\end{array}\end{array}\right).$ (3.20)
Since this covariance matrix pertains to the vector of observables $\mathbf{R}=\left(Q_{j},Q_{k},P_{j},P_{k}\right)$ in the sense that $\mathbf{R}x\sim N\left(0,\left\langle x,\Sigma\left(A_{\left(j,k\right)}\right)x\right\rangle\right)$ (cp. (1.13), we can directly read off the covariances:
$\left\langle Q_{j}Q_{k}\right\rangle_{\rho}$ $=\left\langle P_{j}P_{k}\right\rangle_{\rho}=\frac{1}{2}\operatorname{Re}a_{jk},$
$\left\langle Q_{j}P_{k}\right\rangle_{\rho}$ $=-\frac{1}{2}\operatorname{Im}a_{jk},\;\left\langle P_{j}Q_{k}\right\rangle_{\rho}=\frac{1}{2}\operatorname{Im}a_{jk}.$
From (3.6) we obtain
$Tr\left[\hat{A}_{j}^{\ast}\hat{A}_{k}\rho\right]=\frac{1}{2}\operatorname{Re}a_{jk}-\frac{1}{2}i\operatorname{Im}a_{jk}=\frac{1}{2}\bar{a}_{jk}=\frac{1}{2}a_{kj}.$
(ii) Consider first the case $j<k$. Then in view of (3.3)
$\hat{A}_{j}^{\ast}\hat{A}_{j}\hat{A}_{k}^{\ast}\hat{A}_{k}=\frac{1}{4}\left(Q_{j}^{2}+P_{j}^{2}-\mathbf{1}\right)\left(Q_{k}^{2}+P_{k}^{2}-\mathbf{1}\right),$
hence
$4\cdot\hat{A}_{j}^{\ast}\hat{A}_{j}\hat{A}_{k}^{\ast}\hat{A}_{k}$ $=Q_{j}^{2}Q_{k}^{2}+Q_{j}^{2}P_{k}^{2}-Q_{j}^{2}$
$+P_{j}^{2}Q_{k}^{2}+P_{j}^{2}P_{k}^{2}-P_{j}^{2}$
$-Q_{k}^{2}-P_{k}^{2}+\mathbf{1}.$ (3.21)
Note that on the r.h.s. above, each summand $Q_{j}^{2}Q_{k}^{2}$, $Q_{j}^{2}P_{k}^{2}$ etc. contains only commuting observables, which thus have a joint distribition. In view of (3.20), the joint distribution of $Q_{j},Q_{k}$ is
$\left(Q_{j},Q_{k}\right)\sim N_{2}\left(0,\frac{1}{2}\operatorname{Re}A_{(j,k)}\right).$
From formula (A.77) in Subsection A.6 we obtain
$\left\langle Q_{j}^{2}Q_{k}^{2}\right\rangle_{\rho}=\frac{1}{2}\left(\operatorname{Re}a_{jk}\right)^{2}+\frac{1}{4}a_{jj}a_{kk}.$
Similarly
$\left(Q_{j},P_{k}\right)$ $\sim N_{2}\left(0,\frac{1}{2}\left(\begin{array}[c]{cc}a_{jj}&-\operatorname{Im}a_{jk}\\ -\operatorname{Im}a_{jk}&a_{kk}\end{array}\right)\right),$
$\left\langle Q_{j}^{2}P_{k}^{2}\right\rangle_{\rho}$ $=\frac{1}{2}\left(\operatorname{Im}a_{jk}\right)^{2}+\frac{1}{4}a_{jj}a_{kk},$
$\left\langle P_{j}^{2}Q_{k}^{2}\right\rangle_{\rho}$ $=\frac{1}{2}\left(\operatorname{Im}a_{jk}\right)^{2}+\frac{1}{4}a_{jj}a_{kk},$
$\left\langle P_{j}^{2}P_{k}^{2}\right\rangle_{\rho}$ $=\frac{1}{2}\left(\operatorname{Re}a_{jk}\right)^{2}+\frac{1}{4}a_{jj}a_{kk}.$
Furthermore
$\left\langle Q_{j}^{2}+P_{j}^{2}+Q_{k}^{2}+P_{k}^{2}\right\rangle_{\rho}=a_{jj}+a_{kk}.$
Collecting terms in (3.21), we obtain
$4\cdot\left\langle\hat{A}_{j}^{\ast}\hat{A}_{j}\hat{A}_{k}^{\ast}\hat{A}_{k}\right\rangle_{\rho}=\left|a_{jk}\right|^{2}+a_{jj}a_{kk}-\left(a_{jj}+a_{kk}\right)+1.$
Also from (3.5)
$\left\langle\hat{A}_{j}^{\ast}\hat{A}_{j}\right\rangle_{\rho}=\frac{1}{2}\left(a_{jj}-1\right)\text{, }\left\langle\hat{A}_{k}^{\ast}\hat{A}_{k}\right\rangle_{\rho}=\frac{1}{2}\left(a_{kk}-1\right)$
hence
$4\cdot\mathrm{Cov}_{\rho}\left(\hat{A}_{j}^{\ast}\hat{A}_{j},\hat{A}_{k}^{\ast}\hat{A}_{k}\right)$ $=4\cdot\left(\left\langle\hat{A}_{j}^{\ast}\hat{A}_{j}\hat{A}_{k}^{\ast}\hat{A}_{k}\right\rangle_{\rho}-\left\langle\hat{A}_{j}^{\ast}\hat{A}_{j}\right\rangle_{\rho}\cdot\left\langle\hat{A}_{k}^{\ast}\hat{A}_{k}\right\rangle_{\rho}\right)$
$=\left|a_{jk}\right|^{2}+a_{jj}a_{kk}-\left(a_{jj}+a_{kk}\right)+1-\left(a_{jj}-1\right)\left(a_{kk}-1\right)$
$=\left|a_{jk}\right|^{2}$
which proves the claim for $j<k$. For $j=k$, according to relation (3.4) and the formula for the variance of the geometric (A.58) we have
$\mathrm{Var}_{\rho}\left(\hat{A}_{j}^{\ast}\hat{A}_{j}\right)$ $=\frac{p}{\left(1-p\right)^{2}}=\frac{a_{jj}-1}{a_{jj}+1}\frac{\left(a_{jj}+1\right)^{2}}{4}$
$=\frac{1}{4}\left(a_{jj}^{2}-1\right).$
We note the following consequence of Lemma 3.1:
$\left\langle\hat{A}_{j}\hat{A}_{j}^{\ast}\right\rangle_{\rho}$ $=\left\langle\hat{A}_{j}^{\ast}\hat{A}_{j}+\mathbf{1}\right\rangle_{\rho}=\left\langle\hat{A}_{j}^{\ast}\hat{A}_{j}\right\rangle_{\rho}+1=\frac{1}{2}\left(a_{jj}+1\right),$ (3.22)
$\left\langle\hat{A}_{j}\hat{A}_{k}^{\ast}\right\rangle_{\rho}$ $=\left\langle\hat{A}_{k}^{\ast}\hat{A}_{j}\right\rangle_{\rho}=\frac{1}{2}a_{jk}\text{ for }j\neq k.$ (3.23)
Define vectors of operators
$\mathbf{\hat{A}}$ $=\left(\begin{array}[c]{c}\hat{A}_{1}\\ \ldots\\ \hat{A}_{n}\end{array}\right),$
$\mathbf{\hat{A}}^{{\dagger}}$ $\mathbf{=}\left(\hat{A}_{1}^{\ast},\ldots,\hat{A}_{n}^{\ast}\right).$
For a matrix of operators $\mathbf{C}=\left(C_{jk}\right)$, introduce notation $\left\langle\mathbf{C}\right\rangle_{\rho}=\left(\left\langle C_{jk}\right\rangle_{\rho}\right)$. Then (3.22), (3.23) can be written
$\left\langle\mathbf{\hat{A}\hat{A}}^{{\dagger}}\right\rangle_{\rho}=\frac{1}{2}\left(A+I_{n}\right).$ (3.24)
For the special unitary $U_{n}$ from (2.55) we set
$\mathbf{\hat{B}}=U_{n}^{\ast}\mathbf{\hat{A}}\text{, }\mathbf{\hat{B}}^{{\dagger}}\mathbf{=\hat{A}}^{{\dagger}}U_{n}.$ (3.25)
It then follows that
$\left\langle\mathbf{\hat{B}\mathbf{\hat{B}}^{{\dagger}}}\right\rangle_{\rho}=\frac{1}{2}\left(U_{n}^{\ast}AU_{n}+I_{n}\right).$ (3.26)
Since $\mathbf{\hat{B}}$ represents a discrete Fourier transform of the creation operators, for the components of the vector $\mathbf{\hat{B}}$ we adopt the indexing convention $\mathbf{\hat{B}=}\left(\hat{B}_{j}\right)_{\left|j\right|\leq\left(n-1\right)/2}$. This is in agreement with the form of the unitary $U_{n}$ in (2.55); we then obtain for the components of the vector $\mathbf{\hat{B}}=U_{n}^{\ast}\mathbf{\hat{A}}$
$\hat{B}_{j}=\mathbf{u}_{j}^{\ast}\mathbf{\hat{A}}\text{, }\left|j\right|\leq\left(n-1\right)/2.$
Lemma 3.2
The set of operators $\hat{B}_{j}$, $\left|j\right|\leq\left(n-1\right)/2$ fulfills commutation relations (3.1), (3.2) with $\hat{A}_{j}$ replaced by $\hat{B}_{j-\left(n+1\right)/2}$.
Proof. Relations (3.1), (3.2) can be expressed in concise form as follows: for any $c,d\in\mathbb{C}^{n}$ and $c^{\ast}\mathbf{\hat{A}=}\sum_{j=1}^{n}\bar{c}_{j}\hat{A}_{j}$, $\mathbf{\hat{A}}^{{\dagger}}d=\sum_{j=1}^{n}d_{j}\hat{A}_{j}^{\ast}$ we have
$\left[c^{\ast}\mathbf{\hat{A},\hat{A}}^{{\dagger}}d\right]=\left\langle c,d\right\rangle\;\mathbf{1}.$
Now with definitions (3.25) we have indeed
$\left[c^{\ast}\mathbf{\hat{B},\mathbf{\hat{B}}^{{\dagger}}}d\right]=\left[c^{\ast}U_{n}^{\ast}\mathbf{\hat{A},\mathbf{\hat{A}}^{{\dagger}}}U_{n}d\right]=\left\langle U_{n}c,U_{n}d\right\rangle\;\mathbf{1}=\left\langle c,d\right\rangle\;\mathbf{1.}$
Lemma 3.3
$\hat{B}_{j}^{\ast}\hat{B}_{j}$, $\left|j\right|\leq\left(n-1\right)/2$ is a commuting set of observables, fulfilling
$\hat{B}_{j}^{\ast}\hat{B}_{j}=\hat{B}_{j}\hat{B}_{j}^{\ast}-\mathbf{1.}$ (3.27)
Proof. The first claim follows from (3.2) and the previous lemma. The claimed equality follows from (3.1) applied to $\hat{B}_{j}$, $\hat{B}_{j}^{\ast}$.
Lemma 3.4
Assume the conditions of Lemma 3.1. Then we have for $\left|j\right|,\left|k\right|\leq\left(n-1\right)/2\;$
$\left\langle\hat{B}_{j}^{\ast}\hat{B}_{j}\right\rangle_{\rho}=\frac{1}{2}\left(\mathbf{u}_{j}^{\ast}A\mathbf{u}_{j}-1\right),$
[…]
Proof. For (i), we note that (3.26) implies
$\left\langle\hat{B}_{j}\hat{B}_{j}^{\ast}\right\rangle_{\rho}=\frac{1}{2}\left(\mathbf{u}_{j}^{\ast}A\mathbf{u}_{j}^{\ast}+I_{n}\right).$
so that the claim follows from (3.27). For (ii), note that this claim can be formulated as: if in Lemma 3.1 the $\hat{A}_{j}$ are replaced by $\hat{B}_{j}$ then the assertion (ii) holds with the matrix $A$ replaced by $U_{n}^{\ast}AU_{n}$. Define a set of observables $\tilde{Q}_{j},\tilde{P}_{j}$, $j=1,\ldots,n$ by
$\tilde{Q}_{j-(n+1)/2}=\frac{1}{\sqrt{2}}\left(\hat{B}_{j}+\hat{B}_{j}^{\ast}\right)\text{, }\tilde{P}_{j-(n+1)/2}=\frac{1}{i\sqrt{2}}\left(\hat{B}_{j}-\hat{B}_{j}^{\ast}\right).$ (3.28)
These are related to $\hat{B}_{j}$ and $\hat{B}_{j}^{\ast}$ in the same way as the original canonical observables $P_{j},Q_{j}$ are related to the creation and annihilation operators $\hat{A}_{j}$ and $\hat{A}_{j}^{\ast}$. Due to Lemma 3.2, the set $\tilde{P}_{j},\tilde{Q}_{j}$, $j=1,\ldots,n$ fulfills the same basic commutation relations (1.7). Note that the proof of Lemma 3.1 is based on moment properties of the set of canonical observables $P_{j},Q_{j}$, implied by the fact that their covariance matrix is $\Sigma\left(A\right)$ from (1.16). Hence it suffices to show that the covariance matrix of $\tilde{P}_{j},\tilde{Q}_{j}$, $j=1,\ldots,n$ is $\Sigma\left(U_{n}^{\ast}AU_{n}\right)$. To see this, define the vector of observables
$\mathbf{\tilde{R}}:=\left(\tilde{Q}_{1},\ldots,\tilde{Q}_{n},\tilde{P}_{1},\ldots,\tilde{P}_{n}\right)$
in analogy to the $\mathbf{R}$ occurring in (1.8). Then for every $x\in\mathbb{R}^{2n}$ we have to show, for $\rho=\mathfrak{N}_{n}\left(0,A\right)$
$\mathrm{Tr}\;\rho\exp\left(i\mathbf{\tilde{R}}x\right)=\exp\left(-\frac{1}{2}\left\langle x,\Sigma\left(U_{n}^{\ast}AU_{n}\right)x\right\rangle\right).$ (3.29)
Recall that in connection with (1.15) for $u\in\mathbb{C}^{n}$ we set $\underline{u}:=\left(-\operatorname{Im}u\right)\oplus\operatorname{Re}u$. Setting $x=\underline{u}$ for some $u\in\mathbb{C}^{n}$, we note that (1.12) and (1.18) imply
$\left\langle\underline{u},\Sigma\left(A\right)\underline{u}\right\rangle=\frac{1}{2}\left\langle u,Au\right\rangle\text{, }u\in\mathbb{C}^{n}$
for every symbol matrix $A$, so that (3.29) is equivalent to
$\mathrm{Tr}\;\rho\exp\left(i\mathbf{\tilde{R}}\underline{u}\right)=\exp\left(-\frac{1}{4}\left\langle u,U_{n}^{\ast}AU_{n}u\right\rangle\right)\text{, }u\in\mathbb{C}^{n}.$ (3.30)
Define
$\mathbf{\tilde{R}}_{Q}:=\left(\tilde{Q}_{1},\ldots,\tilde{Q}_{n}\right),\;\mathbf{\tilde{R}}_{P}=\left(\tilde{P}_{1},\ldots,\tilde{P}_{n}\right)$
and set $x=x_{1}\oplus x_{2}$, $x_{i}\in\mathbb{R}^{n}$, $i=1,2$. Then
$\mathbf{\tilde{R}}x$ $=\mathbf{\tilde{R}}_{Q}x_{1}+\mathbf{\tilde{R}}_{P}x_{2}$
$=\frac{1}{\sqrt{2}}\left(x_{1}^{\prime}\mathbf{\hat{B}+\mathbf{\hat{B}}^{{\dagger}}}x_{1}\right)+\frac{1}{i\sqrt{2}}\left(x_{2}^{\prime}\mathbf{\hat{B}-\mathbf{\hat{B}}^{{\dagger}}}x_{2}\right)$
Define $u_{x}\in\mathbb{C}^{n}$ by $u_{x}=x_{2}-ix_{1}$. Then we obtain
$i\mathbf{\tilde{R}}x=2^{-1/2}\left(u_{x}^{\ast}\mathbf{\hat{B}-\mathbf{\hat{B}}^{{\dagger}}}u_{x}\right).$ (3.31)
Analogously one shows for $\mathbf{R}$
$i\mathbf{R}x=2^{-1/2}\left(u_{x}^{\ast}\mathbf{\hat{A}-\mathbf{\hat{A}}^{{\dagger}}}u_{x}\right),$
and thus the Weyl unitaries can be written
$W\left(x\right)=\exp\left(i\mathbf{R}x\right)=\exp\left(2^{-1/2}\left(u_{x}^{\ast}\mathbf{\hat{A}-\mathbf{\hat{A}}^{{\dagger}}}u_{x}\right)\right).$
It turns out that $\underline{u_{x}}=x$, $x\in\mathbb{R}^{2n}$, and since $V\left(u\right)=W\left(\underline{u}\right)$, the above relation can be written
$V\left(u\right)=\exp\left(2^{-1/2}\left(u^{\ast}\mathbf{\hat{A}-\mathbf{\hat{A}}^{{\dagger}}}u\right)\right),\;u\in\mathbb{C}^{n}.$
Now (3.31) in connection with (3.25) yields
$\exp\left(i\mathbf{\tilde{R}}x\right)$ $=\exp\left(2^{-1/2}\left(u_{x}^{\ast}U_{n}^{\ast}\mathbf{\hat{A}-\hat{A}}^{{\dagger}}U_{n}u_{x}\right)\right)$
$=V\left(U_{n}u_{x}\right).$
so that (1.18) implies
$\mathrm{Tr}\;\rho\exp\left(i\mathbf{\tilde{R}}\underline{u}\right)$ $=\exp\left(-\frac{1}{4}\left\langle U_{n}u,AU_{n}u\right\rangle\right)$
$=\exp\left(-\frac{1}{4}\left\langle u,U_{n}^{\ast}AU_{n}u\right\rangle\right)$
establishing (3.30).
3.2 Unbiased covariance estimation
Again assume that $n$ is odd. We will see that in the case of a Toeplitz symbol matrix $A$ (shift invariant time series), the set of observables $\hat{B}_{j}^{\ast}\hat{B}_{j}$, $\left|j\right|\leq\left(n-1\right)/2$ allows an unbiased estimator of the coefficients $a_{j}=a_{k,k+j}$, i.e. the analogs of the autocovariances of a classical time series (cf. (3.40) below).
For the vectors $\mathbf{u}_{j}=\left(u_{j,k}\right)_{k=1,\ldots,n},$ $j\in\mathbb{Z}$ given by (2.49) for $m=n$ we note
$u_{j,k}=n^{-1/2}\epsilon_{j}^{k-1}=n^{-1/2}\exp\left(2\pi ij\left(k-1\right)/n\right)=n^{-1/2}\exp\left(i\left(k-1\right)\omega_{j,n}\right)$
for the Fourier frequencies $\omega_{j,n}$ defined in (2.53). Using the Toeplitz property of $A_{n}=\left(a_{l-k}\right)_{k=1,\ldots,n}^{l=1,\ldots,n}$ we obtain for $\left|j\right|\leq\left(n-1\right)/2$
$\mathbf{u}_{j}^{\ast}A_{n}\mathbf{u}_{j}$ $=\sum_{k,l=1}^{n}\bar{u}_{j,k}\;u_{j,l}\;a_{l-k}=\sum_{k,l=1}^{n}a_{l-k}n^{-1}\exp\left(i\left(l-k\right)\omega_{j,n}\right)$
$=\sum_{s=-\left(n-1\right)}^{n-1}\frac{n-\left|s\right|}{n}a_{s}\exp\left(is\omega_{j,n}\right)=\sum_{s=-\left(n-1\right)}^{n-1}\left(1-\frac{\left|s\right|}{n}\right)a_{s}\phi_{s}\left(\omega_{j,n}\right),$ (3.32)
$\phi_{s}$ being defined by (2.52). Define a commuting set of observables
$\Pi_{j}=2\hat{B}_{j}^{\ast}\hat{B}_{j}+\mathbf{1,}\text{ }\left|j\right|\leq\left(n-1\right)/2.$ (3.33)
Then from Lemma 3.4 (i) and (3.32) we obtain
$\left\langle\Pi_{j}\right\rangle_{\rho}=\sum_{s=-\left(n-1\right)}^{n-1}\left(1-\frac{\left|s\right|}{n}\right)a_{s}\phi_{s}\left(\omega_{j,n}\right).$ (3.34)
Recalling the series representation (1.22) of the spectral density, we see that $\left\langle\Pi_{j}\right\rangle_{\rho}$ is an approximation to the spectral density at the Fourier frequency $\omega_{j,n}$. In particular, assuming that our quantum time series is $d$-dependent, i.e. $a_{j}=0$ for $\left|j\right|>d$, we have for sufficiently large $n$
$\left\langle\Pi_{j}\right\rangle_{\rho}=a\left(\omega_{j,n}\right)+O\left(n^{-1}\right),\text{ }\left|j\right|\leq\left(n-1\right)/2,$
i.e. the estimator $\Pi_{j}$ of $a\left(\omega_{j}\right)$ is asymptotically unbiased of order $O\left(n^{-1}\right)$. Furthermore from (3.34) we can obtain asymptotically unbiased estimates of the symbol coefficients $a_{j}$ (we may informally call them the covariances). Define vectors
$\mathbf{v}_{j,n}:=n^{-1/2}\left(\phi_{j}\left(\omega_{k,n}\right)\right)_{\left|k\right|\leq\left(n-1\right)/2}\text{, }j\in\mathbb{Z}.$ (3.35)
Then $\mathbf{v}_{j,n}$, $\left|j\right|\leq\left(n-1\right)/2$ is an orthonormal system, thus
$\mathbf{v}_{j,n}^{\ast}\mathbf{v}_{k,n}=\delta_{jk}\text{, }\left|j\right|,\left|k\right|\leq\left(n-1\right)/2.$ (3.36)
Indeed set $c_{k-j,n}:=\exp\left(i\left(k-j\right)\frac{2\pi}{n}\right)$; then it can be shown that
$c_{k-j,n}\mathbf{v}_{j,n}^{\ast}\mathbf{v}_{k,n}=c_{k-j,n}n^{-1}\sum_{\left|s\right|\leq\left(n-1\right)/2}\exp\left(i\left(k-j\right)\frac{2\pi s}{n}\right)=\mathbf{v}_{j,n}^{\ast}\mathbf{v}_{k,n}$
so that $\mathbf{v}_{j,n}^{\ast}\mathbf{v}_{k,n}$ must be zero unless $k=j$.
Define the vector of observables
$\mathbf{\Pi}_{n}=\left(\Pi_{j}\right)_{\left|j\right|\leq\left(n-1\right)/2};\;$ (3.37)
then (3.34) can be written, for $\varrho=\mathfrak{N}_{n}\left(0,A_{n}\right)$
$\left\langle\mathbf{\Pi}_{n}\right\rangle_{\rho}=n^{1/2}\sum_{j=-(n-1)}^{n-1}\left(1-\frac{\left|j\right|}{n}\right)a_{j}\mathbf{v}_{j,n}.$ (3.38)
At this point, by $d$-dependency for fixed $d$ and $n$ sufficiently large, we can assume that the above sum extends only over $\left|j\right|\leq d\leq\left(n-1\right)/2$. Then, defining the estimator
$\check{a}_{j,n}=\frac{n^{1/2}}{n-\left|j\right|}\mathbf{v}_{j,n}^{\ast}\mathbf{\Pi}_{n}\mathbf{,}\text{ for }\left|j\right|\leq d,$ (3.39)
we have by the orthogonality (3.36)
$E_{\rho}\check{a}_{j,n}=\frac{n^{1/2}}{n-\left|j\right|}\mathbf{v}_{j}^{\ast}\left\langle\mathbf{\Pi}_{n}\right\rangle_{\rho}=\frac{n^{1/2}}{n-\left|j\right|}n^{1/2}\left(1-\frac{\left|j\right|}{n}\right)\mathbf{v}_{j,n}^{\ast}\mathbf{v}_{j,n}a_{j}=a_{j}\text{.}$ (3.40)
The estimate $\check{a}_{j.n}$ is the analog of the basic unbiased covariance estimate in a classical time series (cf. [Shi19], Sec 6.4).
3.3 A preliminary estimator
3.3.1 Real parameters
We will take the unbiased estimator (3.39) as a starting point for constructing a preliminary estimator in the $d$-dependent case. Since our parameter vector $\left(a_{j}\right)_{\left|j\right|\leq d}$ is complex with $a_{-j}=\bar{a}_{j}$, we will transform it to a real vector as follows: $\theta=\left(\theta_{j}\right)_{\left|j\right|\leq d}$ where
$\theta_{0}=a_{0}\text{, }\theta_{j}=\sqrt{2}\operatorname{Re}a_{j}\text{, }\theta_{-j}=-\sqrt{2}\operatorname{Im}a_{j}\text{, }1\leq j\leq d.$ (3.41)
Let us also define a set of functions on $\left[-\pi,\pi\right]$ as
$\psi_{0}$ $=\phi_{j}=1\text{, }$ (3.42a)
$\psi_{j}$ $=\frac{1}{\sqrt{2}}\left(\phi_{j}+\phi_{-j}\right)=\sqrt{2}\cos\left(j\cdot\right),$ (3.42b)
$\psi_{-j}$ $=\frac{1}{i\sqrt{2}}\left(\phi_{j}-\phi_{-j}\right)=\sqrt{2}\sin\left(j\cdot\right),$ (3.42c)
for $j\in\mathbb{N}$. These functions fulfill
$\frac{1}{2\pi}\int_{\left[-\pi,\pi\right]}\psi_{j}\left(\omega\right)\psi_{l}\left(\omega\right)d\omega=\delta_{jl},\;j,l\in\mathbb{N}.$ (3.43)
Recalling (1.22), we can then write the spectral density as follows:
$a\left(\omega\right)$ $=\sum_{\left|j\right|\leq d}\phi_{j}\left(\omega\right)a_{j}$
$=a_{0}+\sum_{1\leq j\leq d}\left(\phi_{j}\left(\omega\right)+\phi_{-j}\left(\omega\right)\right)\operatorname{Re}a_{j}+i\sum_{1\leq j\leq d}\left(\phi_{j}\left(\omega\right)-\phi_{-j}\left(\omega\right)\right)\operatorname{Im}a_{j}$
$=\sum_{\left|j\right|\leq d}\psi_{j}\left(\omega\right)\theta_{j}=:a_{\theta}\left(\omega\right).$ (3.44)
The above defines the spectral density as a function $a_{\theta}$ of a parameter $\theta\in\mathbb{R}^{2d+1}$. The assumption $a\in\Theta_{2}\left(d,M\right)$ is then equivalent to
$\theta$ $\in\Theta_{2}^{\prime}\left(d,M\right):=\left\{\theta:\left\|\theta\right\|^{2}\leq M\right\}\cap\mathcal{L}_{M}^{\prime},$ (3.45)
$\mathcal{L}_{M}^{\prime}$ $:=\left\{\theta:\inf_{\omega\in\left[-\pi,\pi\right]}a_{\theta}\left(\omega\right)\geq 1+M^{-1}\right\}.$ (3.46)
This parameter space will often be written just $\Theta_{2}^{\prime}$, considering $d$ and $M$ fixed henceforth.
The next Lemma is an analog of Lemma 2.10.
Lemma 3.5
Suppose $\theta\in\Theta_{2}^{\prime}\left(d,M\right)$ for $M>1$. Then
$\left(1+M^{-1}\right)\;I\leq A_{n}\left(a_{\theta}\right)\leq\left(2d+1\right)^{1/2}M^{1/2}\;I.$ (3.47)
Proof. For $\omega\in\left[-\pi,\pi\right]$ we have
$a_{\theta}\left(\omega\right)=\sum_{\left|j\right|\leq d}\psi_{j}\left(\omega\right)\theta_{j}\leq\left(\sum_{\left|j\right|\leq d}\psi_{j}^{2}\left(\omega\right)\right)^{1/2}\left\|\theta\right\|\leq\left(2d+1\right)^{1/2}M^{1/2}.$
Set $C=\left(2d+1\right)^{1/2}M^{1/2}$; then analogously to the proof of Lemma 2.10 for every $x\in\mathbb{C}^{n}$ with $\left\|x\right\|=1$
$\left\langle x,A_{n}\left(a\right)x\right\rangle\leq\frac{C}{2\pi}\int_{-\pi}^{\pi}\left|\sum_{j=1}^{n}x_{j}\exp\left(ij\omega\right)\right|^{2}d\omega=C.$
Analogously we obtain from the first inequality in (2.10) $\left\langle x,A_{n}\left(a\right)x\right\rangle\geq\left(1+M^{-1}\right)$.
Define vectors, in analogy to $\mathbf{v}_{j}$ in (3.35),
$\mathbf{w}_{j,n}:=n^{-1/2}\left(\psi_{j}\left(\omega_{k}\right)\right)_{\left|k\right|\leq\left(n-1\right)/2}\text{, }\left|j\right|\leq\left(n-1\right)/2.$ (3.48)
We then have $\mathbf{w}_{0}=\mathbf{v}_{0}$ and
$\mathbf{w}_{j,n}=\frac{1}{\sqrt{2}}\left(\mathbf{v}_{j,n}+\mathbf{v}_{-j,n}\right)\text{, }\mathbf{w}_{-j,n}=\frac{1}{i\sqrt{2}}\left(\mathbf{v}_{j,n}-\mathbf{v}_{-j,n}\right)\text{, }1\leq j\leq\left(n-1\right)/2$
or equivalently
$\mathbf{v}_{j,n}=\frac{1}{\sqrt{2}}\left(\mathbf{w}_{j,n}+i\mathbf{w}_{-j,n}\right),\;\mathbf{v}_{-j,n}=\frac{1}{\sqrt{2}}\left(\mathbf{w}_{j,n}-i\mathbf{w}_{-j,n}\right).$ (3.49)
It follows that $\mathbf{w}_{j,n}$, $\left|j\right|\leq\left(n-1\right)/2$ are orthonormal; indeed they satisfy
$\mathbf{w}_{j,n}^{\prime}\mathbf{w}_{k,n}=\delta_{jk},\;\left|j\right|\leq\left(n-1\right)/2.$ (3.50)
Since $a_{j}\mathbf{v}_{j,n}+a_{-j}\mathbf{v}_{-j,n}=\theta_{j}\mathbf{w}_{j,n}+\theta_{-j}\mathbf{w}_{-j,n}$ for $0\leq j\leq d$, we can rewrite (3.38) under $d$-dependence as
$E_{\rho}\mathbf{\Pi}_{n}=n^{1/2}\sum_{j=-d}^{d}\left(1-\frac{\left|j\right|}{n}\right)\theta_{j}\mathbf{w}_{j,n}$ (3.51)
for $\rho=\mathfrak{N}_{n}\left(0,A_{n}\left(a_{\theta}\right)\right)$. Also the estimator (3.39) can be rewritten as
$\check{\theta}_{j,n}=\frac{n^{1/2}}{n-\left|j\right|}\mathbf{w}_{j,n}^{\prime}\mathbf{\Pi}_{n}\mathbf{,\;}\left|j\right|\leq d.$ (3.52)
Unbiasedness then follows from (3.51): for $\rho=\mathfrak{N}_{n}\left(0,A_{n}\left(a_{\theta}\right)\right)$
$E_{\rho}\check{\theta}_{j,n}=\theta_{j}\text{, }\left|j\right|\leq d\text{.}$ (3.53)
3.3.2 Partition into independent blocks
Recall that the $n$ pairs of operators $\left(\hat{A}_{j},\hat{A}_{j}^{\ast}\right)$, $j=1,\ldots,n$ define the $n$ modes of the quantum Gaussian state; we will subdivide this sequence into blocks as follows. Set
$m_{n}=2\left[\log n/2\right]+1,\;r_{n}=\left[n/\left(m_{n}+d\right)\right]$ (3.54)
so that $m_{n}$ is odd; we will write $m$ and $r$ hence forth. Consider sets of pairs
$S_{1}$ $:=\left\{\left(\hat{A}_{1},\hat{A}_{1}^{\ast}\right),\ldots,\left(\hat{A}_{m},\hat{A}_{m}^{\ast}\right)\right\},S_{2}:=\left\{\left(\hat{A}_{m+d+1},\hat{A}_{m+d+1}^{\ast}\right),\ldots,\left(\hat{A}_{2m+d},\hat{A}_{2m+d}^{\ast}\right)\right\},\ldots$
$S_{r}$ $:=\left\{\left(\hat{A}_{\left(r-1\right)\left(m+d\right)+1},\hat{A}_{\left(r-1\right)\left(m+d\right)+1}^{\ast}\right),\ldots,\left(\hat{A}_{rm+(r-1)d},\hat{A}_{rm+(r-1)d}^{\ast}\right)\right\}$
Note that operators from two different blocks $S_{j}$, $S_{h}$ are uncorrelated: considering e.g. the last pair $\left(\hat{A}_{m},\hat{A}_{m}^{\ast}\right)$ from $S_{1}$ and the first pair $\left(\hat{A}_{m+d+1},\hat{A}_{m+d+1}^{\ast}\right)$ from $S_{2}$, we have according to Lemma 3.1 (i)
$\left\langle\hat{A}_{m}^{\ast}\hat{A}_{m+d+1}\right\rangle_{\rho}=\frac{1}{2}a_{m,m+d+1}=\frac{1}{2}a_{d+1}=0$
in view of the $d$-dependence ($a_{h}=0$ for $\left|h\right|>d$). Similarly, applying (1.16)
$\left\langle\hat{A}_{m}\hat{A}_{m+d+1}\right\rangle_{\rho}$ $=\frac{1}{2}\left\langle\left(Q_{m}+iP_{m}\right)\left(Q_{m+d+1}+iP_{m+d+1}\right)\right\rangle_{\rho}$
$=\frac{1}{2}\left(\left\langle Q_{m}Q_{m+d+1}\right\rangle_{\rho}+i\left\langle Q_{m}P_{m+d+1}\right\rangle_{\rho}+i\left\langle P_{m}Q_{m+d+1}\right\rangle_{\rho}-\left\langle P_{m}P_{m+d+1}\right\rangle_{\rho}\right)$
$=\frac{1}{2}\left(\operatorname{Re}a_{m,m+d+1}-i\operatorname{Im}a_{m,m+d+1}+i\operatorname{Im}a_{m,m+d+1}-\operatorname{Re}a_{m,m+d+1}\right)=0.$
Intuitively, when we ”omit” all pairs $\left(\hat{A}_{j},\hat{A}_{j}^{\ast}\right)$ between the blocks, and also those after the last block $S_{r}$, then, because of the $d$-dependence, the remaining blocks $S_{1},\ldots,S_{r}$ should be ”independent”. To make this rigorous in the quantum context, we take a partial trace of the state $\mathfrak{N}_{n}\left(0,A_{n}\right)$, tracing out all the modes corresponding to the pairs $\left(\hat{A}_{j},\hat{A}_{j}^{\ast}\right)$ in question. What we get is a Gaussian state with $rm$ modes and symbol matrix $I_{r}\otimes A_{\left(m\right)}$ (in view of the Toeplitz form of $A_{n}$, where $A_{\left(m\right)}$ is the upper central $m\times m$ submatrix of $A_{n}$, i.e. we obtain the gauge invariant state $\mathfrak{N}_{rm}\left(0,I_{r}\otimes A_{\left(m\right)}\right)$. The details of this reasoning are given in Subsection A.2.1. Using characteristic functions , it is easy to show that this state is equivalent to an $r$-fold tensor product $\left(\mathfrak{N}_{m}\left(0,A_{\left(m\right)}\right)\right)^{\otimes r}$.
Recall the basic model assumption (1.23), i.e. $A_{n}=A_{n}\left(a\right)$, $n\rightarrow\infty$ for a given spectral density $a$ (with current assumption $a=a_{\theta}$, $\theta\in\Theta_{2}^{\prime}$, cf. (3.45)). It follows that $A_{\left(m\right)}=A_{m}\left(a_{\theta}\right)$, or $A_{\left(m\right)}=A_{m}$ for short, and we now have the parametric model of states $\left(\mathfrak{N}_{m}\left(0,A_{m}\left(a_{\theta}\right)\right)\right)^{\otimes r}$, $\theta\in\Theta_{2}^{\prime}$.
For each of the $r$ component states of $\left(\mathfrak{N}_{m}\left(0,A_{m}\left(a_{\theta}\right)\right)\right)^{\otimes r}$, we now form the vector of observables $\mathbf{\Pi}_{m}$ corresponding to (3.37) for $n=m$, obtaining an $r$-tuple of such vectors $\mathbf{\Pi}_{m,j}$, $j=1,\ldots,r$, and we form the average
$\mathbf{\bar{\Pi}}_{n}:=r^{-1}\sum_{j=1}^{r}\mathbf{\Pi}_{m,j}.$ (3.55)
We will modify the estimator (3.52), essentially substituting $\mathbf{\bar{\Pi}}_{n}$ for $\mathbf{\Pi}_{n}$. To write it in vector form, consider the vectors $\mathbf{w}_{j,n}$ of (3.48) for dimension $n=m$ and define the $m\times\left(2d+1\right)$ real matrix
$W_{m}=\left(\mathbf{w}_{-d,m},\ldots,\mathbf{w}_{0},\ldots,\mathbf{w}_{d,m}\right),$ (3.56)
fulfilling $W_{m}^{\prime}W_{m}=I_{2d+1}$ by (3.50). Furthermore define the diagonal $\left(2d+1\right)\times\left(2d+1\right)$ matrix
$F_{m}:=\mathrm{diag}\left(\frac{m}{m-\left|j\right|}\right)_{\left|j\right|\leq d}.$ (3.57)
Definition 3.6
The preliminary estimator of the parameter vector $\theta$ from (3.41) is
$\hat{\theta}_{n}:=m^{-1/2}F_{m}W_{m}^{\prime}\mathbf{\bar{\Pi}}_{n}$ (3.58)
with $\mathbf{\bar{\Pi}}_{n}$ from (3.55)
Since $E_{\rho}\mathbf{\Pi}_{m,j}$ coincides with $E_{\rho}\mathbf{\Pi}$ (cf. (3.51)) if the latter is taken at dimension $n=m$, from (3.53) we immediately obtain unbiasedness: $E_{\rho}\hat{\theta}_{n}=\theta$.
Let $P_{n,\theta}$ be the joint distribution of the $\mathbb{R}^{m}$-valued random vectors $\mathbf{\Pi}_{m,j}$, $j=1,\ldots,r$ from (3.55) under the state $\rho=\mathfrak{N}_{n}\left(0,A_{n}\left(a_{\theta}\right)\right)$. Here $\mathbf{\bar{\Pi}}_{n}$ will function as the basic observable for asymptotic inference about $\theta$, so that distributions of further random variables in this section can be described in terms of $P_{n,\theta}$ and corresponding expectations $E_{n,\theta}$.
3.3.3 Asymptotic covariance matrix
We have
$n^{1/2}\left(\hat{\theta}_{n}-\theta\right)=\sum_{j=1}^{r}r^{-1}n^{1/2}m^{-1/2}F_{m}W_{m}^{\prime}\left(\mathbf{\Pi}_{m,j}-E_{\rho}\mathbf{\Pi}_{m,j}\right)$ (3.59)
where it follows from (3.51) that
$E_{n,\theta}\mathbf{\Pi}_{m,j}=m^{1/2}W_{m}F_{m}^{-1}\theta\mathbf{.}$ (3.60)
[…]
Lemma 3.7
Under $\rho=\mathfrak{N}_{n}\left(0,A_{n}\left(a_{\theta}\right)\right)$, $\theta\in\Theta_{2}^{\prime}$ we have
$\lim_{n\rightarrow\infty}\mathrm{Cov}_{n,\theta}\left(n^{1/2}\left(\hat{\theta}_{n}-\theta\right)\right)=\Phi_{\theta}^{0}:=\left(\Phi_{\theta,jk}^{0}\right)_{\left|j\right|,\left|k\right|\leq d},\text{ }$
where
$\Phi_{\theta,jk}^{0}=\frac{1}{2\pi}\int_{\left(-\pi,\pi\right)}\left(a_{\theta}^{2}\left(\omega\right)-1\right)\psi_{j}\left(\omega\right)\psi_{k}\left(\omega\right)d\omega,$ (3.61)
$a_{\theta}\left(\omega\right)$ is the spectral density depending on $\theta\in\Theta_{2}^{\prime}$ according to (3.44), and functions $\psi_{h}$ are defined by (3.42). The convergence is uniform over $\theta\in\Theta_{2}^{\prime}$.
Proof. Note that in (3.59) we have $r^{-1}n^{1/2}m^{-1/2}\sim r^{-1/2}$ and $F_{m}\longrightarrow I_{2d+1}$, hence writing $\mathbf{\Pi}_{m}=\mathbf{\Pi}_{1,m}$ we obtain
$\mathrm{Cov}_{n,\theta}\left(n^{1/2}\left(\hat{\theta}_{n}-\theta\right)\right)=W_{m}^{\prime}\mathrm{Cov}_{n,\theta}\left(\mathbf{\Pi}_{m}\right)W_{m}\left(1+o\left(1\right)\right).$ (3.62)
To obtain the covariance matrix appearing on the r.h.s., consider the result of Lemma 3.4 for $n=m$. For a $m\times m$ matrix $M=\left(M_{jl}\right)_{j,l=1}^{m}$, define the real matrix
$M^{\left[2\right]}=\left(\left|M_{jl}\right|^{2}\right)_{j,l=1}^{m}.$ (3.63)
.Then the result of 3.4 (ii) can be written, with $A_{m}=A_{m}\left(a_{\theta}\right)$,
$\mathrm{Cov}_{n,\theta}\left(\hat{B}_{1}^{\ast}\hat{B}_{1},\ldots,\hat{B}_{m}^{\ast}\hat{B}_{m}\right)=\frac{1}{4}\left(\left(U_{m}^{\ast}A_{m}U_{m}\right)^{\left[2\right]}-I_{m}\right).$
Now recall the definition of the observable vector $\mathbf{\Pi}_{m}$ in (3.33), (3.37) and identify $\mathbf{\Pi}_{1,m}$ with $\Pi_{m}$. We obtain
$\mathrm{Cov}_{n,\theta}\left(\mathbf{\Pi}_{1,m}\right)=\left(U_{m}^{\ast}A_{m}U_{m}\right)^{\left[2\right]}-I_{m},$ (3.64)
with $U_{m}$ from (2.55) and $m$ from (3.54). Then (3.62) can be written
$\mathrm{Cov}_{n,\theta}\left(n^{1/2}\left(\hat{\theta}_{n}-\theta\right)\right)\sim W_{m}^{\prime}\left(U_{m}^{\ast}A_{m}U_{m}\right)^{\left[2\right]}W_{m}-I_{2d+1}.$ (3.65)
To treat the first term on the r.h.s. of (3.65), recall that the Hilbert-Schmidt norm $\left\|M\right\|_{2}$ of an $m\times m$ matrix $M$ is defined as $\left\|M\right\|_{2}^{2}=\mathrm{Tr}\;M^{\ast}M=\sum_{j,l}\left|M_{jk}\right|^{2}$. Note that, under $d$-dependence, the symbol matrix $A_{m}$ is banded in the terminology of [Gra06]. Then for $A_{m}=A_{m}\left(a_{\theta}\right)$ and its circulant approximation $\tilde{A}_{m}=\tilde{A}_{m}\left(a_{\theta}\right)$ defined in (2.58) we have for $\theta\in\Theta_{2}^{\prime}$
$m^{-1}\left\|A_{m}-\tilde{A}_{m}\right\|_{2}^{2}=m^{-1}2\sum_{j=1}^{d}j\left|a_{j}\right|^{2}=m^{-1}\sum_{j=-d}^{d}j\theta_{j}^{2}\rightarrow 0\text{ as }m\rightarrow\infty;$
by a reasoning similar to (2.63) when $m=n$ and $a_{j}=0$ for $j>d$ (or referring to Lemma 4.2 in [Gra06]). The convergence is uniform over $\left\|\theta\right\|\leq C$, hence over $\theta\in\Theta_{2}^{\prime}$. Let $\tilde{A}_{m}=U_{m}\tilde{\Lambda}_{m}U_{m}^{\ast}$ be the spectral decomposition of $\tilde{A}_{m}$; then according to (2.61) we have for sufficiently large $m$ (such that $m>2d+1$),
$\tilde{\Lambda}_{m}=\Lambda_{m}:=\mathrm{diag}\left(\;a_{\theta}\left(\omega_{j,m}\right)_{\left|j\right|\leq\left(m-1\right)/2}\right)$ (3.66)
where $\omega_{j,m}$ are the Fourier frequencies $\omega_{j,m}=2\pi j/m$, $\left|j\right|\leq\left(m-1\right)/2$. Since $\left\|M\right\|_{2}^{2}=\left\|U_{m}^{\ast}MU_{m}\right\|_{2}^{2}$ for any $m\times m$ matrix $M$, we obtain
$m^{-1}\left\|U_{m}^{\ast}A_{m}U_{m}-\Lambda_{m}\right\|_{2}^{2}\rightarrow 0\text{ as }n\rightarrow\infty.$ (3.67)
uniformly over $\theta\in\Theta_{2}^{\prime}$. Consider the elelement with index $\left(j,k\right)$ of $W_{m}^{\prime}\left(U_{m}^{\ast}A_{m}U_{m}\right)^{\left[2\right]}W_{m}$; this is
$\mathbf{w}_{j,m}^{\prime}\left(U_{m}^{\ast}A_{m}U_{m}\right)^{\left[2\right]}\mathbf{w}_{k,m}$ $=\mathbf{w}_{j,m}^{\prime}\Lambda_{m}^{2}\mathbf{w}_{k,m}+\mathbf{w}_{j,m}^{\prime}D_{m}\mathbf{w}_{k,m}\text{ where }$
$D_{m}$ $:=\left(U_{m}^{\ast}A_{m}U_{m}\right)^{\left[2\right]}-\Lambda_{m}^{2}.$ (3.68)
Note that, since all components of $\mathbf{w}_{j,m}$ and $\mathbf{w}_{k,m}$ are bounded in modulus by $\sqrt{2}m^{-1/2}$, we have
$\left|\mathbf{w}_{j,m}^{\ast}D_{m}\mathbf{w}_{k,m}\right|\leq 2m^{-1}\sum_{s,t=1}^{m}\left|\left(D_{m}\right)_{st}\right|=2m^{-1}\sum_{s,t=1}^{m}\left|\left(U_{m}^{\ast}A_{m}U_{m}\right)_{st}^{\left[2\right]}-\left(\Lambda_{m}^{2}\right)_{st}\right|.$
Note that for any complex $x,y$
$\left|\left|x\right|^{2}-\left|y\right|^{2}\right|$ $=\left|\left(\left|x\right|-\left|y\right|\right)\left(\left|x\right|+\left|y\right|\right)\right|$
$\leq\left|x-y\right|\;\left(\left|x\right|+\left|y\right|\right).$
Applying this bound to each term $\left|\left(U_{m}^{\ast}A_{m}U_{m}\right)_{st}^{\left[2\right]}-\left(\Lambda_{m}^{2}\right)_{st}\right|$, we obtain,
$m^{-1}\sum_{s,t=1}^{m}\left|\left(U_{m}^{\ast}A_{m}U_{m}\right)_{st}^{\left[2\right]}-\left(\Lambda_{m}^{2}\right)_{st}\right|$
$\leq m^{-1}\sum_{s,t=1}^{m}\left|\left(U_{m}^{\ast}A_{m}U_{m}-\Lambda_{m}\right)_{st}\right|\left(\left|\left(U_{m}^{\ast}A_{m}U_{m}\right)_{st}\right|+\left|\left(\Lambda_{m}\right)_{st}\right|\right).$
Applying the Cauchy-Schwartz inequality, we obtain an upper bound
$\left(m^{-1}\left\|U_{m}^{\ast}A_{m}U_{m}-\Lambda_{m}\right\|_{2}^{2}\right)^{1/2}\left(2m^{-1}\left(\left\|U_{m}^{\ast}A_{m}U_{m}\right\|^{2}+\left\|\Lambda_{m}\right\|_{2}^{2}\right)\right)^{1/2}.$ (3.69)
Here the first factor is $o\left(1\right)$ uniformly over $\theta\in\Theta_{2}^{\prime}$ by (3.67). The second factor is bounded by the following reasoning. In view of $d$-dependence
$m^{-1}\left\|U_{m}^{\ast}A_{m}U_{m}\right\|_{2}^{2}=m^{-1}\left\|A_{m}\right\|_{2}^{2}=\sum_{\left|j\right|\leq d}\frac{m-\left|j\right|}{m}\left|a_{j}\right|^{2}=\sum_{\left|j\right|\leq d}\frac{m-\left|j\right|}{m}\theta_{j}^{2}\leq M$
by $\theta\in\Theta_{2}^{\prime}$. Similarly
$m^{-1}\left\|\Lambda_{m}\right\|_{2}^{2}$ $\leq\max_{-\pi\leq\omega\leq\pi}a_{\theta}^{2}\left(\omega\right)=\max_{-\pi\leq\omega\leq\pi}\left(\sum_{\left|j\right|\leq d}\psi_{j}\left(\omega\right)\theta_{j}\right)^{2}$
$\leq\left(2d+1\right)\left\|\theta\right\|^{2}\leq\left(2d+1\right)M$
by $\theta\in\Theta_{2}^{\prime}$. As a consequence, (3.69) is $o\left(1\right)$ uniformly over $\theta\in\Theta_{2}^{\prime}$, which implies
$\mathbf{w}_{j,m}^{\prime}\left(U_{m}^{\ast}A_{m}U_{m}\right)^{\left[2\right]}\mathbf{w}_{k,m}$ $=\mathbf{w}_{j,m}^{\prime}\Lambda_{m}^{2}\mathbf{w}_{k,m}+o\left(1\right)$ (3.70)
$=m^{-1}\sum_{\left|j\right|\leq\left(m-1\right)/2}a_{\theta}^{2}\left(\omega_{j,m}\right)\psi_{j}\left(\omega_{j,m}\right)\psi_{k}\left(\omega_{j,m}\right)+o\left(1\right).$
Since the set of functions $\left\{a_{\theta},\;\theta\in\Theta_{2}^{\prime},\psi_{j},\;\left|j\right|\leq d\right\}$ is uniformly bounded and Lipschitz, we now have
$\mathbf{w}_{j,m}^{\prime}\left(U_{m}^{\ast}A_{m}U_{m}\right)^{\left[2\right]}\mathbf{w}_{k,m}=\frac{1}{2\pi}\int_{\left(-\pi,\pi\right)}a_{\theta}^{2}\left(\omega\right)\psi_{j}\left(\omega\right)\psi_{k}\left(\omega\right)d\omega+o\left(1\right).$
uniformly over $\theta\in\Theta_{2}^{\prime}$. In view of (3.65) and (3.43), the claim follows.
Lemma 3.8
Let $\gamma_{n}\rightarrow\infty$ be a sequence such that $\gamma_{n}=o\left(n^{1/2}\right)$. Then for every $\varepsilon>0$ we have
$\sup_{\theta\in\Theta_{2}^{\prime}}P_{n,\theta}\left(\gamma_{n}\left\|\hat{\theta}_{n}-\theta\right\|\geq\varepsilon\right)\rightarrow 0.$
Proof. We have
$P_{n,\theta}\left(\gamma_{n}\left\|\hat{\theta}_{n}-\theta\right\|\geq\varepsilon\right)\leq\frac{\gamma_{n}^{2}}{n}\frac{E_{n,\theta}n\left\|\hat{\theta}_{n}-\theta\right\|^{2}}{\varepsilon^{2}}=\frac{\gamma_{n}^{2}}{n}\frac{\mathrm{Tr\;Cov}_{\rho}\left(n^{1/2}\left(\hat{\theta}_{n}-\theta\right)\right)}{\varepsilon^{2}}$
so the claim follows from Lemma 3.7.
3.4 A one-step improvement estimator
The estimator $\hat{\theta}_{n}$ can be shown to be asymptotically normal, but it is not optimal; indeed will turn out that the optimal covariance matrix is not $\Phi_{\theta}^{0}$ but the inverse of the matrix
$\Phi_{\theta}$ $=\left(\Phi_{\theta,jk}\right)_{\left|j\right|,\left|k\right|\leq d},$ (3.71)
$\Phi_{\theta,jk}$ $:=\frac{1}{2\pi}\int_{\left(-\pi,\pi\right)}\left(a_{\theta}^{2}\left(\omega\right)-1\right)^{-1}\psi_{j}\left(\omega\right)\psi_{k}\left(\omega\right)d\omega.$ (3.72)
where $a_{\theta}\left(\omega\right)$ is the spectral density depending on $\theta\in\Theta_{2}^{\prime}$ according to (3.44).
Lemma 3.9
There are constants $0<C_{1,M}<C_{2,M}$ depending only on $M$ and $d$ such that for all $\theta\in\Theta_{2}^{\prime}.$
$C_{1,M}\;\leq\lambda_{\min}\left(\Phi_{\theta}\right)\text{,\ \ \ }\lambda_{\max}\left(\Phi_{\theta}\right)\leq C_{2,M}\text{.}$
Proof. Note that in view of $a_{\theta}\left(\omega\right)\geq 1+M^{-1}$ we have
$a_{\theta}^{2}\left(\omega\right)-1\geq\left(1+M^{-1}\right)^{2}-1\geq\left(1+M^{-1}\right)-1\geq M^{-1}.$
Furthermore
$a_{\theta}^{2}\left(\omega\right)=\left(\sum_{\left|j\right|\leq d}\psi_{j}\left(\omega\right)\theta_{j}\right)^{2}\leq\left\|\theta\right\|^{2}\sum_{\left|j\right|\leq d}\psi_{j}^{2}\left(\omega\right)=\left\|\theta\right\|^{2}\left(2d+1\right)\leq M\left(2d+1\right).$ (3.73)
The last two displays imply
$\left(M\left(2d+1\right)\right)^{-1}\leq\left(a_{\theta}^{2}\left(\omega\right)-1\right)^{-1}\leq M.$ (3.74)
Now for $x=\left(x_{j}\right)_{\left|j\right|\leq d}\in\mathbb{R}^{2d+1}$ we have
$x^{\prime}\Phi_{\theta}x$ $=\frac{1}{2\pi}\int_{\left(-\pi,\pi\right)}\left(a_{\theta}^{2}\left(\omega\right)-1\right)^{-1}\left(\sum_{\left|j\right|\leq d}x_{j}\psi_{j}\left(\omega\right)\right)^{2}d\omega$
$\leq M\frac{1}{2\pi}\int_{\left(-\pi,\pi\right)}\left(\sum_{\left|j\right|\leq d}x_{j}\psi_{j}\left(\omega\right)\right)^{2}d\omega=M$
and the bound $x^{\prime}\Phi_{\theta}x\geq\left(2M\left(d+1\right)\right)^{-1}$ follows analogously. Setting $C_{1,M}=\left(M\left(2d+1\right)\right)^{-1}$, $C_{2,M}=M$ completes the proof.
In order to modify the preliminary estimator $\hat{\theta}_{n}$ given by (3.58) in a suitable way, we will need estimates of the parameter dependent diagonal matrices
$\Delta_{m,\theta}:=\mathrm{diag}\left(\;a_{\theta}^{2}\left(\omega_{j,m}\right)_{\left|j\right|\leq\left(m-1\right)/2}-1\right).$ (3.75)
In order to replace $\theta$ there by a suitable estimator, consider the following lemma.
Lemma 3.10
The set $\Theta_{2}^{\prime}=\Theta_{2}^{\prime}\left(M,d\right)$ given by (3.45) is a compact convex subset of $\mathbb{R}^{2d+1}$.
Proof. The set $B_{M}:=\left\{\theta\in\mathbb{R}^{2d+1}:\left\|\theta\right\|^{2}\leq M,\;\right\}$ is compact and convex. for each $\omega$, the set $\left\{\theta\in\mathbb{R}^{2d+1}:a_{\theta}\left(\omega\right)\geq 1+M^{-1}\right\}$ is convex and closed, since the map $\theta\rightarrow a_{\theta}\left(\omega\right)$ is linear. Since the intersection of closed sets is closed, and convex if each set is convex, $\Theta_{2}^{\prime}\left(M,d\right)$ is a closed convex subset of $B_{M}$, from which the claim follows.
Define an estimator $\bar{\theta}_{n}$ as the projection of $\hat{\theta}_{n}$ onto the compact convex set $\Theta_{2}^{\prime}$ and set
$\hat{\Delta}_{m}:=\Delta_{m,\bar{\theta}_{n}}=\mathrm{diag}\left(\;a_{\bar{\theta}_{n}}^{2}\left(\omega_{j,m}\right)-1\right)_{\left|j\right|\leq\left(m-1\right)/2}.$
Note that do to (3.74), $\hat{\Delta}_{m}$ is nonsingular.
Definition 3.11
The improved estimator of the parameter vector $\theta$ from (3.41) is
$\tilde{\theta}_{n}:=m^{-1/2}F_{m}\left(W_{m}^{\prime}\hat{\Delta}_{m}^{-1}W_{m}\right)^{-1}W_{m}^{\prime}\hat{\Delta}_{m}^{-1}\mathbf{\bar{\Pi}}_{n}$ (3.76)
with $\mathbf{\bar{\Pi}}_{n}$ from (3.55).
Refer to Appendix A.3 for the definition of convergence in distribution (with symbol $\Longrightarrow_{d}$) uniformly in $\theta$, and some associated results.
Theorem 3.12
The estimator $\tilde{\theta}_{n}$ is asymptotically normal
$n^{1/2}\left(\tilde{\theta}_{n}-\theta\right)\Longrightarrow_{d}N_{2d+1}\left(0,\Phi_{\theta}^{-1}\right)$
uniformly in $\theta\in\Theta_{2}^{\prime}$.
We will begin the proof with a series of technical lemmas.
Lemma 3.13
On the compact set $\Theta_{2}^{\prime}\in\mathbb{R}^{2d+1}$, the map $\theta\longrightarrow\Phi_{\theta}^{-1}$ is continuous in Hilbert-Schmidt norm.
Proof. For $x\in\mathbb{R}^{2d+1}$ with $\left\|x\right\|=1$, we have in view of (3.74), setting $M_{1}=$
$x^{\prime}\Phi_{\theta}x$ $=\frac{1}{2\pi}\int_{\left(-\pi,\pi\right)}\left(a_{\theta}^{2}\left(\omega\right)-1\right)^{-1}\left(\sum_{j=-d}^{d}x_{j}\psi_{j}\left(\omega\right)\right)^{2}d\omega$
$\leq M\frac{1}{2\pi}\int_{\left(-\pi,\pi\right)}\left(\sum_{j=-d}^{d}x_{j}\psi_{j}\left(\omega\right)\right)^{2}d\omega=M$
and similarly, for $M_{1}=$ $\left(M\left(2d+1\right)\right)^{-1}$
$x^{\prime}\Phi_{\theta}x\geq M_{1}.$
It follows that
$s_{1}$ $:=\inf\left\{\lambda_{\min}\left(\Phi_{\theta}\right):\theta\in\Theta_{2}^{\prime}\right\}\geq M_{1}>0.$
$s_{2}$ $:=\sup\left\{\lambda_{\max}\left(\Phi_{\theta}\right):\theta\in\Theta_{2}^{\prime}\right\}\leq M.$
Clearly the map $\theta\longrightarrow\Phi_{\theta}$ is continuous on $\Theta_{2}^{\prime}$. For nonsingular matrices $\Phi_{1},\Phi_{2}$ we have
$\Phi_{1}^{-1}-\Phi_{2}^{-1}=\Phi_{1}^{-1}\left(\Phi_{2}-\Phi_{1}\right)\Phi_{2}^{-1}$
which for the Hilbert-Schmidt norm $\left\|\cdot\right\|_{2}$ implies, if both $\Phi_{1},\Phi_{2}$ are positive,
$\left\|\Phi_{1}^{-1}-\Phi_{2}^{-1}\right\|_{2}\leq\lambda_{\max}\left(\Phi_{1}^{-1}\right)\lambda_{\max}\left(\Phi_{2}^{-1}\right)\left\|\Phi_{2}-\Phi_{2}\right\|_{2}.$
Thus for , $\theta_{j}\in\Theta_{2}^{\prime}$, $j=1,2$ we have
$\left\|\Phi_{\theta_{1}}^{-1}-\Phi_{\theta_{2}}^{-1}\right\|_{2}\leq s_{1}^{-2}\left\|\Phi_{\theta_{1}}-\Phi_{\theta_{2}}\right\|_{2}$
showing that the map $\theta\longrightarrow\Phi_{\theta}^{-1}$ is continuous on $\Theta_{2}^{\prime}$.
Define the function
$g\left(\theta,\omega\right):=\left(a_{\theta}^{2}\left(\omega\right)-1\right)^{-1}\text{, }\theta\in\Theta_{2}^{\prime},\omega\in\left[-\pi,\pi\right].$
Lemma 3.14
There exists $L>0$ depending only on $m$ and $d$ such that
$\sup_{\omega\in\left[-\pi,\pi\right]}\left|g\left(\theta_{1},\omega\right)-g\left(\theta_{2},\omega\right)\right|\leq L\left\|\theta_{1}-\theta_{2}\right\|\text{, }\theta_{1},\theta_{2}\in\Theta_{2}^{\prime}.$
Proof. We first claim that $\left\|\partial_{\theta}g\left(\theta,\omega\right)\right\|^{2}\leq 2M^{2}.$Indeed for $\theta=\left(\theta_{j}\right)_{\left|j\right|\leq d}$ we have for any $\left|j\right|\leq d$, recalling $a_{\theta}\left(\omega\right)=\sum_{\left|j\right|\leq d}\theta_{j}\psi_{j}\left(\omega\right)$, where we used the bounds (3.74) and (3.73). Consequently
$\left\|\partial_{\theta}g\left(\theta,\omega\right)\right\|^{2}$ $=\sum_{\left|j\right|\leq d}\left(\partial_{\theta_{j}}g\left(\theta,\omega\right)\right)^{2}\leq\left(2d+1\right)M^{5}\sum_{\left|j\right|\leq d}\psi_{j}^{2}\left(\omega\right)$
$=M^{5}(2d+1)^{2}.$
Noting also that $\partial_{\theta}g\left(\theta,\omega\right)$ is continuous in $\theta$, the claim follows.
Proof of Theorem 3.12. Step 1. Lemma 3.13 in conjunction with Lemma A.7 shows that the mapping $\theta\rightarrow N_{2m+1}\left(0,\Phi_{\theta}^{-1}\right)$ is continuous in total variation norm on the compact $\Theta_{2}^{\prime}$. According to Lemma A.5 (iii), it suffices to prove that for every sequence $\left\{\theta_{n}\right\}$ such that $\theta_{n}\rightarrow\theta$ for some $\theta\in\Theta$, one has
$n^{1/2}\left(\tilde{\theta}_{n}-\theta_{n}\right)\Longrightarrow_{d}N_{2d+1}\left(0,\Phi_{\theta}^{-1}\right)\text{ under }P_{n,\theta_{n}}.$
From (3.60) we obtain
$E_{n,\theta}\mathbf{\bar{\Pi}}_{n}=m^{1/2}W_{m}F_{m}^{-1}\theta$
and hence
$\theta$ $=m^{-1/2}F_{m}\left(W_{m}^{\prime}\hat{\Delta}_{m}^{-1}W_{m}\right)^{-1}W_{m}^{\prime}\hat{\Delta}_{m}^{-1}\mathbf{\;}E_{n,\theta}\mathbf{\bar{\Pi}}_{n},$
$n^{1/2}\left(\tilde{\theta}_{n}-\theta\right)$ $=n^{1/2}\left(mr\right)^{-1/2}F_{m}\left(W_{m}^{\prime}\hat{\Delta}_{m}^{-1}W_{m}\right)^{-1}W_{m}^{\prime}\hat{\Delta}_{m}^{-1}r^{1/2}\left(\mathbf{\bar{\Pi}}_{n}-E_{n,\theta}\mathbf{\bar{\Pi}}_{n}\right).$
Here $n^{1/2}\left(mr\right)^{-1/2}=1+o\left(1\right)$ due to (3.54) and $F_{m}\rightarrow I_{2d+1}$ due to (3.57). Hence it suffices to prove that for all sequences $\theta_{n}$ converging to some $\theta$
$\left(W_{m}^{\prime}\hat{\Delta}_{m}^{-1}W_{m}\right)^{-1}W_{m}^{\prime}\hat{\Delta}_{m}^{-1}r^{1/2}\left(\mathbf{\bar{\Pi}}_{n}-E_{n,\theta_{n}}\mathbf{\bar{\Pi}}_{n}\right)\Longrightarrow_{d}N_{2d+1}\left(0,\Phi_{\theta}^{-1}\right)\text{ under }P_{n,\theta_{n}}.$ (3.77)
The sequence $\left\{\theta_{n}\right\}\subset\Theta_{2}^{\prime}$ will be considered fixed henceforth and $P_{n,\theta_{n}}$ is assumed to be joint distribution of the $\mathbb{R}^{m}$-valued random vectors $\mathbf{\Pi}_{j,m}$, $j=1,\ldots,r$ from (3.55).
Step 2. We claim
$\left(W_{m}^{\prime}\hat{\Delta}_{m}^{-1}W_{m}\right)^{-1}\rightarrow_{p}\Phi_{\theta}^{-1}$ (3.78)
(convergence in probability of the $\left(2d+1\right)\times\left(2d+1\right)$ matrix). Note that
$\left\|W_{m}^{\prime}\hat{\Delta}_{m}^{-1}W_{m}-W_{m}^{\prime}\Delta_{m,\theta_{n}}^{-1}W_{m}\right\|_{2}^{2}\leq\left\|\hat{\Delta}_{m}^{-1}-\Delta_{m,\theta_{n}}^{-1}\right\|_{2}^{2}$
$\leq m\sup_{\omega\in\left[-\pi,\pi\right]}\left(\left(a_{\bar{\theta}_{n}}^{2}\left(\omega\right)-1\right)^{-1}-\left(a_{\theta_{n}}^{2}\left(\omega\right)-1\right)^{-1}\right)^{2}$ (3.79)
[…]
$\leq m\;L^{2}\left\|\hat{\theta}_{n}-\theta_{n}\right\|^{2}\text{ (projection property of }\bar{\theta}_{n}\text{) }$
$\rightarrow_{p}0$ (3.80)
where the last claim follows from Lemma 3.8) and $m\sim\log n=o\left(n\right)$. Furthermore note that for each element $(j,k)$ of $W_{m}^{\prime}\Delta_{m,\theta_{n}}^{-1}W_{m}$ we have
$\mathbf{w}_{j,m}^{\prime}\Delta_{m,\theta_{n}}^{-1}\mathbf{w}_{k,m}$ $=m^{-1}\sum_{s\leq\left(m-1\right)/2}\left(a_{\theta_{n}}^{2}\left(\omega_{s,m}\right)-1\right)^{-1}\psi_{j}\left(\omega_{s,m}\right)\psi_{k}\left(\omega_{s,m}\right)$
$=\frac{1}{2\pi}\int_{\left(-\pi,\pi\right)}\left(a_{\theta}^{2}\left(\omega\right)-1\right)^{-1}\psi_{j}\left(\omega\right)\psi_{k}\left(\omega\right)d\omega+o\left(1\right).$
where the convergence to the integral follows from Lemma 3.14 and $\theta_{n}\rightarrow\theta$. Hence by (3.72)
$\mathbf{w}_{j,m}^{\prime}\Delta_{m,\theta_{n}}^{-1}\mathbf{w}_{k,m}=\Phi_{\theta,jk}+o\left(1\right).$
The last relation and (3.80) imply (3.78). For (3.77) it now suffices to prove
$W_{m}^{\prime}\hat{\Delta}_{m}^{-1}r^{1/2}\left(\mathbf{\bar{\Pi}}_{n}-E_{n,\theta_{n}}\mathbf{\bar{\Pi}}_{n}\right)\Longrightarrow_{d}N_{2d+1}\left(0,\Phi_{\theta}\right).\text{ }$ (3.81)
Step 3. We claim
$W_{m}^{\prime}\left(\hat{\Delta}_{m}^{-1}-\Delta_{m,\theta_{n}}^{-1}\right)r^{1/2}\left(\mathbf{\bar{\Pi}}_{n}-E_{n,\theta_{n}}\mathbf{\bar{\Pi}}_{n}\right)\rightarrow_{p}0$ (3.82)
(convergence in probability of a $2d+1$-vector). Indeed we have
$\left\|W_{m}^{\prime}\left(\hat{\Delta}_{m}^{-1}-\Delta_{m,\theta_{n}}^{-1}\right)r^{1/2}\left(\mathbf{\bar{\Pi}}_{n}-E_{n,\theta_{n}}\mathbf{\bar{\Pi}}_{n}\right)\right\|^{2}\leq\lambda_{\max}\left(\hat{\Delta}_{m}^{-1}-\Delta_{m,\theta_{n}}^{-1}\right)^{2}\left\|r^{1/2}\left(\mathbf{\bar{\Pi}}_{n}-E_{n,\theta_{n}}\mathbf{\bar{\Pi}}_{n}\right)\right\|^{2}.$ (3.83)
Here analogously to (3.79)- (3.80) one obtains, in view of $m^{2}\sim\left(\log n\right)^{2}=o\left(n\right),$
$m^{2}\;\lambda_{\max}\left(\hat{\Delta}_{m}^{-1}-\Delta_{m,\theta_{n}}^{-1}\right)^{2}\rightarrow_{p}0.$ (3.84)
Recall that $\mathbf{\bar{\Pi}}_{n}=r^{-1}\sum_{j=1}^{r}\mathbf{\Pi}_{m,j}$ (cf. (3.55)) where $\mathbf{\Pi}_{m,j}$ are i.i.d. vectors; hence
$\mathrm{Cov}_{n,\mathbf{\theta}_{n}}\left(r^{1/2}\left(\mathbf{\bar{\Pi}}_{n}-E_{n,\theta_{n}}\mathbf{\bar{\Pi}}_{n}\right)\right)=\mathrm{Cov}_{n,\mathbf{\theta}_{n}}\left(\mathbf{\Pi}_{m,1}\right)$
and consequently
$E_{n,\mathbf{\theta}_{n}}\left\|r^{1/2}\left(\mathbf{\bar{\Pi}}_{n}-E_{n,\theta_{n}}\mathbf{\bar{\Pi}}_{n}\right)\right\|^{2}$ $=\mathrm{Tr}\;\mathrm{Cov}_{n,\mathbf{\theta}_{n}}\left(\mathbf{\Pi}_{m,1}\right)$
$=\mathrm{Tr}\;\left(\left(U_{m}^{\ast}A_{m}\left(a_{\theta_{n}}\right)U_{m}\right)^{\left[2\right]}-I_{m}\right),$
in view of (3.64), where $A_{m}\left(a_{\theta_{n}}\right)$ is the $m\times m$ symbol matrix pertaining to spectral density $a_{\theta_{n}}$ and For a $m\times m$ matrix $M$, the real matrix $M^{\left[2\right]}$ is defined in (3.63). Hence
$\mathrm{Tr}\;\mathrm{Cov}_{n,\mathbf{\theta}_{n}}\left(\mathbf{\Pi}_{m,1}\right)=\sum_{\left|j\right|\leq\left(m-1\right)/2}\left(\mathbf{u}_{j}^{\ast}A_{m}\left(a_{\theta_{n}}\right)\mathbf{u}_{j}\right)^{2}-m$
where $\mathbf{u}_{j}$ are the $m$-vectors defined in (2.49), (2.55) for the current value of $m$. Then Lemma 2.10 implies that for a constant $C_{M}$ depending only on $M$ we have $\mathbf{u}_{j}^{\ast}A_{m}\left(a_{\theta_{n}}\right)\mathbf{u}_{j}\leq C_{M}$ and hence
$m^{-2}\;E_{n,\mathbf{\theta}_{n}}\left\|r^{1/2}\left(\mathbf{\bar{\Pi}}_{n}-E_{n,\theta_{n}}\mathbf{\bar{\Pi}}_{n}\right)\right\|^{2}\leq m^{-1}\left(C_{M}^{2}-1\right)=o\left(1\right).$
Hence
$m^{-2}\;\left\|r^{1/2}\left(\mathbf{\bar{\Pi}}_{n}-E_{n,\theta_{n}}\mathbf{\bar{\Pi}}_{n}\right)\right\|^{2}\rightarrow_{p}0$
which in conjunction with (3.84) and (3.83) implies (3.82). For (3.81) it now suffices to prove
$S_{n}:=W_{m}^{\prime}\Delta_{m,\theta_{n}}^{-1}r^{1/2}\left(\mathbf{\bar{\Pi}}_{n}-E_{n,\theta_{n}}\mathbf{\bar{\Pi}}_{n}\right)\Longrightarrow_{d}N_{2d+1}\left(0,\Phi_{\theta}\right).\text{ }$ (3.85)
Step 4. We claim that
$\lim_{n\rightarrow\infty}\mathrm{Cov}_{n,\mathbf{\theta}_{n}}\left(S_{n}\right)=\Phi_{\theta}.$ (3.86)
Indeed, following the steps in the proof of Lemma 3.7, we obtain
$\mathrm{Cov}_{n,\mathbf{\theta}_{n}}\left(T_{1,n}\right)$ $=W_{m}^{\prime}\Delta_{m,\theta_{n}}^{-1}\mathrm{Cov}_{n,\mathbf{\theta}_{n}}\left(\mathbf{\Pi}_{1,m}\right)\Delta_{m,\theta_{n}}^{-1}W_{m}$
$=W_{m}^{\prime}\Delta_{m,\theta_{n}}^{-1}\left(U_{m}^{\ast}A_{m}U_{m}\right)^{\left[2\right]}\Delta_{m,\theta_{n}}^{-1}W_{m}-W_{m}^{\prime}\Delta_{m,\theta_{n}}^{-2}W_{m}.$ (3.87)
According to (3.56), the column vectors of the matrix $\Delta_{m,\theta_{n}}^{-1}W_{m}$ are
$\mathbf{\tilde{w}}_{j,m}:=\Delta_{m,\theta_{n}}^{-1}\mathbf{w}_{j,m}=\left(\;\left(a_{\theta_{n}}^{2}\left(\omega_{s,m}\right)-1\right)^{-1}m^{-1/2}\psi_{j}\left(\omega_{s,m}\right)\right)_{\left|s\right|\leq\left(m-1\right)/2}.$
In the proof of Lemma 3.7, relation (3.70) it has been shown that the element $(j,k)$ of the matrix $W_{m}^{\prime}\left(U_{m}^{\ast}A_{m}U_{m}\right)^{\left[2\right]}W_{m}$ satisfies
$\mathbf{w}_{j,m}^{\prime}\left(U_{m}^{\ast}A_{m}U_{m}\right)^{\left[2\right]}\mathbf{w}_{k,m}=\mathbf{w}_{j,m}^{\prime}\Lambda_{m}^{2}\mathbf{w}_{k,m}+o\left(1\right)$
where $\Lambda_{m}=\Lambda_{m,\theta}$ is defined by (3.66), with $\theta=\theta_{n}$ currently. For the vectors $\mathbf{w}_{j,m}$ that proof only used the fact that all components of $\mathbf{w}_{j,m}$ and $\mathbf{w}_{k,m}$ are bounded in modulus by $\sqrt{2}m^{-1/2}$. Replacing $\mathbf{w}_{j,m}$ by $\mathbf{\tilde{w}}_{j,m}$, we note that all components are bounded in modulus by $M\sqrt{2}m^{-1/2}$, due to (3.74). Therefore we have
$\mathbf{\tilde{w}}_{j,m}^{\prime}\left(U_{m}^{\ast}A_{m}U_{m}\right)^{\left[2\right]}\mathbf{\tilde{w}}_{k,m}=\mathbf{\tilde{w}}_{j,m}^{\prime}\Lambda_{m}^{2}\mathbf{\tilde{w}}_{k,m}+o\left(1\right),$
hence from (3.87) the element $(j,k)$ of $\mathrm{Cov}_{n,\mathbf{\theta}_{n}}\left(T_{1,n}\right)$ is
$\mathbf{\tilde{w}}_{j,m}^{\prime}\left(\Lambda_{m,\theta_{n}}^{2}-I_{2d+1}\right)\mathbf{\tilde{w}}_{k,m}+o\left(1\right)$ $=\mathbf{\tilde{w}}_{j,m}^{\prime}\Delta_{m,\theta_{n}}\mathbf{\tilde{w}}_{k,m}+o\left(1\right)$
$=\mathbf{w}_{j,m}^{\prime}\Delta_{m,\theta_{n}}^{-1}\mathbf{w}_{k,m}+o\left(1\right)$
$m^{-1}\sum_{\left|j\right|\leq\left(m-1\right)/2}\left(a_{\theta}^{2}\left(\omega_{j,m}\right)-1\right)^{-1}\psi_{j}\left(\omega_{j,m}\right)\psi_{k}\left(\omega_{j,m}\right)+o\left(1\right).$
This expression converges to
$\frac{1}{2\pi}\int_{\left(-\pi,\pi\right)}\left(a_{\theta}^{2}\left(\omega\right)-1\right)^{-1}\psi_{j}\left(\omega\right)\psi_{k}\left(\omega\right)d\omega,$
in view of Lemma 3.14. The claim (3.86) is proved.
Step 5. We use the Lindeberg-Feller Theorem to show (3.85). Consider independent random $d$-vectors
$X_{n,j}=W_{m}^{\prime}\Delta_{m,\theta_{n}}^{-1}\left(\mathbf{\Pi}_{m,j}-E_{n,\theta_{n}}\mathbf{\Pi}_{m,j}\right)\text{, }j=1,\ldots,r\text{ }$
with $\mathbf{\Pi}_{m,j}$ from (3.55). Then $X_{n,j}$ are identically distributed with $E_{\theta_{n}}X_{n,j}=0,$ and $\sum_{j=1}^{r}r^{-1/2}X_{n,j}=S_{n}$. In view of (3.86), it suffices to establish the Lindeberg condition: for every $\varepsilon>0$
$r^{-1}\sum_{j=1}^{r}E_{n,\mathbf{\theta}_{n}}\left\|X_{n,j}\right\|^{2}\mathbf{1}\left\{r^{-1}\left\|X_{n,j}\right\|^{2}>\varepsilon\right\}\rightarrow 0$ (3.88)
or equivalently
$E_{n,\mathbf{\theta}_{n}}\left\|X_{n,1}\right\|^{2}\mathbf{1}\left\{\left\|X_{n,1}\right\|^{2}>\varepsilon r\right\}\rightarrow 0.$ (3.89)
Define
$Y_{n}:=\mathbf{\Pi}_{m,1}-E_{n,\theta_{n}}\mathbf{\Pi}_{m,1},$
then in view of (3.74) we have $\left\|X_{n,1}\right\|\leq M\left\|Y_{n}\right\|$ and hence for (3.89) it suffices to show
$E_{n,\mathbf{\theta}_{n}}\left\|Y_{n}\right\|^{2}\mathbf{1}\left\{\left\|X_{n,1}\right\|^{2}>\varepsilon r\right\}\rightarrow 0.$
Applying the Cauchy-Schwarz and Markov inequalities, we obtain
$E_{n,\mathbf{\theta}_{n}}\left\|Y_{n}\right\|^{2}\mathbf{1}\left\{\left\|X_{n,1}\right\|^{2}>\varepsilon r\right\}\leq\left(E_{n,\mathbf{\theta}_{n}}\left\|Y_{n}\right\|^{4}\right)^{1/2}\left(\frac{E_{n,\mathbf{\theta}_{n}}\left\|X_{n,1}\right\|^{2}}{\varepsilon r}\right)^{1/2}.$
Here, since $\mathrm{Cov}_{n,\mathbf{\theta}_{n}}\left(S_{n}\right)=\mathrm{Cov}_{n,\mathbf{\theta}_{n}}\left(X_{n,1}\right)$, we have
$E_{n,\mathbf{\theta}_{n}}\left\|X_{n,1}\right\|^{2}=\mathrm{Tr}\;\mathrm{Cov}_{n,\mathbf{\theta}_{n}}\left(S_{n}\right)=O\left(1\right)$
due to (3.86). It now suffices to show
$r^{-1}E_{n,\mathbf{\theta}_{n}}\left\|Y_{n}\right\|^{4}=o\left(1\right).$ (3.90)
Recall that according to Subsection 3.3.2, the random vector $\mathbf{\Pi}_{m,1}$ has the same distribution as $\mathbf{\Pi}_{m}=\left(\Pi_{j}\right)_{\left|j\right|\leq\left(m-1\right)/2}$ given by (3.37) with $n$ replaced by $m$, where according to (3.33).
$\Pi_{j}=2\hat{B}_{j}^{\ast}\hat{B}_{j}+\mathbf{1,}\text{ }\left|j\right|\leq\left(m-1\right)/2.$
Hence
$r^{-1}E_{n,\mathbf{\theta}_{n}}\left\|Y_{n}\right\|^{4}$ $=r^{-1}E_{n,\mathbf{\theta}_{n}}\left(\sum_{\left|j\right|\leq\left(m-1\right)/2}\left(\Pi_{j}-E_{n,\mathbf{\theta}_{n}}\Pi_{j}\right)^{2}\right)^{2}$
$\leq\frac{m}{r}\sum_{\left|j\right|\leq\left(m-1\right)/2}E_{n,\mathbf{\theta}_{n}}\left(\Pi_{j}-E_{n,\mathbf{\theta}_{n}}\Pi_{j}\right)^{4}.$
Further note that for the observables $\tilde{Q}_{j},\tilde{P}_{j}$, $j=1,\ldots,m$ defined in (3.28) for $n=m$, one has
$\hat{B}_{j}^{\ast}\hat{B}_{j}=\frac{1}{2}\left(\tilde{Q}_{j+\left(m+1\right)/2}^{2}+\tilde{P}_{j+\left(m+1\right)/2}^{2}-\mathbf{1}\right),\;\left|j\right|\leq\left(m-1\right)/2$
in analogy to (3.3), by the argument about $\tilde{Q}_{j},\tilde{P}_{j}$ used in the proof of Lemma 3.4. To shorten notation, we now write $s(j):=j-\left(m+1\right)/2$ for $j=1,\ldots,m$. Hence
$\Pi_{s(j)}=\tilde{Q}_{j}^{2}+\tilde{P}_{j}^{2}\text{, }j=1,\ldots,m$
and for $j=1,\ldots,m$
$E_{n,\mathbf{\theta}_{n}}\left(\Pi_{s(j)}-E_{n,\mathbf{\theta}_{n}}\Pi_{s(j)}\right)^{4}=E_{n,\mathbf{\theta}_{n}}\left(\tilde{Q}_{j}^{2}-E_{n,\mathbf{\theta}_{n}}\tilde{Q}_{j}^{2}+\tilde{P}_{j}^{2}-E_{n,\mathbf{\theta}_{n}}\tilde{P}_{j}^{2}\right)^{4}$
$\leq 8\;\left(E_{n,\mathbf{\theta}_{n}}\left(\tilde{Q}_{j}^{2}-E_{n,\mathbf{\theta}_{n}}\tilde{Q}_{j}^{2}\right)^{4}+E_{n,\mathbf{\theta}_{n}}\left(\tilde{P}_{j}^{2}-E_{n,\mathbf{\theta}_{n}}\tilde{P}_{j}^{2}\right)^{4}\right).$
By (3.29), $\tilde{Q}_{j}$ has a normal distribution $\tilde{Q}_{j}\sim N\left(0,u_{s\left(j\right)}^{\ast}A_{m}u_{s\left(j\right)}\right)$ where $A_{m}=A_{m}\left(a_{\theta_{n}}\right)$. Writing $\tilde{Q}_{j}=\left(u_{s\left(j\right)}^{\ast}Au_{s\left(j\right)}\right)^{1/2}Z$ for a standard normal $Z$, we obtain
$E_{\mathbf{\theta}}\left(\tilde{Q}_{j}^{2}-E_{\mathbf{\theta}}\tilde{Q}_{j}^{2}\right)^{4}=\left(u_{s\left(j\right)}^{\ast}A_{m}u_{s\left(j\right)}\right)^{4}\mu_{4}$
where $\mu_{4}$ is the fourth central moment of $N\left(0,1\right)$. Applying the same reasoning to $\tilde{P}_{j}\sim N\left(0,u_{s\left(j\right)}^{\ast}Au_{s\left(j\right)}\right)$, we obtain
$\frac{m}{r}\sum_{\left|j\right|\leq\left(m-1\right)/2}E_{n,\mathbf{\theta}_{n}}\left(\Pi_{j}-E_{n,\mathbf{\theta}_{n}}\Pi_{j}\right)^{4}\leq 8\mu_{4}\frac{m^{2}}{r}\max_{\left|k\right|\leq\left(m-1\right)/2}\left(u_{k}^{\ast}A_{m}u_{k}\right)^{4}.$
To bound $u_{k}^{\ast}A_{m}u_{k}$, apply an Lemma 3.5 to conclude that $\left(u_{k}^{\ast}A_{m}\left(a_{\theta}\right)u_{k}\right)^{2}\leq\left(2d+1\right)M$, for $\left|k\right|\leq\left(m-1\right)/2$ and spectral densities $a_{\theta}\in\Theta_{2}\left(d,M\right)$, Since $m^{2}/r\rightarrow 0$, we obtain (3.90) and hence (3.85).
3.5 A deficiency bound from limit distributions
We now show how uniform asymptotic normality an estimator can be used to establish a bound on the one sided Le Cam deficiency. The result is inspired by the two theorems in [M8̈0].
Theorem 3.15
Consider a sequence of experiments $\mathcal{P}_{n}=\left\{\Omega_{n},\mathcal{X}_{n},P_{n,\theta},\theta\in\Theta\right\}$ where $P_{n,\theta}$ are probability measures on $\left(\Omega_{n},\mathcal{X}_{n}\right)$ and $\Theta$ is a compact subset of $\mathbb{R}^{d}$. Assume that for a sequence of statistics $\hat{\theta}_{n}:\left(\Omega_{n},\mathcal{X}_{n}\right)\rightarrow\left(\mathbb{R}^{d},\mathfrak{B}^{d}\right)$ one has
$\mathcal{L}\left(\sqrt{n}\left(\hat{\theta}_{n}-\theta\right)|P_{n,\theta}\right)\Longrightarrow_{d}N_{d}\left(0,\Sigma_{\theta}\right)\text{ uniformly in }\theta\in\Theta$ (3.91)
where the map $\theta\rightarrow\Sigma_{\theta}$ is continuous in the norm $\left\|\cdot\right\|_{2}$ for covariance matrices and $\Sigma_{\theta}>0,\theta\in\Theta$ . Then for experiments
$\mathcal{Q}_{n}=\left\{\mathbb{R}^{d},\mathfrak{B}^{d},N_{d}\left(\theta,n^{-1}\Sigma_{\theta}\right),\theta\in\Theta\right\}$ (3.92)
one has
$\delta\left(\mathcal{P}_{n},\mathcal{Q}_{n}\right)\rightarrow 0.$ (3.93)
Proof. Let $f$ be a measurable function on $\mathbb{R}^{d}$ with $\left\|f\right\|_{\infty}\leq 1$, set $X_{n,\theta}:=\sqrt{n}\left(\hat{\theta}_{n}-\theta\right)$, and let $Y_{n}$ be a random vector on $\left(\mathbb{R}^{d},\mathfrak{B}^{d}\right)$ with $\mathcal{L}\left(Y_{n}\right)=N_{d}\left(0,\Sigma_{\theta}\right)$. Consider the following Markov kernel: for $x\in\mathbb{R}^{d}$, $A\in\mathfrak{B}^{d}$ and some $\gamma\in\left(0,1\right)$ set
$H_{\gamma}\left(A,x\right)=N_{d}\left(x,\gamma^{2}I_{d}\right)\left(A\right).$
Set $P_{n,\theta}^{\prime}:=\mathcal{L}\left(\sqrt{n}\left(\hat{\theta}_{n}-\theta\right)|P_{n,\theta}\right)$, then the law $H_{\gamma}P_{n,\theta}^{\prime}$ can be described by
$H_{\gamma}P_{n,\theta}^{\prime}$ $=\int H_{\gamma}\left(\cdot,x\right)dP_{n,\theta}^{\prime}\left(x\right)$
$=\mathcal{L}\left(X_{n,\theta}+\gamma Z|P_{n,\theta}\right)$ (3.94)
where $Z$ is a standard normal $d$-vector independent of $X_{n,\theta}$. Analogously we have
$H_{\gamma}N_{d}\left(0,\Sigma_{\theta}\right)=\mathcal{L}\left(Y_{n}+\gamma Z^{\prime}\right).$ (3.95)
where $Z^{\prime}$ is a standard normal $d$-vector independent of $Y_{n}$. Now
$\left\|H_{\gamma}P_{n,\theta}^{\prime}-N_{d}\left(0,\Sigma_{\theta}\right)\right\|_{TV}$ (3.96)
$\leq\left\|H_{\gamma}P_{n,\theta}^{\prime}-H_{\gamma}N_{d}\left(0,\Sigma_{\theta}\right)\right\|_{TV}+\left\|H_{\gamma}N_{d}\left(0,\Sigma_{\theta}\right)-N_{d}\left(0,\Sigma_{\theta}\right)\right\|_{TV}.$ (3.97)
For the first term on the r.h.s. we have (cp. (A.47))
$\left\|H_{\gamma}P_{n,\theta}^{\prime}-H_{\gamma}N_{d}\left(0,\Sigma_{\theta}\right)\right\|_{TV}=\frac{1}{2}\sup_{\left\|f\right\|_{\infty}\leq 1}\left|\int f\;dH_{\gamma}P_{n,\theta}^{\prime}-\int f\;dH_{\gamma}N_{d}\left(0,\Sigma_{\theta}\right)\right|.$ (3.98)
Here
$\int f\;dH_{\gamma}P_{n,\theta}^{\prime}=\int g_{f}\left(x\right)P_{n,\theta}^{\prime}\left(dx\right)$
where
$g_{f}\left(x\right)=\int f\left(t\right)\;dH_{\gamma}\left(dt,x\right)=Ef\left(x+\gamma Z\right)$
and similarly
$\int f\;dH_{\gamma}N_{d}\left(0,\Sigma_{\theta}\right)=\int g_{f}\left(x\right)N_{d}\left(0,\Sigma_{\theta}\right)\left(dx\right).$
We claim that $g_{f}\left(x\right)$ is a Lipschitz function. Indeed for $h\in\mathbb{R}^{d}$
$\left|g_{f}\left(x+h\right)-g_{f}\left(x\right)\right|=\left|Ef\left(x+h+\gamma Z\right)-Ef\left(x+\gamma Z\right)\right|$
[…]
By a well known formula
$H^{2}\left(N\left(x+h,\gamma^{2}I_{d}\right),N\left(h,\gamma^{2}I_{d}\right)\right)=2\left(1-\exp\left(-\frac{1}{8\gamma^{2}}\left\|h\right\|^{2}\right)\right)\leq\frac{\left\|h\right\|^{2}}{4\gamma^{2}}$
so that
$\left|g_{f}\left(x+h\right)-g_{f}\left(x\right)\right|\leq\frac{\left\|h\right\|}{\gamma}.$
It follows that for $\gamma\leq 1$ the function $\gamma g_{f}/2$ satisfies $\left\|f\right\|_{BL}\leq 1$. By (3.98)
$\left\|H_{\gamma}P_{n,\theta}^{\prime}-H_{\gamma}N_{d}\left(0,\Sigma_{\theta}\right)\right\|_{TV}$
$\leq\sup_{\left\|f\right\|_{\infty}\leq 1}\left|\int g_{f}\left(x\right)P_{n,\theta}^{\prime}\left(dx\right)-\int g_{f}\left(x\right)N_{d}\left(0,\Sigma_{\theta}\right)\left(dx\right)\right|$
$\leq 2\gamma^{-1}\;\beta\left(P_{n,\theta}^{\prime}\left(dx\right),N_{d}\left(0,\Sigma_{\theta}\right)\right).$
By Lemma A.5 and (3.91) one obtains for every fixed $\gamma\in\left(0,1\right)$
$\sup_{\theta\in\Theta}\left\|H_{\gamma}P_{n,\theta}^{\prime}-H_{\gamma}N_{d}\left(0,\Sigma_{\theta}\right)\right\|_{TV}\rightarrow 0.$
Hence there is a sequence $\gamma_{n}\rightarrow 0$ such that
$\sup_{\theta\in\Theta}\left\|H_{\gamma_{n}}P_{n,\theta}^{\prime}-H_{\gamma_{n}}N_{d}\left(0,\Sigma_{\theta}\right)\right\|_{TV}\rightarrow 0.$ (3.99)
Now consider the second term in (3.97) for $\gamma=\gamma_{n}$: in view of (3.95) we have
$H_{\gamma}N_{d}\left(0,\Sigma_{\theta}\right)=N_{d}\left(0,\Sigma_{\theta}+\gamma^{2}I_{d}\right)$
and thus
$\left\|H_{\gamma_{n}}N_{d}\left(0,\Sigma_{\theta}\right)-N_{d}\left(0,\Sigma_{\theta}\right)\right\|_{TV}$
$=\left\|N_{d}\left(0,\Sigma_{\theta}+\gamma_{n}^{2}I_{d}\right)-N_{d}\left(0,\Sigma_{\theta}\right)\right\|_{TV}.$
Since the map $\theta\rightarrow\Sigma_{\theta}$ is continuous and $\Theta\subset\mathbb{R}^{d}$ is compact, the set $\left\{\Sigma_{\theta},\theta\in\Theta\right\}$ is compact in Hilbert-Schmidt norm. Then $\Sigma_{\theta}>0,\theta\in\Theta$ implies that (analogously to (A.55))
$s_{1}:=\inf\left\{\lambda_{\min}\left(\Sigma_{\theta}\right):\theta\in\Theta\right\}>0,$
and by compactness we also have
$s_{2}:=\sup\left\{\lambda_{\max}\left(\Sigma_{\theta}\right):\theta\in\Theta\right\}<\infty.$
Then by (A.48), Lemma A.6 and $\gamma_{n}\rightarrow 0$
$\left\|N_{d}\left(0,\Sigma_{\theta}+\gamma_{n}^{2}I_{d}\right)-N_{d}\left(0,\Sigma_{\theta}\right)\right\|_{TV}.$
$\leq C\left\|\gamma_{n}^{2}I_{d}\right\|_{2}=Cd^{1/2}\gamma_{n}^{2}\rightarrow 0$
since $d$ is fixed here. In conjunction with (3.99) and (3.97) this implies
$\sup_{\theta\in\Theta}\left\|H_{\gamma_{n}}P_{n,\theta}^{\prime}-N_{d}\left(0,\Sigma_{\theta}\right)\right\|_{TV}\rightarrow 0.$ (3.100)
Consider now a one-to-one transformation of the sample space $\left(\mathbb{R}^{d},\mathfrak{B}^{d}\right)$ as $T_{\theta}\left(x\right)=n^{-1/2}x+\theta$. For any probability measure $P$ on $\left(\mathbb{R}^{d},\mathfrak{B}^{d}\right)$ consider the induced measure $\left(T_{\theta}\circ P\right)\left(A\right)=P\left(T_{\theta}^{-1}\left(A\right)\right)$, equivalently described by $T_{\theta}\circ P=\mathcal{L}\left(T_{\theta}\left(X\right)\right)$ if $P=\mathcal{L}\left(X\right)$. Note the total variation distance then is invariant: for any $P,Q$
$\left\|P-Q\right\|_{TV}=\left\|T_{\theta}\circ P-T_{\theta}\circ Q\right\|_{TV}.$ (3.101)
Now $T_{\theta}\circ N_{d}\left(0,\Sigma_{\theta}\right)=N_{d}\left(\theta,n^{-1}\Sigma_{\theta}\right)$ and by (3.94)
$T_{\theta}\circ H_{\gamma_{n}}P_{n,\theta}^{\prime}=T_{\theta}\circ\mathcal{L}\left(\sqrt{n}\left(\theta_{n}-\theta\right)+\gamma_{n}Z|P_{n,\theta}\right)$
where $Z$ is a standard normal vector, independent of $\theta_{n}$. Thus
$T_{\theta}\circ H_{\gamma_{n}}P_{n,\theta}^{\prime}$ $=\mathcal{L}\left(T_{\theta}\left(\sqrt{n}\left(\hat{\theta}_{n}-\theta\right)+\gamma_{n}Z\right)|P_{n,\theta}\right)$
$=\mathcal{L}\left(\hat{\theta}_{n}+n^{-1/2}\gamma_{n}Z|P_{n,\theta}\right)$
so that from (3.100) and (3.101) we obtain
$\sup_{\theta\in\Theta}\left\|\mathcal{L}\left(\hat{\theta}_{n}+n^{-1/2}\gamma_{n}Z|P_{n,\theta}\right)-N_{d}\left(\theta,n^{-1}\Sigma_{\theta}\right)\right\|_{TV}\rightarrow 0.$
The transition from $P_{n,\theta}$ to $\mathcal{L}\left(\hat{\theta}_{n}+n^{-1/2}\gamma_{n}Z|P_{n,\theta}\right)$ represents a Markov kernel operation, so that the claim (3.93) follows.
3.6 Le Cam’s globalization method
The ”heteroskedastic normal experiment” (3.92) resulting from Theorem 3.15 arises as a global approximation, roughly speaking, in regular parametric models with asymptotic normalized information matrix $\Sigma_{\theta}^{-1}$; cf. [LC75] and discussions in [Mam86], [Nus96]. We will utilize this result as a tool in our quest for lower information bounds for the quantum time series. Below we cite Le Cam’s original result and then give an application in our context.
For an experiment $\mathcal{P}=\left\{\Omega,\mathcal{X},P_{\theta},\theta\in\Theta\right\}$ and a $S\subset\Theta$ we denote the ”localized” experiment by $\mathcal{P}_{S}:=$ $\left\{\Omega,\mathcal{X}\text{,}P_{\theta},\theta\in S\right\}$. We will frequently omit the sample spaces from notation, with the understanding that they may be different for different experiments. All experiments are assumed to be dominated by sigma-finite measures on their respective sample spaces.
Proposition 3.16
(Theorem 1 in [LC75]) Let $\mathcal{P}=\left\{P_{\theta},\theta\in\Theta\right\}$ and $\mathcal{Q}=\left\{Q_{\theta},\theta\in\Theta\right\}$ be two dominated experiments indexed by the set $\Theta$. Assume that $\Theta$ is metrized by $W$, that $0\leq a<b$ are given. Assume also that
(i) any subset of diameter $4b+2a$ of $\Theta$ can be covered by no more than $C$ sets of diameter $b$,
(ii) if $S\subset\Theta$ has a diameter $3b$ then the deficiency $\delta\left(\mathcal{P}_{S},\mathcal{Q}_{S}\right)$ does not exceed $\varepsilon_{1},$
(iii) there is an estimator $\hat{\theta}_{n}$ available on $\mathcal{P}$ such that $P_{\theta}\left(W\left(\hat{\theta}_{n},\theta\right)>a\right)\leq\varepsilon_{2}$ for all $\theta\in\Theta.$
Then
$\delta\left(\mathcal{P},\mathcal{Q}\right)\leq\varepsilon_{1}+\varepsilon_{2}+\frac{1}{2}\frac{a}{b}C.$
The coverage condition on $\Theta$ is well known to be related to the dimension of $\Theta$. A set $S\subset\Theta$ has diameter $b$ if $b=\sup_{s,t\in S}W\left(s,t\right)$. Since $a<b$, a stronger condition than (i) above is: any subset of diameter $6b$ of $\Theta$ can be covered by no more than $C$ sets of diameter $b$. If $\Theta\subset\mathbb{R}^{d}$, a crude bound for $C$ can be given as follows. [^1] In the paper, this lemma will have to be replaced by a reference.
Lemma 3.17
Assume $\Theta\subset\mathbb{R}^{d}$ and $W\left(\theta_{1},\theta_{2}\right)$ is euclidean distance. Then $C$ can be chosen as $\left(12d\right)^{d}$.
Proof. Assume $S\subset\Theta$ has diameter $6b$. Then it is contained in a ball of radius $6b$. This ball is contained in a square of side length $12b$. The square can be partitioned into $12^{d}$ squares of side length $b$. Each of these squares has radius $\sqrt{d}b$. Each of these squares can be further partitioned into $d^{d}$ smaller squares with side length $b/d$, such that the diameter of these squares is $\sqrt{d}b/d=b/\sqrt{d}\leq b$. Then $S$ can be covered by the totality of these smaller squares, i.e. by $\left(12d\right)^{d}$ sets of diameter $b$.
We will apply Proposition 3.16 when $\mathcal{P}$ is an element of the sequence $\mathcal{P}_{n}=\left\{N_{d}\left(\theta,n^{-1}\Sigma_{\theta}\right),\theta\in\Theta\right\}$ and $\Theta$ is a subset of $\mathbb{R}^{d}$. The claim of Lemma 3.17 remains valid if the euclidean metric $\left\|\theta_{1}-\theta_{2}\right\|$ is replaced by $c\left\|\theta_{1}-\theta_{2}\right\|$ for any $c>0;$ in particular for $W\left(\theta_{1},\theta_{2}\right)=\sqrt{n}\left\|\theta_{1}-\theta_{2}\right\|$. Consider some other sequence of dominated experiments $\mathcal{Q}_{n}=\left\{Q_{n,\theta},\theta\in\Theta\right\}$ and consider localized versions: for $\theta_{0}\in\Theta$ and $r>0$ set
$S_{n}\left(\theta_{0},r\right)=\left\{\theta\in\mathbb{R}^{d}:n^{1/2}\left\|\theta-\theta_{0}\right\|\leq r\right\},$ (3.102)
$\mathcal{P}_{n}\left(\theta_{0},r\right)$ $:=\left\{N_{d}\left(\theta,n^{-1}\Sigma_{\theta}\right),\;,\theta\in\Theta\cap S_{n}\left(\theta_{0},r\right)\right\},$ (3.103)
$\mathcal{Q}_{n}\left(\theta_{0},r\right)$ $:=\left\{Q_{n,\theta},\;\theta\in\Theta\cap S_{n}\left(\theta_{0},r\right)\right\}.$ (3.104)
Lemma 3.18
Assume that the sequence $\mathcal{P}_{n}$ fulfills
$s_{2}:=\sup_{\theta\in\Theta}\lambda_{\max}\left(\Sigma_{\theta}\right)<\infty$ (3.105)
and for every $r>0$
$\sup_{\theta_{0}\in\Theta}\delta\left(\mathcal{P}_{n}\left(\theta_{0},r\right),\mathcal{Q}_{n}\left(\theta_{0},r\right)\right)\rightarrow 0.$ (3.106)
Then
$\delta\left(\mathcal{P}_{n},\mathcal{Q}_{n}\right)\rightarrow 0.$
Proof. First we show that in $\mathcal{P}_{n}$ an estimator $\hat{\theta}_{n}$ is available such that for $P_{n,\theta}=N_{d}\left(\theta,n^{-1}\Sigma_{\theta}\right)$
$\sup_{\theta\in\Theta}P_{n,\theta}\left(n^{1/2}\left\|\hat{\theta}_{n}-\theta\right\|>a\right)\rightarrow 0\text{ as }a\rightarrow\infty$ (3.107)
($\hat{\theta}_{n}$ is uniformly $\sqrt{n}$-consistent). Indeed let $\hat{\theta}_{n}$ be the identity map on $\left(\mathbb{R}^{d},\mathfrak{B}^{d}\right)$, i.e. a random $d$-vector such that $\mathcal{L}\left(\hat{\theta}_{n}|P_{n,\theta}\right)=N_{d}\left(\theta,n^{-1}\Sigma_{\theta}\right)$. Then
$\mathcal{L}\left(n^{1/2}\left(\hat{\theta}_{n}-\theta\right)|P_{n,\theta}\right)=N_{d}\left(0,\Sigma_{\theta}\right),$
hence for a standard normal $d$-vector $Z$
$\sup_{\theta\in\Theta}P_{n,\theta}\left(n^{1/2}\left\|\hat{\theta}_{n}-\theta\right\|>a\right)=\sup_{\theta\in\Theta}P\left(\left\|\Sigma_{\theta}^{1/2}Z\right\|>a\right)$
$\leq\sup_{\theta\in\Theta}P\left(\lambda_{\max}^{1/2}\left(\Sigma_{\theta}\right)\left\|Z\right\|>a\right)\leq P\left(s_{2}^{1/2}\left\|Z\right\|>a\right)\rightarrow 0\text{ as }a\rightarrow\infty$
so (3.107) is shown. Now (3.106) implies that there is a sequence $r_{n}\rightarrow\infty$ such that
$\sup_{\theta_{0}\in\Theta}\delta\left(\mathcal{P}_{n}\left(\theta_{0},r_{n}\right),\mathcal{Q}_{n}\left(\theta_{0},r_{n}\right)\right)\rightarrow 0.$
Let $\varepsilon>0$ and choose $n_{1}$ such that for $n\geq n_{1}$
$\sup_{\theta_{0}\in\Theta}\delta\left(\mathcal{P}_{n}\left(\theta_{0},r_{n}\right),\mathcal{P}_{n}\left(\theta_{0},r_{n}\right)\right)\leq\varepsilon/3.$
Set $b_{n}=2r_{n}/3$; then the diameter of $S_{n}\left(\theta_{0},r_{n}\right)$ is $3b_{n}$. Then choose $n_{2}\geq n_{1}$ such that for $n\geq n_{2}$ and $a_{n}=$ $\sqrt{b_{n}}$
$\sup_{\theta\in\Theta}P_{n,\theta}\left(n^{1/2}\left\|\hat{\theta}_{n}-\theta\right\|>a_{n}\right)\leq\varepsilon/3.$
Finally choose $n_{3}\geq n_{2}$ such that for $n\geq n_{3}$ and the constant $C$ described in Lemma 3.17
$\frac{1}{2}\frac{a_{n}}{b_{n}}C=\frac{C}{2\sqrt{b_{n}}}\leq\varepsilon/3.$
By Proposition 3.16, for $n\geq n_{3}$ we then have $\delta\left(\mathcal{P}_{n},\mathcal{Q}_{n}\right)\leq\varepsilon.$
3.7 Proof of the lower informativity bound
Consider again the set $\Theta_{2}^{\prime}=\Theta_{2}^{\prime}\left(M,d\right)$ given by (3.45) and let $\theta_{0}\in\Theta_{2}^{\prime}$ be a fixed parameter point therein. Recall that the distribution $Q_{n.2}\left(a,a_{0}\right)$ was described by (2.96); with a slight abuse of notation, we write $Q_{n,2}\left(\theta,\theta_{0}\right)$ for this distribution when $a=a_{\theta}$ and $a_{0}=a_{\theta_{0}}$, so that $Q_{n,2}\left(\theta,\theta_{0}\right)$ is described by
$dY_{\omega}=a_{\theta}\left(\omega\right)d\omega+\left(2\pi/n\right)^{1/2}\left(a_{\theta_{0}}^{2}-1\right)^{1/2}dW_{\omega}\text{, }\omega\in\left[-\pi,\pi\right].$ (3.108)
For a subset $S\subset\mathbb{R}^{2d+1}$, define experiments
$\mathcal{G}_{n,2}\left(\theta_{0},S\right)$ $:=\left\{Q_{n,2}\left(\theta,\theta_{0}\right),\;\theta\in S\right\},$ (3.109)
$\mathcal{G}_{n,3}\left(\theta_{0},S\right)$ $:=\left\{N_{2d+1}\left(\theta,n^{-1}\Phi_{\theta_{0}}^{-1}\right),\theta\in S\right\}.$ (3.110)
with $\Phi_{\theta_{0}}$ given by (3.71),
Lemma 3.19
For any $S\subset\mathbb{R}^{2d+1}$, $\theta_{0}\in\Theta_{2}^{\prime}$ and each $n$, we have
$\Delta\left(\mathcal{G}_{n,2}\left(\theta_{0},S\right),\mathcal{G}_{n,3}\left(\theta_{0},S\right)\right)=0.$
Proof. For $\theta=\left(\theta_{j}\right)_{\left|j\right|\leq d}$ we have according to (3.44)
$a_{\theta}\left(\omega\right)=\sum_{\left|j\right|\leq d}\psi_{j}\left(\omega\right)\theta_{j}.$
Define a vector of functions $\Psi:=\left(\psi_{j}\right)_{\left|j\right|\leq d}$ and write $a_{\theta}\left(\omega\right)=\theta^{\prime}\Psi\left(\omega\right)$. For the likelihood ratio in the model (3.108) we have
$\frac{dQ_{n,2}\left(\theta,\theta_{0}\right)}{dQ_{n,2}\left(0,\theta_{0}\right)}\left(Y\right)=\exp\left(\frac{n}{2\pi}\int_{\left[-\pi,\pi\right]}a_{\theta}\left(a_{\theta_{0}}^{2}-1\right)^{-1}dY_{\omega}-\frac{n}{4\pi}\int_{\left[-\pi,\pi\right]}a_{\theta}^{2}\left(a_{\theta_{0}}^{2}-1\right)^{-1}d\omega\right).$
Here we can write
$\int_{\left[-\pi,\pi\right]}a_{\theta}\left(a_{\theta_{0}}^{2}-1\right)^{-1}dY_{\omega}=\theta^{\prime}\int_{\left[-\pi,\pi\right]}\Psi\left(\omega\right)\left(a_{\theta_{0}}^{2}\left(\omega\right)-1\right)^{-1}dY_{\omega}.$
By the Neyman factorization criterion, the random $2d+1$-vector
$T\left(Y\right)=\frac{1}{2\pi}\int_{\left[-\pi,\pi\right]}\Psi\left(a_{\theta_{0}}^{2}-1\right)^{-1}dY_{\omega}$
is a sufficient statistic. Then the distributions of $T\left(Y\right)$ under $Q_{n,2}\left(\theta,\theta_{0}\right)$ for $\theta\in S$ form an equivalent experiment. Clearly these distributions are $2d+1$-variate normal. We have
$E_{n,\theta}T=\frac{1}{2\pi}\int_{\left[-\pi,\pi\right]}\Psi\left(a_{\theta_{0}}^{2}-1\right)^{-1}\Psi^{\prime}\theta d\omega$
In view of (3.71), we have
$E_{n,\theta}T=\Phi_{\theta_{0}}\theta.$
To find the covariance matrix, observe that for $T\left(Y\right)=\left(T_{j}\left(Y\right)\right)_{\left|j\right|\leq d}$ we have
$2\pi\left(T_{j}\left(Y\right)-E_{n,\theta}T_{j}\left(Y\right)\right)$ $=\int_{\left[-\pi,\pi\right]}\psi_{j}\left(a_{\theta_{0}}^{2}-1\right)^{-1}\left(2\pi/n\right)^{1/2}\left(a_{\theta_{0}}^{2}-1\right)^{1/2}dW_{\omega}$
$=\left(2\pi/n\right)^{1/2}\int_{\left[-\pi,\pi\right]}\psi_{j}\left(a_{\theta_{0}}^{2}-1\right)^{-1/2}dW_{\omega}.$
Consequently
$\mathrm{Cov}_{n,\theta}\left(T_{j}\left(Y\right),T_{k}\left(Y\right)\right)$
$=\frac{1}{2\pi n}\int_{\left[-\pi,\pi\right]}\psi_{j}\left(a_{\vartheta_{0}}^{2}-1\right)^{-1}\psi_{k}d\omega$
$=n^{-1}\Phi_{\theta_{0},jk}.$
by (3.72). Hence
$\mathcal{L}\left(T\left(Y\right)|Q_{n,2}\left(\theta,\theta_{0}\right)\right)=N_{2d+1}\left(\Phi_{\theta_{0}}\theta,n^{-1}\Phi_{\theta_{0}}\right)$
and the respective experiment with $\theta\in S$ is equivalent to $\mathcal{G}_{1,n}\left(\theta_{0},S\right)$. Define
$\tilde{T}\left(Y\right):=\Phi_{\theta_{0}}^{-1}T\left(Y\right);$ (3.111)
then
$\mathcal{L}\left(\tilde{T}\left(Y\right)|Q_{n,2}\left(\theta,\theta_{0}\right)\right)=N_{2d+1}\left(\theta,n^{-1}\Phi_{\theta_{0}}^{-1}\right).$ (3.112)
Since (3.111) is a one-to-one transformation of the data, giving an equivalent experiment, and (3.112) with $\theta\in S$ describes $\mathcal{G}_{2,n}\left(\theta_{0},S\right)$, the claim is proved.
Recall that the distribution $Q_{n.1}\left(a\right)$ was described by (1.33); with a slight abuse of notation, we write $Q_{n,1}\left(\theta\right)$ for this distribution when $a=a_{\theta}$ so that $Q_{n,1}\left(\theta\right)$ is described by
$dY_{\omega}=\mathrm{arc\cosh}\left(a_{\theta}\left(\omega\right)\right)d\omega+\left(2\pi/n\right)^{1/2}dW_{\omega},\omega\in\left[-\pi,\pi\right].$
Analogously to (3.109), (3.110) for $S\subset\Theta_{2}^{\prime}\left(d,M\right)$, define an experiment
$\mathcal{G}_{n,1}\left(S\right)=\left\{Q_{n,1}\left(\theta\right),\;\theta\in S\right\}$ (3.113)
and also
$\mathcal{G}_{n,4}\left(S\right):=\left\{N_{2d+1}\left(\theta,n^{-1}\Phi_{\theta}^{-1}\right),\theta\in S\right\}.$ (3.114)
Lemma 3.20
(i) For all $M>0,$ there exists $M^{\prime}>0$ such that $\left\{a_{\theta},\;\theta\in\Theta_{2}^{\prime}\left(d,M\right)\right\}\subset\Theta_{1,c}\left(1,M^{\prime}\right).$
(ii) For all $M>0,$ there exists $M^{\prime}>0$ and a sequence $\gamma_{n}=O\left(n^{-1/2}\right)$ such that for all $\theta_{0}\in\Theta_{2}^{\prime}\left(d,M\right)$
$\left\{a_{\theta},\;\theta\in\Theta_{2}^{\prime}\left(d,M\right)\cap S_{n}\left(\theta_{0},r\right)\right\}\subset\Theta_{1,c}\left(1,M^{\prime}\right)\cap B\left(a_{\theta_{0}},\gamma_{n}\right).$
Proof. (i) Recall the definition of $\Theta_{1,c}\left(1,M^{\prime}\right)$ in (2.85). If $\theta\in\Theta_{2}^{\prime}\left(d,M\right)$ then
$a_{\theta}\left(\omega\right)$ $=\sum_{\left|j\right|\leq d}\theta_{j}\psi_{j}\left(\omega\right),$
$\left|a_{\theta}\left(\omega\right)\right|$ $\leq\left(2d+1\right)^{1/2}\left\|\theta\right\|\leq\left(2d+1\right)^{1/2}M^{1/2},$ (3.115)
$\left|a_{\theta}^{\prime}\left(\omega\right)\right|$ $\leq\left(2d\right)^{1/2}\left\|\theta\right\|\leq\left(2d\right)^{1/2}M^{1/2},$
hence for $\alpha=1$
$\left\|a_{\theta}\right\|_{C^{\alpha}}\leq\left\|a_{\theta}\right\|_{\infty}+\left\|a_{\theta}^{\prime}\right\|_{\infty}\leq 2\left(2d+1\right)^{1/2}M^{1/2}.$
If $\theta\in\Theta_{2}^{\prime}\left(d,M\right)$ then we also have $\inf_{\omega\in\left[-\pi,\pi\right]}a_{\theta}\left(\omega\right)\geq 1+M^{-1}$, so by choosing $M^{\prime}=\max\left(2\left(2d+1\right)^{1/2}M^{1/2},M\right)$ we have $\left\|a_{\theta}\right\|_{C^{\alpha}}\leq M^{\prime}$ and $a_{\theta}\in\mathcal{F}_{M^{\prime}}$, i.e. $a_{\theta}\in\Theta_{1,c}\left(1,M^{\prime}\right)$.
(ii) If $\theta\in S_{n}\left(\theta_{0},r\right)$ then then we have analogously to (3.115)
$\left\|a_{\theta}-a_{\theta_{0}}\right\|_{\infty}\leq\left(2d+1\right)^{1/2}\left\|\theta-\theta_{0}\right\|\leq\left(2d+1\right)^{1/2}n^{-1/2}r=:\gamma_{n}$
and $\gamma_{n}=O\left(n^{-1/2}\right)$. In conjunction with (i) the claim is proved.
Recall that for $\theta_{0}\in\Theta_{2}^{\prime}\left(d,M\right)$ neighborhoods $S_{n}\left(\theta_{0},r\right)$ for $r>0$ are defined by (3.102).
Lemma 3.21
For any $r>0$ and $\Theta_{n}=\Theta_{2}^{\prime}\left(d,M\right)\cap S_{n}\left(\theta_{0},r\right)$ we have
$\sup_{\theta_{0}\in\Theta_{2}^{\prime}}\Delta\left(\mathcal{G}_{n,1}\left(\Theta_{n}\right),\mathcal{G}_{n,4}\left(\Theta_{n}\right)\right)\rightarrow 0.$
Proof. Consider the experiment $\mathcal{G}_{n,2}\left(\theta_{0},\Theta_{n}\right)$ defined by (3.109).
$\mathcal{G}_{n,2}\left(\theta_{0},\Theta_{n}\right):=\left\{Q_{n,2}\left(\theta,\theta_{0}\right),\;\theta\in\Theta_{n}\right\}$
with $Q_{n,2}\left(\theta,\theta_{0}\right)$ given by (3.108). We claim that Lemma 2.22 implies that
$\sup_{\theta_{0}\in\Theta_{2}^{\prime}}\Delta\left(\mathcal{G}_{n,1}\left(\Theta_{n}\right),\mathcal{G}_{n,2}\left(\theta_{0},\Theta_{n}\right)\right)\rightarrow 0.$ (3.116)
Indeed it can be seen that $\mathcal{G}_{n,1}\left(\Theta_{n}\right)$, as a set of probability measures, can be considered a subset of $\mathcal{G}_{n,1}\left(\tilde{\Theta}_{n}\right)$ as defined in Lemma 2.21 for $\tilde{\Theta}_{n}=\Theta_{1,c}\left(\alpha,M^{\prime}\right)\cap B\left(a_{0},\gamma_{n}\right)$ for a certain sequence $\gamma_{n},$ a certain $M^{\prime}>0$ and $\alpha=1$, upon setting $a=a_{\theta}$ and $a_{0}=a_{\theta_{0}}$. (Note that when writing $\mathcal{G}_{n,1}\left(\Theta_{n}\right)$ with $\Theta_{n}\subset\mathbb{R}^{2d+1}$, we understand the measures to be indexed by $\theta\in\mathbb{R}^{2d+1}$, whereas when writing $\mathcal{G}_{n,1}\left(\tilde{\Theta}_{n}\right)$ where $\tilde{\Theta}_{n}$ is a set of functions $a$ on $\left[-\pi,\pi\right]$, we understand the measures to be indexed by functions $a$. But we can compare $\mathcal{G}_{n,1}\left(\Theta_{n}\right)$ and $\mathcal{G}_{n,1}\left(\tilde{\Theta}_{n}\right)$ a sets of probability measures on the same sample space.) With that understanding, the claim
$\mathcal{G}_{n,1}\left(\Theta_{2}^{\prime}\left(d,M\right)\cap S_{n}\left(\theta_{0},r\right)\right)\subset\mathcal{G}_{n,1}\left(\Theta_{1,c}\left(1,M^{\prime}\right)\cap B\left(a_{\theta_{0}},\gamma_{n}\right)\right)\text{, for all }\theta_{0}\in\Theta_{2}^{\prime}\left(d,M\right)$
follows from Lemma 3.20 for a certain $M^{\prime}>0$ and a sequence $\gamma_{n}=O\left(n^{-1/2}\right)$. Analogously we obtain, for the same $M^{\prime}$ and $\gamma_{n}$
$\mathcal{G}_{n,2}\left(\theta_{0},\Theta_{2}^{\prime}\left(d,M\right)\cap S_{n}\left(\theta_{0},r\right)\right)\subset\mathcal{G}_{n,2}\left(a_{\theta_{0}},\Theta_{1,c}\left(1,M^{\prime}\right)\cap B\left(a_{\theta_{0}},\gamma_{n}\right)\right)\text{, for all }\theta_{0}\in\Theta_{2}^{\prime}\left(d,M\right)$
where the experiment on the r.h.s.is defined in (2.97). Since $\gamma_{n}$ fulfills the condition $\gamma_{n}=o\left(\left(n/\log n\right)^{-\alpha/(2\alpha+1)}\right)$ for $\alpha=1$, Lemma 2.22 indeed implies (3.116). Now Lemma 3.19 implies
$\sup_{\theta_{0}\in\Theta_{2}^{\prime}}\Delta\left(\mathcal{G}_{n,2}\left(\theta_{0},\Theta_{n}\right),\mathcal{G}_{n,3}\left(\theta_{0},\Theta_{n}\right)\right)=0.$ (3.117)
We now claim
$\sup_{\theta_{0}\in\Theta_{2}^{\prime}}\Delta\left(\mathcal{G}_{n,3}\left(\theta_{0},\Theta_{n}\right),\mathcal{G}_{n,4}\left(\Theta_{n}\right)\right)\rightarrow 0.$ (3.118)
For that consider the total variation distance, for $\theta,\theta_{0}\in\Theta_{2}^{\prime}\left(d,M\right)$
$\left\|N_{2d+1}\left(\theta,n^{-1}\Phi_{\theta}^{-1}\right)-N_{2d+1}\left(\theta,n^{-1}\Phi_{\theta_{0}}^{-1}\right)\right\|_{TV}$
$=\left\|N_{2d+1}\left(0,\Phi_{\theta}^{-1}\right)-N_{2d+1}\left(0,\Phi_{\theta_{0}}^{-1}\right)\right\|_{TV}$
where the equality is obtained by applying the one-to-one map $x\rightarrow n^{1/2}\left(x-\theta\right)$. By (A.48) the above is upperbounded by
$H\left(N_{2d+1}\left(0,\Phi_{\theta}^{-1}\right),N_{2d+1}\left(0,\Phi_{\theta_{0}}^{-1}\right)\right).$
Now by Lemma A.6 and 3.9 the above is upperbounded by
$C\;\left\|\Phi_{\theta}^{-1}-\Phi_{\theta_{0}}^{-1}\right\|_{2}$
where $\left\|\cdot\right\|_{2}$ denotes Hilbert-Schmidt norm for matrices and $C$ only depends on $M$ and $d$. By Lemma 3.13, the mapping $\theta\longrightarrow\Phi_{\theta}^{-1}$ is continuous in Hilbert-Schmidt norm on the compact set $\Theta_{2}^{\prime}\left(d,M\right)\in\mathbb{R}^{2d+1}$, and thus uniformly continuous ([Die60], 3.16.5). Hence
$\sup_{\theta_{0}\in\Theta_{2}^{\prime}}\Delta\left(\mathcal{G}_{n,3}\left(\theta_{0},\Theta_{n}\right),\mathcal{G}_{n,4}\left(\Theta_{n}\right)\right)$
$\leq\sup_{\theta_{0}\in\Theta_{2}^{\prime}}\sup_{\theta\in\Theta_{2}^{\prime}}\left\|N_{2d+1}\left(\theta,n^{-1}\Phi_{\theta}^{-1}\right)-N_{2d+1}\left(\theta,n^{-1}\Phi_{\theta_{0}}^{-1}\right)\right\|_{TV}$
$\leq\sup_{\theta,\theta_{0}\in\Theta_{2}^{\prime}}C\;\left\|\Phi_{\theta}^{-1}-\Phi_{\theta_{0}}^{-1}\right\|_{2}\rightarrow 0$
confirming (3.118). Now relations (3.116) -(3.118) establish the claim.
Lemma 3.22
We have
$\delta\left(\mathcal{G}_{n,4}\left(\Theta_{2}^{\prime}\right),\mathcal{G}_{n,1}\left(\Theta_{2}^{\prime}\right)\right)\rightarrow 0.$
Proof. Apply 3.18 Lemma with $\mathcal{P}_{n}=\mathcal{G}_{n,4}\left(\Theta_{2}^{\prime}\right)$, $\mathcal{Q}_{n}=\mathcal{G}_{n,1}\left(\Theta_{2}^{\prime}\right)$ and
$\mathcal{P}_{n}\left(\theta_{0},r\right)$ $=\mathcal{G}_{n,4}\left(\Theta_{2}^{\prime}\left(d,M\right)\cap S_{n}\left(\theta_{0},r\right)\right),$
$\mathcal{Q}_{n}\left(\theta_{0},r\right)$ $=\mathcal{G}_{n,1}\left(\Theta_{2}^{\prime}\left(d,M\right)\cap S_{n}\left(\theta_{0},r\right)\right).$
Then condition (3.106) is guaranteed by Lemma 3.21, while condition (3.105) is guaranteed by Lemma 3.9.
Proof of Theorem 1.3.. Identify the experiment $\mathcal{E}_{n}\left(\Theta_{2}\left(d,M\right)\right)$ of (1.28) with a set of states indexed by $\theta\in\Theta_{2}^{\prime}\left(d,M\right)$, i.e. with
$\mathcal{E}_{n,1}\left(\Theta_{2}^{\prime}\right):=\left\{\mathfrak{N}_{n}\left(0,A_{n}\left(a_{\theta}\right)\right),\theta\in\Theta_{2}^{\prime}\right\}.$
In the same way, we can identify the $\mathcal{G}_{n}\left(\Theta_{2}\left(d,M\right)\right)$ in the Theorem with $\mathcal{G}_{n,1}\left(\Theta_{2}^{\prime}\right)$ defined in (3.113). Then the claim is
$\mathcal{G}_{n,1}\left(\Theta_{2}^{\prime}\right)\preceq\mathcal{E}_{n,1}\left(\Theta_{2}^{\prime}\right).$ (3.119)
Consider the observable $\mathbf{\bar{\Pi}}_{n}$ defined in (3.55) and the experiment formed by its distributions under the state $\mathfrak{N}_{n}\left(0,A_{n}\left(a_{\theta}\right)\right)$, i.e.
$\mathcal{E}_{n,k}\left(\Theta_{2}^{\prime}\right)=\left\{\mathcal{L}\left(\mathbf{\bar{\Pi}}_{n}|\theta\right),\;\theta\in\Theta_{2}^{\prime}\right\}.$
Since $\mathbf{\bar{\Pi}}_{n}$ is based on a measurement of the state, the map from $\mathfrak{N}_{n}\left(0,A_{n}\left(a_{\theta}\right)\right)$ to $\mathcal{L}\left(\mathbf{\bar{\Pi}}_{n}|\theta\right)$ is given by a quantum channel, hence
$\mathcal{E}_{n,k}\left(\Theta_{2}^{\prime}\right)\preceq\mathcal{E}_{n,1}\left(\Theta_{2}^{\prime}\right).$ (3.120)
Consider now estimator $\tilde{\theta}_{n}$ according to Definition 3.11, which is a function of $\mathbf{\bar{\Pi}}_{n}$. According to Theorem 3.12, $\tilde{\theta}_{n}$ is asymptotically normal
$n^{1/2}\left(\tilde{\theta}_{n}-\theta\right)\Longrightarrow_{d}N_{2d+1}\left(0,\Phi_{\theta}^{-1}\right)$
uniformly in $\theta\in\Theta_{2}^{\prime}$, so condition (3.91) of Theorem 3.15 is fulfilled. Furthermore $\Theta_{2}^{\prime}$ is compact according to Lemma 3.10, the map $\theta\rightarrow\Phi_{\theta}^{-1}$ is continuous in norm $\left\|\cdot\right\|_{2}$ according to Lemma 3.13, and $\Phi_{\theta}^{-1}>0,\;\theta\in\Theta_{2}^{\prime}$ holds according to Lemm 3.9. Then, with $\mathcal{G}_{n,4}\left(S\right)$ defined by (3.92), Theorem 3.15 gives
$\delta\left(\mathcal{E}_{n,k}\left(\Theta_{2}^{\prime}\right),\mathcal{G}_{n,4}\left(\Theta_{2}^{\prime}\right)\right)\rightarrow 0,$
or in semiordering notation
$\mathcal{G}_{n,4}\left(\Theta_{2}^{\prime}\right)\precsim\mathcal{E}_{n,k}\left(\Theta_{2}^{\prime}\right).$ (3.121)
Now Lemma 3.22 states
$\mathcal{G}_{n,1}\left(\Theta_{2}^{\prime}\right)\precsim\mathcal{G}_{n,4}\left(\Theta_{2}^{\prime}\right).$ (3.122)
Relations (3.120), (3.121) and (3.122) establish the claim (3.119).
Appendix A Appendix
A.1 States, channels, observables
A.1.1 Von Neumann algebras
Let $\mathcal{A}$ be a von Neumann algebra of bounded linear operators on a complex Hilbert space $\mathcal{H}$ ([Con00], §46). $\mathcal{H}$ will be assumed separable in the sequel. The two examples we will consider are (i) the set $\mathcal{L}(\mathcal{H})$ of bounded linear operators on $\mathcal{H}$ ([Con90], IX.7.2) ), (ii) the set of functions $L^{\infty}\left(\mu\right)$ on a $\sigma$-finite measure space $\left(X,\Omega,\mu\right)$, construed as linear operators on $\mathcal{H}=L^{2}\left(\mu\right)$ by pointwise multiplication ([Con90], IX.7.2 for both cases). In the former case, $\mathcal{H}$ will always be a symmetric Fock space $\mathcal{F}\left(\mathbb{C}^{n}\right)$, which his separable ([Par92], 19.3, cf. also Lemma A.2 below). In the latter case, the measurable space $\left(X,\Omega\right)$ will be a Polish space with the respective Borel $\sigma$-algebra, so that $L^{2}\left(\mu\right)$ is separable ([Coh13], 3.4.5).
A.1.2 The predual
For every von Neumann algebra $\mathcal{A}$ there is a Banach space $\mathcal{A}_{\ast}$ such that $\mathcal{A}$ is the dual Banach space of $\mathcal{A}_{\ast}$ ([Sak98], 1.1.2). $\mathcal{A}_{\ast}$ is unique up to an isometric isomorphism ([Sak98] 1.13.3, [SW99] VI.6.9, Corollary 1). $\mathcal{A}_{\ast}$ is called the predual of $\mathcal{A}$; the pertaining duality is
$\left\langle a,\tau\right\rangle=a\left(\tau\right)\text{, }a\in\mathcal{A}\text{, }\tau\in\mathcal{A}_{\ast}.$ (A.1)
The norm on $\mathcal{A}_{\ast}$, written $\left\|\cdot\right\|_{1}$ here, is derived from the norm of the dual Banach space $\mathcal{A}^{\ast}$ ( [BR87], 2.4.18), i. e.
$\left\|\tau\right\|_{1}:=\sup_{\left\|a\right\|\leq 1}\left|\left\langle a,\tau\right\rangle\right|\text{, }\tau\in\mathcal{A}_{\ast}.$ (A.2)
On the other hand, since $\mathcal{A}$ is the dual of $\mathcal{A}_{\ast}$, the norm of $\mathcal{A}$ fulfills
$\left\|a\right\|:=\sup_{\left\|\tau\right\|\leq 1}\left|\left\langle a,\tau\right\rangle\right|\text{.}$
In case (i), if $\mathcal{A=L}(\mathcal{H})$ then $\mathcal{A}_{\ast}\mathcal{=L}^{1}(\mathcal{H}),$ the Banach space of trace class operators $R$ on $\mathcal{H}$ with norm $\left\|R\right\|_{1}=\mathrm{Tr\,}\left(R^{\ast}R\right)^{1/2}$, and (A.1), (A.2) take the form
$\left\langle a,R\right\rangle$ $=\mathrm{Tr\,}aR\text{, }a\in\mathcal{L}(\mathcal{H})\text{, }R\in\mathcal{L}^{1}(\mathcal{H}),$ (A.3)
$\left\|R\right\|_{1}$ $=\mathrm{Tr\,}\left(R^{\ast}R\right)^{1/2},$ (A.4)
([SW99] VI.6, [Cha15] 2.1.6). In case (ii), if $\mathcal{A=}L^{\infty}\left(\mu\right)$ then $\mathcal{A}_{\ast}\mathcal{=}L^{1}\left(\mu\right)$, and (A.1), (A.2) are given by
$\left\langle a,f\right\rangle$ $=\int afd\mu\text{, }a\in L^{\infty}\left(\mu\right)\text{, }f\in L^{1}\left(\mu\right),$ (A.5)
$\left\|f\right\|_{1}$ $=\int\left|f\right|d\mu,$ (A.6)
([Sak98], 1.13.3, [SW99], VI.6.8, [BR87], 2.4.17, [Cha15], 2.1.12)
A.1.3 States
([BR87], [Cha15], sec. 2.2). An element $a$ of $\mathcal{A}$ is positive ($a\geq 0$) if $a$ is self-adjoint and $\left\langle x|ax\right\rangle\geq 0$ for every $x\in\mathcal{H}_{A}$ ([Con90], VIII, §3). A linear functional $\tau:$ $\mathcal{A}\rightarrow\mathbb{C}$ is said to be positive if $\tau(a)\geq 0$ for all $a\geq 0$. Such functionals are continuous (bounded) on $\mathcal{A}$ ([BR87] , 2.3.11). A state on $\mathcal{A}$ is a positive element of $\mathcal{A}_{\ast}$ which takes value $1$ on the unit of $\mathcal{A}$. In case (i), by (A.3) $\tau$ is given by a positive element $\rho_{\tau}$ of $\mathcal{L}^{1}(\mathcal{H})$ with $\mathrm{Tr\,}\rho=1$ (a density operator) such that $\tau\left(A\right)=\mathrm{Tr\,}\rho A$. In case (ii), by by (A.5) $\tau$ is given by a positive function $f_{\tau}$ in $L^{1}\left(\mu\right)$ with $\int f_{\tau}d\mu=1$ (a probability density function) such that $\tau\left(\phi\right)=\int\phi f_{\tau}d\mu$.
A.1.4 Normal maps
For the strong and weak operator topologies on $\mathcal{A}$ (SOT, WOT) cf. [Con00], §8; for the weak* topology cf. [Con00], §20 or its equivalent definition as the $\sigma$-weak topology in [BR87], 2.4.2. For two von Neumann algebras $\mathcal{A},\mathcal{B}$, a linear map $\alpha:\mathcal{A}\rightarrow\mathcal{B}$ is positive if $\alpha\left(a\right)\geq 0$ for every $a\geq 0.$Such maps are bounded [Con00], 33.4. A positive linear map $\alpha:\mathcal{A}\rightarrow\mathcal{B}$ is said to be normal if for every increasing net $\left\{a_{\gamma}\right\}$ such that $a_{\gamma}\rightarrow a$ (SOT) one has $\alpha\left(a_{\gamma}\right)\rightarrow\alpha\left(a\right)$ (SOT) ([Con00], 46.1). If the respective Hilbert spaces $\mathcal{H}_{A}$, $\mathcal{H}_{B}$ are separable then the SOT is metrizable on bounded subsets ([Con90] , IX.1.3) and hence nets can be replaced by sequences. A positive linear map $\alpha$ is normal if and only if it is weak* continuous ([Con00], 46.5). It is clear that compositions of bounded positive normal maps are normal. Consider the special case of $\mathcal{B}=\mathbb{C}$, when $\alpha$ is a positive linear functional on $\mathcal{A}$. The predual of $\mathcal{A}$ can be taken as the Banach space generated by all normal linear forms on $\mathcal{A}$ ([SW99], VI.6.9, or [BR87], 2.4.18, 2.4.21). Thus states on $\mathcal{A}$ can also be described as positive normal linear forms on $\mathcal{A}$ which take value $1$ on the unit of $\mathcal{A}$ (cf. also [Con00] 46.4 or [Cha15], 2.1.7).
A.1.5 Complete positivity
Let $\mathcal{A}$, $\mathcal{B}$ be a von Neumann algebras of operators on respective Hilbert spaces $\mathcal{H}_{A},$ $\mathcal{H}_{B}$. The algebra $M_{n}\left(\mathcal{A}\right)$ of all $n\times n$ matrices with entries from $\mathcal{A}$ acting on the $n$-fold direct sum $\mathcal{H}_{A}^{(n)}:=\mathcal{H}_{A}\oplus\ldots\oplus\mathcal{H}_{A}$ is a von Neumann algebra, with norm derived from its being a subalgebra of $\mathcal{L}\left(\mathcal{H}_{A}^{(n)}\right)$ ([Con00], §34, §44). An element $a=\left(a_{ij}\right)_{i,j=1}^{n}\in$ $M_{n}\left(\mathcal{A}\right)$ is called positive if the associated linear operator on $\mathcal{H}_{A}^{(n)}$ is positive, i.e $a$ is self-adjoint and $\left\langle x|ax\right\rangle\geq 0$ for every $x\in\mathcal{H}_{A}^{(n)}$. For a linear map $\alpha:\mathcal{A}\rightarrow\mathcal{B}$, define an associated map $\alpha_{n}:M_{n}\left(\mathcal{A}\right)\rightarrow M_{n}\left(\mathcal{B}\right)$ by $\alpha_{n}\left(a\right)=\left(\alpha\left(a_{ij}\right)\right)_{i,j=1}^{n}$. The map $\alpha$ is completely positive if for every $n\geq 1$, the map $\alpha_{n}$ is positive ([Cha15], sec. 5.4). Compositions of completely positive maps are completely positive ([Cha15], 5.4.9). If either $\mathcal{A}$ or $\mathcal{B}$ are commutative then every positive linear map is completely positive ([Cha15], 5.4.6).
This concept can be developed in parallel for the preduals $\mathcal{A}_{\ast}$, $\mathcal{B}_{\ast}$. The predual of $M_{n}\left(\mathcal{A}\right)$ is the Banach space $M_{n}\left(\mathcal{A}\right)_{\ast}$ of $n\times n$ matrices with entries from $\mathcal{A}_{\ast}$, acting on $M_{n}\left(\mathcal{A}\right)$ according to
[…]
where $a=\left(a_{ij}\right)_{i,j=1}^{n}$, $\tau=\left(\tau_{ij}\right)_{i,j=1}^{n}$. The norm of $M_{n}\left(\mathcal{A}\right)_{\ast}$ is
$\left\|\tau\right\|_{1}=\sup_{a\in M_{n}\left(\mathcal{A}\right),\left\|a\right\|=1}\left|\left\langle a,\tau\right\rangle\right|\text{, }\tau\in M_{n}\left(\mathcal{A}\right)_{\ast}.$
An element $\tau\in$ $M_{n}\left(\mathcal{A}\right)_{\ast}$ is positive if $\left\langle a,\tau\right\rangle\geq 0$ for every $a\geq 0$, $a\in M_{n}\left(\mathcal{A}\right)$. Let $\mathbf{1}$ be the unit of $\mathcal{A}$ and let $\mathbf{1}_{n}$ be the unit of $M_{n}\left(\mathcal{A}\right)$, i.e. the diagonal matrix with diagonal entries all $\mathbf{1}$. Let $\tau\in$ $M_{n}\left(\mathcal{A}\right)_{\ast}$, $\tau\geq 0$; then
$\left\|\tau\right\|_{1}=\left\langle\mathbf{1}_{n},\tau\right\rangle=\sum_{i=1}^{n}\left\langle\mathbf{1},\tau_{ii}\right\rangle=\sum_{i=1}^{n}\left\|\tau_{ii}\right\|_{1}.$
For a linear map $T:\mathcal{A}_{\ast}\rightarrow\mathcal{B}_{\ast}$, define an associated map $T_{n}:M_{n}\left(\mathcal{A}\right)_{\ast}\rightarrow M_{n}\left(\mathcal{B}\right)_{\ast}$ by $T_{n}\left(a\right)=\left(T\left(a_{ij}\right)\right)_{i,j=1}^{n}$. The map $T$ is completely positive if for every $n\geq 1$, the map $T_{n}$ is positive.
A.1.6 Channels
([OP93], chap. 8). Consider a linear map $\alpha:\mathcal{A}\rightarrow\mathcal{B}$. The mapping $\alpha$ is unital if it maps the unit of $\mathcal{A}$ into the unit of $\mathcal{B}$. A quantum channel is a linear, completely positive, unital and normal map $\alpha:\mathcal{A}\rightarrow\mathcal{B}$. Here boundedness of $\alpha$ follows from positivity ([Con00], 33.4). Compositions of channels are channels again. Channels have the Kraus representation
$\alpha\left(a\right)=\sum_{j=1}^{\infty}V_{j}^{\ast}aV_{j}\text{, }a\in\mathcal{A}$
where $\left\{V_{j}\right\}_{j\geq 1}$ is a sequence of bounded linear operators $V_{j}:\mathcal{H}_{B}\rightarrow$ $\mathcal{H}_{A}$ such that $\sum_{j=1}^{\infty}V_{j}^{\ast}V_{j}=\mathbf{1}$, and the sums are convergent in SOT ([Par92], 29.8, [Cha15], 5.4.16 ). An important special case with $\mathcal{A}=\mathcal{B=L}(\mathcal{H})$ is
$\alpha\left(a\right)=U^{\ast}aU\text{, }a\in\mathcal{A}$ (A.7)
where $U$ is a unitary operator on $\mathcal{H}_{A}$.
A.1.7 State transitions (TP-CP maps)
Since a state is a channel $\tau:\mathcal{A}\rightarrow\mathbb{C}$, it follows that a composition of a state $\tau$ on $\mathcal{A}$ with a channel $\alpha:\mathcal{B}\rightarrow\mathcal{A}$ gives a state $\tau\circ\alpha$ on $\mathcal{B}$. This mapping of states extends to a linear map of the preduals $T:\mathcal{A}_{\ast}\rightarrow\mathcal{B}_{\ast}$; the map $T$ is called the dual channel of $\alpha$. Since $\alpha$ is completely positive, it can be shown that $T$ is completely positive (CP), and since $\alpha$ is unital, it follows that $T$ is norm preserving on positives:
$\left\|T\left(\sigma\right)\right\|_{1}=\left\|\sigma\right\|_{1},\sigma\geq 0\text{, }\sigma\in\mathcal{A}_{\ast}.$ (A.8)
In the case $\mathcal{A=L}(\mathcal{H}_{A})$, $\mathcal{B=L}(\mathcal{H}_{B})$ the latter property can be written $\mathrm{Tr\,}T\left(\rho\right)=\mathrm{Tr\,}\rho$ for $\rho\geq 0$, $\rho\in\mathcal{L}^{1}(\mathcal{H}_{A})$, thus $T$ is trace preserving (TP) on positives. In this context a dual channel $T$ is often called a TP-CP map; more generally a CP linear map $T:\mathcal{A}_{\ast}\rightarrow\mathcal{B}_{\ast}$ fulfilling (A.8) will be called a state transition. State transitions have the contraction property:
$\left\|T\left(\sigma_{1}\right)-T\left(\sigma_{2}\right)\right\|_{1}\leq\left\|\sigma_{1}-\sigma_{2}\right\|_{1},\sigma_{i}\geq 0\text{, }\sigma_{i}\in\mathcal{A}_{\ast}\text{, }i=1,2.$ (A.9)
The pair $\left(\alpha,T\right)$ is said to be a dual pair if
$\left\langle\alpha\left(b\right),\omega\right\rangle=\left\langle b,T\left(\omega\right)\right\rangle\text{, }b\in\mathcal{B}\text{, }\omega\in\mathcal{A}_{\ast}.$ (A.10)
The above construction shows that for every channel $\alpha:\mathcal{B}\rightarrow\mathcal{A}$ there exists a state transition $T:\mathcal{A}_{\ast}\rightarrow\mathcal{B}_{\ast}$ such that $\left(\alpha,T\right)$ is a dual pair. The converse can also be shown: for every state transition $T:\mathcal{A}_{\ast}\rightarrow\mathcal{B}_{\ast}$ there exists a channel $\alpha:\mathcal{B}\rightarrow\mathcal{A}$ such that $\left(\alpha,T\right)$ is a dual pair.
In the case $\mathcal{A=L}(\mathcal{H}_{A})$, $\mathcal{B=L}(\mathcal{H}_{B})$, the duality (A.10) for a given channel $\alpha$ and a state transition (TP-CP map) $T$ writes as
[…]
In the case described in (A.7) where $\mathcal{A}=\mathcal{B=L}(\mathcal{H})$ one has
$T\left(R\right)=URU^{\ast}\text{, }R\in\mathcal{L}^{1}(\mathcal{\mathcal{H}_{\mathcal{A}}}).$
Consider now the case $\mathcal{A=}L^{\infty}\left(\mu\right)$, $\mathcal{B=}L^{\infty}\left(\nu\right)$ where $\mu,\nu$ are a sigma-finite measures on measurable spaces $\left(X,\Omega_{X}\right)$, $\left(Y,\Omega_{Y}\right)$ respectively. Then a dual pair $\left(\alpha,T\right)$ fulfills
$\int\alpha\left(g\right)fd\mu=\int gT\left(f\right)d\nu\text{, }g\in L^{\infty}\left(\nu\right)\text{, }f\in L^{1}\left(\mu\right).$ (A.12)
This duality is described in Theorems 24.4 and 24.5 of [Str85]. Only real function spaces and maps between them are considered, but then the duality (A.12) extends to the complex spaces and corresponding maps. The equivalent terminology for a channel $\alpha:L^{\infty}\left(\nu\right)\rightarrow L^{\infty}\left(\mu\right)$ there is Markov operator (a linear, positive, unital and normal map) and for a state transition $T:L^{\infty}\left(\nu\right)\rightarrow L^{\infty}\left(\mu\right)$ it is stochastic operator (a linear, positive and $\left\|\cdot\right\|_{1}$-norm preserving map on positives).
Assume that $\Omega_{Y}$ is the Borel sigma-algebra of a Polish space $Y$ and $\nu$ is a measure on $\left(Y,\Omega_{Y}\right)$. Then for every state transition $T:L^{1}\left(\mu\right)\rightarrow L^{1}\left(\nu\right)$ there is a Markov kernel $K\left(B,x\right)$, $B\in\Omega_{Y}$, $x\in X$ such that
$\int_{B}T\left(f\right)d\nu=\int K\left(B,\cdot\right)fd\mu\text{, }B\in\Omega_{Y}\text{, }f\in L^{1}\left(\mu\right)\text{, }f\geq 0.$ (A.13)
holds ([Str85], Remark 55.6(3), [Nus96], Proposition 9.2).
A.1.8 *-Homomorphisms
A bounded linear map $\alpha:\mathcal{B}\rightarrow\mathcal{A}$ is called a *-homomorphism if for any $a,b\in\mathcal{B}$
$\alpha\left(ab\right)$ $=\alpha\left(a\right)\alpha\left(b\right)$
$\alpha\left(a^{\ast}\right)$ $=\alpha\left(a\right)^{\ast}$
([Cha15], 1.5.3). Such maps are completely positive ([Cha15], 5.4.2) and $\sigma$-weakly continuous ([BR87], 2.4.23), hence normal. Thus they are quantum channels; in our application, $\mathcal{B}$ will represent a ”smaller” quantum system compared to $\mathcal{A}$, in the sense that $\mathcal{A=B}\otimes\mathcal{C}$ for a von Neumann algebra $\mathcal{C}$ of linear operators on $\mathcal{H}_{C}$. Setting $\alpha:\mathcal{B}\rightarrow\mathcal{A}$ as $\alpha\left(b\right)=b\otimes\mathbf{1}$ where $\mathbf{1}$ is the unit of $\mathcal{C}$, we obtain a *-homomorphism. The corresponding state transition operates by restricting a state $\rho$ on $\mathcal{A}$ to the subalgebra $\mathcal{B}\otimes\mathbf{1}$, isomorphic to $\mathcal{B}$ (the partial trace).
A.1.9 Measurements and observation channels
A channel $\alpha:\mathcal{A}\rightarrow\mathcal{B}$ is said to be an observation channel if $\mathcal{A}$ is commutative ([OP93], chap 8). Here we focus on the case where $\mathcal{A}$ is given by $L^{\infty}\left(\mu\right)$ pertaining to a measurable space $\left(X,\Omega,\mu\right)$ and $\mathcal{B=\mathcal{L}(\mathcal{H}_{\mathcal{B}})}$. Observation channels arise from a positive operator valued measure (POVM) in the following way. A POVM on $\left(X,\Omega\right)$ is a mapping $M:\Omega\rightarrow\mathcal{\mathcal{L}(\mathcal{H}_{\mathcal{B}})}$ with properties (i) $M\left(A\right)\geq 0$, $A\in\Omega$ (hence $M\left(A\right)$ is self-adjoint), (ii) $M\left(X\right)=1$, (iii) if $\left\{A_{j}\right\}_{j=1}^{\infty}$ are pairwise disjoint set from $\Omega$ then
$M\left({\textstyle\bigcup\nolimits_{j=1}^{\infty}}A_{j}\right)={\textstyle\sum\nolimits_{j=1}^{\infty}}M\left(A_{j}\right)$
where the r.h.s. is an SOT convergent sum. Then for any state $\rho\in\mathcal{\mathcal{L}}^{1}\mathcal{(\mathcal{H}_{\mathcal{B}})}$,
$\nu_{\rho}\left(A\right)=\mathrm{Tr\,}\rho M\left(A\right)\text{, }A\in\Omega$ (A.14)
is a probability measure on $\Omega$. This defines a state transition $T$ for a certain measure $\nu_{0}$ on $\left(X,\Omega\right)$ in the following way. Suppose that $\rho_{0}\in\mathcal{\mathcal{L}}^{1}\mathcal{(\mathcal{H}_{\mathcal{B}})}$ is a faithful state on $\mathcal{\mathcal{L}(\mathcal{H}_{\mathcal{B}})}$, i.e. $\rho_{0}>0$, and set $\nu_{0}=\nu_{\rho_{0}}$. Note that such a $\rho_{0}$ exists if and only if $\mathcal{H}$ is separable ([BR87], 2.5.5). Then $\nu_{\rho}\ll\nu_{0}$ and
$T\left(\rho\right)=\frac{d\nu_{\rho}}{d\nu_{0}}$ (A.15)
defines a transition $T:$ $\mathcal{\mathcal{L}}^{1}\mathcal{(\mathcal{H}_{\mathcal{B}})\rightarrow}L^{1}\left(\nu_{0}\right)$. Then the dual $\alpha_{T}:L^{\infty}\left(\nu_{0}\right)\rightarrow\mathcal{\mathcal{L}(\mathcal{H}_{\mathcal{B}})}$ is an observation channel, satisfying for any state $\rho\in\mathcal{\mathcal{L}}^{1}\mathcal{(\mathcal{H}_{\mathcal{B}})}$
$\mathrm{Tr\,}\rho M\left(A\right)=\int_{A}T\left(\rho\right)d\nu_{0}=\mathrm{Tr\,}\rho\alpha_{T}\left(\mathbf{1}_{A}\right)\text{, }A\in\Omega.$ (A.16)
This in conjunction with (A.10) shows that $M\left(A\right)=\alpha_{T}\left(\mathbf{1}_{A}\right)$, where $A\in\Omega$ and $\mathbf{1}_{A}\in L^{\infty}\left(\mu\right)$ is the indicator function.
Conversely, let $\alpha:L^{\infty}\left(\mu\right)\rightarrow\mathcal{\mathcal{L}(\mathcal{H}_{\mathcal{B}})}$ be an observation channel for a sigma-finite $\mu$ on $\left(X,\Omega\right)$ and let $T_{\alpha}:\mathcal{\mathcal{L}}^{1}\mathcal{(\mathcal{H}_{\mathcal{B}})\rightarrow}L^{1}\left(\mu\right)$ be the dual channel (transition). Then there is a POVM $M$ on $\left(X,\Omega\right)$ such that (A.16) holds for $T=T_{\alpha}$ and any state $\rho\in\mathcal{\mathcal{L}}^{1}\mathcal{(\mathcal{H}_{\mathcal{B}})}$, and it follows that $M\left(A\right)=\alpha\left(\mathbf{1}_{A}\right)$, $A\in\Omega$ .
If $M\left(A\right)$, $A\in\Omega$ are projections then $M$ is called a projection valued measure (PVM) or spectral measure.
A.1.10 Real and vector valued observables
Consider a self-adjoint operator $S$ on $\mathcal{\mathcal{H}}$ , possibly unbounded and densely defined. By the spectral theorem there is a PVM $M$ on $\left(\mathbb{R},\mathfrak{B}_{\mathbb{R}}\right)$ ($\mathfrak{B}_{\mathbb{R}}$ being the Borel $\sigma$-algebra) such that
$Sx=\int_{\mathbb{R}}tdM(t)x$
for all $x$ in the domain of $S$, i.e. all $x\in\mathcal{\mathcal{H}}$ satisfying $\int t^{2}d\left\langle x,M(t)x\right\rangle<\infty$, with $\left\langle\cdot,\cdot\right\rangle$ being the inner product of $\mathcal{\mathcal{H}}$ ([Lax02], 32.1). The operator $S$ is bounded if and only if $M$ is concentrated on a bounded set in $\mathbb{R}$. Consider the state transition $T_{M}:\mathcal{L}^{1}\left(\mathcal{H}\right)\rightarrow L^{1}\left(\nu_{0}\right)$ given by the PVM $M$ according to (A.15); its dual $\alpha_{M}:L^{\infty}\left(\nu_{0}\right)\rightarrow\mathcal{\mathcal{L}(\mathcal{H})}$ is an observation channel. For a given state $\rho\in\mathcal{L}^{1}(\mathcal{H})$, application of $T_{M}$ produces a probability density $T_{M}\left(\rho\right)\in$ $L^{1}\left(\nu_{0}\right)$. If $M$ is absolutely continuous w.r.t. Lebesgue measure $\lambda$ (i.e. $S$ has absolutely continuous spectrum, [Lax02], 31.4) then the measure $\nu_{0}$ in (A.15) is absolutely continuous, and setting $p_{0}=d\nu_{0}/d\lambda$ for Lebesgue measure $\lambda$ on $\mathbb{R}$, for a given state $\rho\in\mathcal{L}^{1}(\mathcal{H})$, the transition $\rho\rightarrow T_{M}\left(\rho\right)p_{0}\in L^{1}\left(\lambda\right)$ produces a Lebesgue density on $\left(\mathbb{R},\mathfrak{B}_{\mathbb{R}}\right)$. If $M$ is concentrated on a discrete set $D\subset\mathbb{R}$ (i.e. $S$ has point spectrum), $\kappa$ is counting measure on $D$ and $p_{0}=d\nu_{0}/d\kappa$, then analogously $T_{M}\left(\rho\right)p_{0}$ is a density w.r.t. counting measure on $D$, i.e. gives a discrete distribution.
The random variable having distribution given by the $\nu_{0}$-density $T_{M}(\rho)$ is commonly identified in notation with the operator $S$. If $S$ is bounded then, with $D$ being the support of $M$, applying the basic duality (A.10) with an function $g\left(t\right)=t\mathbf{1}_{D}\left(t\right)$, $t\in\mathbb{R}$, such that $g\in L^{\infty}\left(\nu_{0}\right)$
$E_{\rho}S=\int_{D}tT_{M}(\rho)\left(t\right)\nu_{0}\left(dt\right)=\left\langle g,T_{M}(\rho)\right\rangle=\left\langle\alpha_{M}(g),\rho\right\rangle.$ (A.17)
From (A.16) which holds for all $A\in\Omega$ it can be seen that in the case of a spectral measure $M$, the channel $\alpha_{M}\left(f\right)$ for $f\in L^{\infty}\left(\nu_{0}\right)$ acts as
$\alpha_{M}\left(f\right)=\int f\left(t\right)dM(t)$
so that from (A.17) we obtain
$E_{\rho}S=\mathrm{Tr\,}\left(\int_{D}tdM(t)\right)\rho=\mathrm{Tr\,}S\rho$ (A.18)
giving the basic trace rule for expectation of bounded observables. If the operator $S$ is unbounded but the density $T_{M}(\rho)$ has an expectation then the trace rule $E_{\rho}S=\mathrm{Tr\,}S\rho$ extends from (A.18) through an approximation of $S$ by bounded operators $\int_{B}tdM(t)$ for bounded $B\subset\mathbb{R}$.
Let $S_{i},$ $i=1,2$ be self-adjoint operators on $\mathcal{H}$ with respective spectral measures $M_{i}$, $i=1,2$ on $\left(\mathbb{R},\mathfrak{B}_{\mathbb{R}}\right)$. The operators $S_{i}$ commute ($S_{1}S_{2}=S_{2}S_{1}$) if and only if the respective spectral measures commute, i.e. if $M_{1}(A_{1})M_{2}(A_{2})=M_{2}(A_{2})M_{1}(A_{1})$ for all Borel sets $A_{i}\in\mathfrak{B}_{\mathbb{R}}$ ([Con00], Theorem 10.2). Then all operators $M_{1}(A_{1})M_{2}(A_{2})$ are projections in $\mathcal{H}$, and setting for cylinder sets $A_{1}\times A_{2}\subset\mathbb{R}^{2}$
$M\left(A_{1}\times A_{2}\right):=M_{1}(A_{1})M_{2}(A_{2}),$
by extension to $\mathfrak{B}_{\mathbb{R}^{2}}$ one defines a PVM $M$ on $\left(\mathbb{R}^{2},\mathfrak{B}_{\mathbb{R}^{2}}\right)$ ([Par92], 10.9). For a given state $\rho\in\mathcal{L}^{1}(\mathcal{H})$, the commuting operators $S_{i}$ give a bivariate probability distribution $\nu_{\rho}$ on $\left(\mathbb{R}^{2},\mathfrak{B}_{\mathbb{R}^{2}}\right)$ by
$\nu_{\rho}\left(A_{1}\times A_{2}\right)=\mathrm{Tr\,}\rho M(A_{1}\times A_{2})\text{, }A_{i}\in\mathfrak{B}_{\mathbb{R}},i=1,2$ (A.19)
in accordance with (A.14). Its marginal distributions are those given by the operators $S_{i}$. Therefore, if self-adjoint operators are to be identified in notation with the corresponding random variables, then (A.19) describes a bivariate random variable $\left(S_{1},S_{2}\right)$.
Consider now the Fock space $\mathcal{F}\left(\mathcal{H}\right)$ where $\mathcal{H}$ is a direct sum $\mathcal{H=H}_{1}\oplus\mathcal{H}_{2}$. In this case $\mathcal{F}\left(\mathcal{H}\right)$ is unitarily isomorphic to $\mathcal{F}\left(\mathcal{H}_{1}\right)\otimes\mathcal{F}\left(\mathcal{H}_{2}\right)$ ([Par92], 19.6). Suppose that $\tilde{S}_{i}$ are self-adjoint operators on $\mathcal{F}\left(\mathcal{H}_{i}\right)$ with respective spectral measures $M_{i}$, $i=1,2$, and let $\mathbf{1}_{i}$ be the unit operators on $\mathcal{F}\left(\mathcal{H}_{i}\right)$. Then $S_{1}:=\tilde{S}_{1}\otimes\mathbf{1}_{2}$, $S_{2}:=\mathbf{1}_{1}\otimes\tilde{S}_{2}$ commute on $\mathcal{F}\left(\mathcal{H}\right)$ and thus generate a bivariate random variable, with marginal distributions those generated by $\tilde{S}_{i}$. If $M_{i}$ are the respective spectral measures for $\tilde{S}_{i}$ then the PVM $M$ on $\left(\mathbb{R}^{2},\mathfrak{B}_{\mathbb{R}^{2}}\right)$ generating the joint distribution is
$M(A_{1}\times A_{2})=M_{1}(A_{1})\otimes M_{2}(A_{2})\text{, }A_{i}\in\mathfrak{B}_{\mathbb{R}},i=1,2.$
In this paper, $\mathcal{H=}\mathbb{C}^{n}$ such that $\mathcal{F}\left(\mathcal{H}\right)$ is identified with $\mathcal{F}\left(\mathbb{C}\right)^{\otimes n}$. Let $\tilde{Q},\tilde{P}$ be the pair of canonical observables in $\mathcal{F}\left(\mathbb{C}\right)$ and let $Q_{i},P_{i}$ be their extension to the whole of $\mathcal{F}\left(\mathbb{C}\right)^{\otimes n}$ such that
$Q_{i}=\mathbf{1}^{\otimes\left(i-1\right)}\otimes\tilde{Q}\otimes\mathbf{1}^{\otimes\left(n-i\right)}$ (A.20)
where $\mathbf{1}$ is the unit operator on $\mathcal{F}\left(\mathbb{C}\right)$, and analogously for $P_{i}$. Then any subset of $\left\{Q_{i},P_{i}\text{, }i=1,\ldots,n\right\}$ which does not contain a pair $\left\{Q_{j},P_{j}\right\}$ is a commuting set, and under a state $\mathfrak{N}_{n}\left(0,A\right)$ (cf. 1.19) the corresponding joint distribution is Gaussian. Let
$\tilde{N}=\frac{1}{2}\left(\tilde{Q}^{2}+\tilde{P}^{2}-\mathbf{1}\right)$
be the number operator on $\mathcal{F}\left(\mathbb{C}\right)$ and let $N_{i}$ be its extension to $\mathcal{F}\left(\mathbb{C}\right)^{\otimes n}$ in analogy to (A.20), for $i=1,\ldots,n$. Then $\left\{N_{i}\text{, }i=1,\ldots,n\right\}$ is a commuting set, and under a state $\mathfrak{N}_{n}\left(0,A\right)$ the corresponding joint distribution is discrete (concentrated on $\mathbb{Z}_{+}^{n}$) with Geometric marginals.
A.1.11 Quantum statistical experiments
A quantum statistical experiment is a family of normal states $\mathcal{E}=\left\{\mathcal{A},\tau_{\theta},\theta\in\Theta\right\}$ on a von Neumann algebra $\mathcal{A}$. The experiment $\mathcal{E}$ is said to be dominated if there exists a normal state
$\omega=\sum_{n=1}^{\infty}\lambda_{n}\tau_{n}$ (A.21)
with $\tau_{n}\in\mathcal{E}$, $\lambda_{n}\geq 0$, $\sum_{n=1}^{\infty}\lambda_{n}=1$ such that
$\mathrm{supp\,}\tau_{\theta}\leq\mathrm{supp\,}\omega\text{ for all }\theta\in\Theta$ (A.22)
[…]
We note that the Fock space $\mathcal{F}\left(\mathbb{C}^{n}\right)$ is separable since the exponential vectors $x_{F}$, $x\in\mathbb{C}^{n}$ (cf. (2.1)) are dense in $\mathcal{F}\left(\mathbb{C}^{n}\right)$. For a separable Hilbert space $\mathcal{H}$, a state on the von Neumann algebra $\mathcal{L}\left(\mathcal{H}\right)$ is faithful if the density operator is strictly positive. The Gaussian states $\mathfrak{N}\left(0,A\right)$ on $\mathcal{L}\left(\mathcal{F}\left(\mathbb{C}^{n}\right)\right)$ have density operator (2.5); Lemma A.2 below then shows that if the Hermitian $n\times n$ matrix AIA-I is strictly positive then $\mathfrak{N}\left(0,A\right)$ is faithful. In Theorem 1.1 we consider the quantum experiment
$\mathcal{E}_{n}\left(\Theta\right)=\left\{\mathcal{L}\left(\mathcal{F}\left(\mathbb{C}^{n}\right)\right),\mathfrak{N}\left(0,A_{n}\left(a\right)\right),\,a\in\Theta\right\}$
for $\Theta=\Theta_{1}\left(\alpha,M\right)$ given by (1.31), (1.32). Here (1.32) and Lemma 2.10 guarantee that $A_{n}\left(a\right)-I>0$ for $a\in\Theta$, hence $\mathcal{E}_{n}\left(\Theta\right)$ is homogeneous and in reduced form. The latter also applies to all Gaussian quantum experiments $\mathcal{E}_{n}\left(\Theta\right)$ occurring in this paper with modified $\Theta$. When $\mathcal{E}=\left\{\mathcal{A},\tau_{\theta},\theta\in\Theta\right\}$ is such that $\mathcal{A=L}\left(\mathcal{F}\left(\mathbb{C}^{n}\right)\right)$ and $\mathcal{E}$ is in reduced form, we will omit $\mathcal{A}$ from notation and simply write $\mathcal{E}$ as a family of density operators $\tau_{\theta}\in\mathcal{L}^{1}\left(\mathcal{F}\left(\mathbb{C}^{n}\right)\right)$.
Consider now the commutative case where $\mathcal{A}=L^{\infty}\left(\mu\right)$ on a $\sigma$-finite measure space $\left(X,\Omega,\mu\right)$, construed as an algebra of linear operators acting on $\mathcal{H}=L^{2}\left(\mu\right)$ by pointwise multiplication. Here every $\tau_{\theta},\theta\in\Theta$ can be identified with a probability density $p_{\theta}\in\mathcal{A}_{\ast}=L^{1}\left(\mu\right)$, and for $\phi\in\mathcal{A}$ we have (cp. (A.5))
$\tau_{\theta}\left(\phi\right)=\int\phi p_{\theta}d\mu.$
The set of probability measures $\mathcal{P}=\left\{P_{\theta}:dP_{\theta}/d\mu=p_{\theta},\theta\in\Theta\right\}$ is then dominated by the measure $\mu$ ($P_{\theta}\ll\mu$, $\theta\in\Theta$). By the Halmos-Savage Theorem ([Str85], 20.3) there exists a probability measure
$Q=\sum_{n=1}^{\infty}\lambda_{n}P_{n}$ (A.23)
with $P_{n}\in\mathcal{P}$, $\lambda_{n}\geq 0$, $\sum_{n=1}^{\infty}\lambda_{n}=1$ such that
$P_{\theta}\ll Q\text{, }\theta\in\Theta.$ (A.24)
Then (A.23) and (A.24) imply that for every $B\in\Omega$
$Q\left(B\right)=0\Longleftrightarrow P_{\theta}\left(B\right)=0\text{ for all }\theta\in\Theta.$
The latter relation is also written $\mathcal{P}\sim Q$. Set $q=dQ/d\mu$; then (A.23) can be written as
$q=\sum_{n=1}^{\infty}\lambda_{n}p_{n}$ (A.25)
where $p_{n}=dP_{n}/d\mu$. For a function $f\in L^{1}\left(\mu\right)$, let $\mathrm{supp\,}f$ be the support projection in $L^{\infty}\left(\mu\right)$: if $f_{0}$ is a function in the $\mu$-equivalence class $f$ then $\mathrm{supp\,}f$ is the $\mu$-equivalence class of $\mathbf{1}\left\{f_{0}\left(x\right)\neq 0\right\}$ ([Con00], 54.5). Then (A.24) is equivalent to
$\mathrm{supp\,}p_{\theta}\leq\mathrm{supp\,}q\text{ for all }\theta\in\Theta$ (A.26)
so that (A.25), (A.26) are the versions of (A.21), (A.22) for the quantum experiment $\mathcal{E}=\left\{L^{\infty}\left(\mu\right),p_{\theta},\theta\in\Theta\right\}$.
Consider now an arbitrary family of probability measures $\mathcal{P}=\left\{P_{\theta},\theta\in\Theta\right\}$ on $\left(X,\Omega\right)$ dominated by sigma-finite measure $\mu$, i.e. a dominated classical statistical experiment. The above reasoning shows that there exists a probability measure $Q$ of form (A.23) with $\mathcal{P}\sim Q$. Then $L^{\infty}\left(Q\right)$, $L^{1}\left(Q\right)$ are an M-space and an L-space of $\mathcal{P}$, respectively ([Str85], 24.6, 24.8). The choice of $Q$ is not unique, but all $L^{\infty}\left(Q\right)$ are isometrically isomorphic Banach spaces, and the same holds for $L^{1}\left(Q\right)$. Moreover all the $L^{\infty}\left(Q\right)$ with $\mathcal{P}\sim Q$ are isomorphic as von Neumann algebras. Thus $\mathcal{P}$ can be identified in a canonical way with a quantum experiment $\mathcal{E}_{\mathcal{P},Q}=\left\{L^{\infty}\left(Q\right),dP_{\theta}/dQ,\theta\in\Theta\right\}$. Here $\mathcal{E}_{\mathcal{P},\mathcal{Q}}$ is in reduced form since $1=dQ/dQ\in L^{1}\left(Q\right)$ is a faithful state on $L^{\infty}\left(Q\right)$. The condition that a quantum experiment $\mathcal{E}=\left\{\mathcal{A},\tau_{\theta},\theta\in\Theta\right\}$ be in reduced form thus generalizes the condition that if a classical dominated family $\mathcal{P}$ is represented as $\left\{L^{\infty}\left(\mu\right),dP_{\theta}/d\mu,\,\theta\in\Theta\right\}$, the space $L^{\infty}\left(\mu\right)$ is an $M$-space of $\mathcal{P}$.
We note that for different $Q$, all quantum experiments $\mathcal{E}_{\mathcal{P},Q}$ are statistically equivalent in the sense of the quantum Le Cam distance (1.25). All classical experiments occurring in this paper are dominated, and the simplifying notation $\mathcal{P}=\left\{P_{\theta},\theta\in\Theta\right\}$ will be used to denote one of the (statistically equivalent) quantum experiments $\mathcal{E}_{\mathcal{P},Q}$.
A.2 Further facts about Gaussian states
A.2.1 Partial trace
In [GNZ10], for the treatment of a classical stationary Gaussian time series $X_{1},\ldots,X_{n}$, an essential step of reasoning has been to consider a series where some observations are omitted, say $X_{m+1},\ldots,X_{n}$, $m<n$ and make the obvious claim that the reduced series is ”less informative” than the original. In the framework of Le Cam theory, this means that there exists a transition (Markov kernel) mapping the law $\mathcal{L}\left(X_{1},\ldots,X_{n}\right)$ into its marginal law $\mathcal{L}\left(X_{1},\ldots,X_{m}\right)$. For a Gaussian, zero mean time series, we then know that $X_{1},\ldots,X_{m}$ is again Gaussian centered, and the covariance matrix is just the pertaining submatrix. We now set out to describe the analog of this reasoning for a quantum Gaussian time series.
We will consider centered gauge invariant Gaussian states $\mathfrak{N}_{n}\left(0,A\right)$ with $n\times n$ symbol matrix $A$, given by characteristic function (1.12). Assume for some $m<n$ we consider $\mathfrak{N}_{m}\left(0,A_{(m)}\right)$ where $A_{(m)}$ is the upper $m\times m$ central submatrix of $A$. Is there a quantum channel, mapping $\mathfrak{N}_{n}\left(0,A\right)$ into $\mathfrak{N}_{m}\left(0,A_{(m)}\right)$ for all (permissible) symbol matrices $A$ ?
Let again $\mathcal{H}_{A}$ be a finite dimensional complex Hilbert space which is a direct sum $\mathcal{H}_{A}=\mathcal{H}_{B}\oplus\mathcal{H}_{C}$. The Fock space $\mathcal{F}\left(\mathcal{H}_{A}\right)$ is unitarily isomorphic to $\mathcal{F}\left(\mathcal{H}_{B}\right)\otimes\mathcal{F}\left(\mathcal{H}_{C}\right)$ ([Par92], 19.6) and the respective Weyl operators $W\left(\cdot\right)$ satisfy
$W\left(u_{1}\oplus u_{2}\right)=W\left(u_{1}\right)\otimes W\left(u_{2}\right)\text{ for }u_{1}\in\mathcal{H}_{B},u_{2}\in\mathcal{H}_{C}.$ (A.27)
([Par92], 20.21).
This means that
$CCR_{W}\left(\mathcal{H}_{A}\right)\simeq CCR_{W}\left(\mathcal{H}_{B}\right)\otimes CCR_{W}\left(\mathcal{H}_{C}\right)$ (A.28)
in the sense of a W*-isomorphism ([SW99], VI.6.9). In view of (A.27), (A.28), we can describe the quantum channel realizing the restriction of a Gaussian state on a system $A$ to a subsystem $B$: it is $\alpha:CCR\left(\mathcal{H}_{B}\right)\rightarrow CCR\left(\mathcal{H}_{A}\right)$ given by
$\alpha\left(W\left(u\right)\right)=W\left(u\right)\otimes\mathbf{1=}W\left(u\oplus 0\right)\text{, }u\in\mathcal{H}_{B}.$ (A.29)
It remains to show that for $\mathcal{H}_{A}=\mathbb{C}^{n}$, $\mathcal{H}_{B}=\mathbb{C}^{m}$ we have
$\mathfrak{N}_{n}\left(0,A\right)\circ\alpha=\mathfrak{N}_{m}\left(0,A_{(m)}\right)$ (A.30)
for all Hermitian $A\geq I$. To this end we compute the characteristic function (1.12).
Let $W\left(u\right)\in CCR\left(\mathcal{H}_{B}\right)$ be a Weyl unitary with $u\in\mathbb{C}^{m}$; then for $\rho=\mathfrak{N}_{n}\left(0,A\right)$ according to (1.10) we have
$\hat{W}\left[\mathfrak{\rho}\circ\alpha\right]\left(u\right)$ $:=\left(\mathfrak{\rho}\circ\alpha\right)\left(W\left(u\right)\right)=\rho\left(\alpha\left(W\left(u\right)\right)\right)$
$=\rho\left(W\left(u\oplus 0\right)\right)=\exp\left(-\frac{1}{4}\left\langle\left(u\oplus 0\right),A\left(u\oplus 0\right)\right\rangle\right)$
$=\exp\left(-\frac{1}{4}\left\langle u,A_{(m)}u\right\rangle\right)$
which confirms (A.30).
A.2.2 The density operator under gauge invariance
Lemma A.1
Consider the gauge invariant centered $n$-mode Gausian state $\mathfrak{N}_{n}\left(0,A\right)$ with symbol $A$, where $A$ is a complex Hermitian $n\times n$ matrix fulfilling $A\geq I$. Its density operator on the symmetric Fock space $\mathcal{F}\left(\mathbb{C}^{n}\right)$ is
$\rho_{A}=\frac{2^{n}}{\det\left(I+A\right)}\left(\frac{A-I}{A+I}\right)_{F}.$
Proof. Write $\mathcal{H}=\mathbb{C}^{n}$ and $H=\mathbb{R}^{2n}$. For any $u\in\mathcal{H}$, consider the exponential vector $u_{F}=\oplus_{k=0}^{\infty}\left(k!\right)^{-1/2}u^{\otimes n}$. Then we have
$\left\langle u_{F},v_{F}\right\rangle=\exp\left\langle u,v\right\rangle.$
Define the coherent vector $\psi\left(u\right):=u_{F}\exp\left(-\left\|u\right\|^{2}/2\right)$. For any $x=x_{1}\oplus x_{2}$, $x_{i}\in\mathbb{R}^{n}$ set $\mathrm{c}\left(x\right)=x_{1}+ix_{2}$. We claim that the coherent vectors $\pi^{-n/2}\psi\left(c\left(x\right)\right)$, $x\in H$ form a resolution of the identity, i.e.
$\frac{1}{\pi^{n}}\int_{H}\left|\left(\psi\left(\mathrm{c}\left(x\right)\right)\right)\right\rangle\left\langle\psi\left(\mathrm{c}\left(x\right)\right)\right|dx=I$ (A.31)
where $I$ is the unit operator on $\mathcal{F}\left(\mathbb{C}^{n}\right)$ and the integral converges in a weak sense in $\mathcal{F}\left(\mathbb{C}^{n}\right)$. For a proof, denote $\mu$ the l.h.s. above and note that for every unit vector $\psi\left(y\right)$, $y\in\mathbb{R}^{2n}$
$\left\langle\psi\left(\mathrm{c}\left(y\right)\right)\right|\mu\left|\left(\psi\left(\mathrm{c}\left(y\right)\right)\right)\right\rangle$ $=\frac{1}{\pi^{n}}\int_{H}\exp\left(2\operatorname{Re}\left\langle\mathrm{c}\left(x\right),\mathrm{c}\left(y\right)\right\rangle\right)dx\exp\left(-\left\|y\right\|^{2}\right)$
$=\frac{1}{\pi^{n}}\int_{H}\exp\left(2\left(x,y\right)-\left\|x\right\|^{2}\right)dx\exp\left(-\left\|y\right\|^{2}\right)$
$=\frac{2^{n}}{\left(2\pi\right)^{n}}\int_{H}\exp\left(-\left\|x-y\right\|^{2}\right)dx$
$=\frac{1}{\left(2\pi\right)^{n}\sigma^{2n}}\int_{H}\exp\left(-\frac{1}{2\sigma^{2}}\left\|x-y\right\|^{2}\right)dx$
for $\sigma^{2}=1/2$. The above expression is the integral of the density of the $N_{2n}\left(y,\sigma^{2}I_{2n}\right)$ law, which is $1$. Since the unit vectors $\psi\left(\mathrm{c}\left(y\right)\right)$ are dense in $\mathcal{F}\left(\mathbb{C}^{n}\right)$, (A.31) is proved.
Since $c:\mathbb{R}^{2n}\rightarrow\mathbb{C}^{n}$ is an isometry, for every unitary $U$ there is an orthogonal matrix $O_{U}$ such that $U\mathrm{c}\left(\mu\right)=\mathrm{c}\left(O_{U}\mu\right)$ .Let $W\left(v\right)$, $v\in\mathbb{C}^{n}$ be an element of the Weyl algebra on $\mathcal{F}\left(\mathbb{C}^{n}\right)$, acting on exponential vectors as
$W\left(v\right)u_{F}=\left(v+u\right)_{F}\exp\left(-\left\langle v,u\right\rangle-\left\|v\right\|^{2}/2\right).$
The state $\rho_{A}$ is centered Gaussian gauge invariant if its characteristic function is
$\phi\left(t\right)=\mathrm{tr\;}W\left(\mathrm{c}\left(t\right)\right)\rho_{A}=\exp\left(-\frac{1}{2}\operatorname{Re}\left\langle A\mathrm{c}\left(t\right),\mathrm{c}\left(t\right)\right\rangle\right)\text{, }t\in\mathbb{R}^{2n}.$ (A.32)
Setting $R:=\left(A-I\right)/\left(A+I\right)$, we then have
$A=\left(I+R\right)/\left(I-R\right)\text{, }\frac{I+A}{2}=1/\left(I-R\right)$
and
$\mathrm{tr\;}R_{F}=\frac{1}{\det\left(I-R\right)}=\det\left(\frac{I+A}{2}\right)$
(see [Mos09], Appendix for the last relation). It follows that
$\rho_{A}=\det\left(I-R\right)R_{F}\text{. }$
If $_{F}, uu\in\mathcal{H}$ is an exponential vector then
$R_{F}u_{F}=\left(Ru\right)_{F}.$
To find the characteristic function of $\rho_{A}$, note that
$\phi\left(t\right)$ $=\frac{1}{\pi^{n}}\int_{\mathbb{R}^{2n}}\mathrm{tr\;}W\left(\mathrm{c}\left(t\right)\right)\left|\left(\psi\left(\mathrm{c}\left(x\right)\right)\right)\right\rangle\left\langle\psi\left(\mathrm{c}\left(x\right)\right)\right|\rho_{A}dx$
$=\frac{1}{\pi^{n}}\int_{\mathbb{R}^{2n}}\mathrm{tr\;}W\left(\mathrm{c}\left(t\right)\right)\left|\mathrm{c}\left(x\right)_{F}\right\rangle\left\langle\mathrm{c}\left(x\right)_{F}\right|R_{F}\exp\left(-\left\|x\right\|^{2}\right)dx$
$=\frac{\det\left(I-R\right)}{\pi^{n}}\int_{\mathbb{R}^{2n}}\mathrm{tr\;}\left|\left(\mathrm{c}\left(x\right)+\mathrm{c}\left(t\right)\right)_{F}\right\rangle\left\langle\left(R\mathrm{c}\left(x\right)\right)_{F}\right|\exp\left(-\left\|x\right\|^{2}-\left\langle\mathrm{c}\left(t\right),\mathrm{c}\left(x\right)\right\rangle\right)dx\exp\left(-\left\|t\right\|^{2}/2\right)$
$=\frac{\det\left(I-R\right)}{\pi^{n}}\int_{\mathbb{R}^{2n}}\exp\left(\left\langle R\mathrm{c}\left(x\right),\mathrm{c}\left(x\right)+\mathrm{c}\left(t\right)\right\rangle-\left\|x\right\|^{2}-\left\langle\mathrm{c}\left(t\right),\mathrm{c}\left(x\right)\right\rangle\right)dx\exp\left(-\left\|t\right\|^{2}/2\right)$
Let $R=UDU^{\ast}$ where $D=\mathrm{Diag}(r_{1},\ldots,r_{n})$ is real diagonal and $U$ is unitary in $\mathbb{C}^{n}$. Let $O$ be orthogonal in $\mathbb{R}^{2n}$ such that $R\mathrm{c}\left(x\right)=\mathrm{c}\left(Ox\right)$. By a change of variable $U\mathrm{c}\left(x\right)=\mathrm{c}\left(y\right)$, or equivalently $x=Oy$ , setting $y=\oplus_{j=1}^{n}y_{j}$, $y_{j}\in\mathbb{R}^{2}$ and $t=Os$, $s=\oplus_{j=1}^{n}s_{j}$, $s_{j}\in\mathbb{R}^{2}$ accordingly, we obtain
$\phi\left(t\right)={\displaystyle\prod\limits_{j=1}^{n}}\frac{\left(1-r_{j}\right)}{\pi^{n}}\int_{\mathbb{R}^{2}}\exp\left(\left\langle r_{j}\mathrm{c}\left(y_{j}\right),\mathrm{c}\left(y_{j}\right)+\mathrm{c}\left(s_{j}\right)\right\rangle-\left\|y_{j}\right\|^{2}-\left\langle\mathrm{c}\left(s_{j}\right),\mathrm{c}\left(y_{j}\right)\right\rangle\right)dy_{j}\exp\left(-\left\|s_{j}\right\|^{2}/2\right)$
We will compute each of the factors above, $\phi_{j}\left(t\right)$ say, omitting the index $j$ for the variables. Then each of the factors can be understood as pertaining to the case $n=1$, where $R=r=r_{j}$ and $A=a=\left(1+r_{j}\right)/\left(1-r_{j}\right)$. Then $r=\left(a-1\right)/\left(a+1\right)$, and
$\phi_{j}\left(t\right)=\frac{\left(1-r\right)}{\pi^{n}}\int_{\mathbb{R}^{2}}\exp\left(-\left(1-r\right)\left\|y\right\|^{2}-\left(1-r\right)\left\langle\mathrm{c}\left(y\right),\mathrm{c}\left(s\right)\right\rangle+2i\operatorname{Im}\left\langle\mathrm{c}\left(y\right),\mathrm{c}\left(s\right)\right\rangle-\left\|s\right\|^{2}/2\right)dy.$
Note that for $y=y_{1}\oplus y_{2}$, $y_{i}\in\mathbb{R}$ we have
$\left\langle\mathrm{c}\left(y\right),\mathrm{c}\left(s\right)\right\rangle=\left(y,s\right)+i\left(y,Js\right)$
where $J$ is the operator in $\mathbb{R}$ satisfying
$J\left(y_{1}\oplus y_{2}\right)=y_{2}\oplus-y_{1}.$
Note that $\left(y,Jy\right)=0$. Now
$-\left(1-r\right)\left\langle\mathrm{c}\left(y\right),\mathrm{c}\left(s\right)\right\rangle+2i\left\langle\operatorname{Im}\mathrm{c}\left(y\right),\mathrm{c}\left(y\right)\right\rangle$
$=-\left(1-r\right)\operatorname{Re}\left\langle\mathrm{c}\left(y\right),\mathrm{c}\left(s\right)\right\rangle+i\left(1+r\right)\left\langle\operatorname{Im}\mathrm{c}\left(y\right),\mathrm{c}\left(s\right)\right\rangle$
$=-\left(1-r\right)\left(y,s\right)+i\left(1+r\right)\left(y,Js\right).$
This gives
$\phi_{j}\left(t\right)$ $=\frac{\left(1-r\right)}{\pi}\int_{H}\exp\left(-\left(1-r\right)\left\|y\right\|^{2}-\left(1-r\right)\left(y,s\right)+i\left(1+r\right)\left(y,Js\right)-\left\|s\right\|^{2}/2\right)dy$
$=\frac{\left(1-r\right)}{\pi^{n}}\int_{H}\exp\left(-\left(1-r\right)\left\|y+s/2\right\|^{2}+i\left(1+r\right)\left(y,Js\right)\right)dy\cdot\exp\left(-\left(1+r\right)\left\|s\right\|^{2}/4\right)$
$=\frac{2\left(1-r\right)}{2\pi}\int_{H}\exp\left(-\frac{2\left(1-r\right)}{2}\left\|y+s/2\right\|^{2}+i\left(1+r\right)\left(y,Js\right)\right)dy\cdot\exp\left(-\left(1+r\right)\left\|t\right\|^{2}/4\right)$
The expression before the second exponential factor is the characteristic function of the $N_{2}\left(-t/2,1/2\left(1-r\right)\right)$ law at position $w=\left(1+r\right)Js\in\mathbb{R}^{2}$, which is
$\exp\left(i\left(-s/2,w\right)-\frac{1}{2\cdot 2\left(1-r\right)}\left\|w\right\|^{2}\right)$
$=\exp\left(-\left(1+r\right)i\left(s,Js\right)-\frac{\left(1+r\right)^{2}}{2\cdot 2\left(1-r\right)}\left\|s\right\|^{2}\right)$
$=\exp\left(-a\left(1+r\right)\left\|s\right\|^{2}/4\right).$
Hence
$\phi_{j}\left(t\right)=\exp\left(-\left(a+1\right)\left(1+r\right)\left\|s\right\|^{2}/4\right).$
Since $\left(1+r\right)=2a/\left(a+1\right)$, we obtain
$\phi_{j}\left(t\right)=\exp\left(-a\left\|s\right\|^{2}/2\right)$
Hence, setting $a_{j}=\left(1+r_{j}\right)/\left(1-r_{j}\right)$, we obtain
$\phi\left(t\right)={\displaystyle\prod\limits_{j=1}^{n}}\phi_{j}\left(t\right)=\exp\left(-\sum_{j=1}^{n}a_{j}\left\|s_{j}\right\|^{2}/2\right).$ (A.33)
Here $\left\|s_{j}\right\|^{2}=\left|e_{j}^{\prime}\mathrm{c}\left(O^{\prime}t\right)\right|^{2}$ where $e_{j}$ is the $j$-th standard unit vector in $\mathbb{C}^{n}$. But then
$e_{j}^{\prime}\mathrm{c}\left(O^{\prime}t\right)=e_{j}^{\prime}U^{\ast}\mathrm{c}\left(t\right)=\mathbf{u}_{j}^{\ast}\mathrm{c}\left(t\right)$
where is an $\mathbf{u}_{j}^{\ast}$ eigenvector of $A$ pertaining to eigenvalue $\alpha_{j}$. Then
$a_{j}\sum_{j=1}^{n}\left\|s_{j}\right\|^{2}$ $=a_{j}\sum_{j=1}^{n}\left|\mathbf{u}_{j}^{\ast}\mathrm{c}\left(t\right)\right|^{2}$
$=\left\langle A\mathrm{c}\left(t\right),\mathrm{c}\left(t\right)\right\rangle=\operatorname{Re}\left\langle A\mathrm{c}\left(t\right),\mathrm{c}\left(t\right)\right\rangle$
such that (A.33) yields the claimed form of $\phi\left(t\right)$.
A.2.3 Some facts on Fock operators
The following technical result for finite dimensional $B$ allows to relate the spectral decompositions of $B_{F}$ and $B$ (cp. (A1), (A3) of [Mos09]). Define the multiindex set $D\left(m\right):=\left\{\mathbf{m}\in\mathbb{Z}_{+}^{d}:m_{1}+\ldots+m_{d}=m\right\}$ and for any $\mathbf{m}\in D\left(m\right)$ let $\Pi\left(\mathbf{m},d\right)$ be the set of partitions of $m$ objects into $d$ distinct groups, each of size $m_{j}$, $j=1,\ldots,d$. It is well known that
[…]
For each $\nu\in\Pi\left(\mathbf{m},d\right)$ and $j\in\left\{1,\ldots,m\right\}$, let $\nu\left(j\right)\in\left\{1,\ldots,d\right\}$ be the index of the group to which the $j$-th object has been assigned.
Lemma A.2
Let $B$ be Hermitian on $\mathcal{H}=\mathbb{C}^{d}$ with spectral decomposition $B=\sum_{k=1}^{d}\lambda_{k}\left|e_{k}\right\rangle\left\langle e_{k}\right|$. Then the spectral decomposition of $\vee^{m}B$ is
$\vee^{m}B=\sum_{\mathbf{m}\in D\left(m\right)}\lambda_{\mathbf{m}}\left|e_{\mathbf{m}}\right\rangle\left\langle e_{\mathbf{m}}\right|$ (A.34)
where
$\lambda_{\mathbf{m}}$ $:=\lambda_{1}^{m_{1}}\ldots\lambda_{d}^{m_{d}}\text{, }$ (A.35)
$e_{\mathbf{m}}$ $=\frac{1}{\sqrt{d_{\mathbf{m}}}}\sum_{\nu\in\Pi\left(\mathbf{m},d\right)}e_{\nu\left(1\right)}\otimes\ldots\otimes e_{\nu\left(m\right)}.$ (A.36)
Proof. For $\mathcal{H}=\mathbb{C}^{d}$, consider the symmetrization operator in $\mathcal{H}^{\otimes m}$: let $\mathbf{k}\in\left[1,d\right]^{\times m}$ be a multiindex and let
$\tilde{e}_{\mathbf{k}}:=e_{k(1)}\otimes\ldots\otimes e_{k(m)}$
be an orthonormal basis of $\mathcal{H}^{\otimes m}$; then, if $U_{\sigma}^{(m)}$, $\sigma\in S_{m}$ denotes the standard unitary representation of the symmetric group $S_{m}$ on $\mathcal{H}^{\otimes m}$,
$\Pi_{m}\tilde{e}_{\mathbf{k}}:=\frac{1}{m!}\sum_{\sigma\in S_{m}}U_{\sigma}^{(m)}\left(e_{k(1)}\otimes\ldots\otimes e_{k(m)}\right)$
is the symmetrization operator in $\mathcal{H}^{\otimes m}$, where every $U_{\sigma}^{(m)}\left(e_{k(1)}\otimes\ldots\otimes e_{k(m)}\right)$ gives just a permutation of the tensor components. It is a projection, and the space $\vee^{m}\mathcal{H}$ is the eigenspace. Note that for $\mathbf{k}_{1}$, $\mathbf{k}_{2}\in$ $\left[1,d\right]^{\times m}$ we have $\Pi_{m}\tilde{e}_{\mathbf{k}_{1}}=\Pi_{m}\tilde{e}_{\mathbf{k}_{2}}$ if and only if there exists a multiindex $\mathbf{m}\in\mathbb{Z}_{+}^{d},m_{1}+\ldots+m_{d}=m$ such that both $\tilde{e}_{\mathbf{k}_{j}}$ are permutations of $e_{1}^{\otimes m_{1}}\otimes\ldots\otimes e_{d}^{\otimes m_{d}}$, in other words there exist permutations $\sigma_{1},\sigma_{2}\in S_{m}$ such that
$U_{\sigma_{j}}^{(m)}\left(e_{1}^{\otimes m_{1}}\otimes\ldots\otimes e_{d}^{\otimes m_{d}}\right)=\tilde{e}_{\mathbf{k}_{j}}\text{, }j=1,2.$
If $\Pi_{m}\tilde{e}_{\mathbf{k}_{1}}\neq\Pi_{m}\tilde{e}_{\mathbf{k}_{2}}$ then the images are orthogonal, i.e. $\left\langle\Pi_{m}\tilde{e}_{\mathbf{k}_{1}},\Pi_{m}\tilde{e}_{\mathbf{k}_{2}}\right\rangle=0$. This implies that the set
$\left\{f_{\mathbf{m}},\mathbf{m}\in\mathbb{Z}_{+}^{d},m_{1}+\ldots+m_{d}=m\right\}\text{ where }$
$f_{\mathbf{m}}$ $:=\Pi_{m}\left(e_{1}^{\otimes m_{1}}\otimes\ldots\otimes e_{d}^{\otimes m_{d}}\right)=\frac{1}{m!}\sum_{\sigma\in S_{m}}U_{\sigma}^{(m)}\left(e_{1}^{\otimes m_{1}}\otimes\ldots\otimes e_{d}^{\otimes m_{d}}\right)$
[…]
$\hat{e}_{\nu}:=e_{\nu\left(1\right)}\otimes\ldots\otimes e_{\nu\left(m\right)}\text{, }\nu\in\Pi\left(\mathbf{m},d\right);$ (A.37)
they are orthogonal to each other. Hence
$f_{\mathbf{m}}=\frac{1}{d_{\mathbf{m}}}\sum_{\nu\in\Pi\left(\mathbf{m},d\right)}\hat{e}_{\nu},$
$\left\|f_{\mathbf{m}}\right\|^{2}=\left(\frac{1}{d_{\mathbf{m}}}\right)^{2}d_{\mathbf{m}}=\frac{1}{d_{\mathbf{m}}},$
which implies that the vectors
$e_{\mathbf{m}}:=f_{\mathbf{m}}/\left\|f_{\mathbf{m}}\right\|=\frac{1}{\sqrt{d_{\mathbf{m}}}}\sum_{\nu\in\Pi\left(\mathbf{m},d\right)}\hat{e}_{\nu}$ (A.38)
are an orthonormal basis of $\vee^{m}\mathcal{H}$. To see that they are an eigenbasis of $\vee^{m}B$ for eigenvalues $\lambda_{\mathbf{m}}$, note that each $\hat{e}_{\nu}$ is an eigenvector of $B^{\otimes m}$ for eigenvalue $\lambda_{\mathbf{m}}$, hence $e_{\mathbf{m}}$ is also an eigenvector for $\lambda_{\mathbf{m}}$. Since the $e_{\mathbf{m}}$ are an orthonormal basis of $\vee^{m}\mathcal{H}$, they are an eigenbasis of $\vee^{m}B$.
Lemma A.3
Let $A,B$ be Hermitian operators on $\mathcal{H=}\mathbb{C}^{d}$ such that $0<A<I$, and let $\Gamma\left(B\right):=\oplus_{m=0}^{\infty}\Gamma_{m}\left(B\right)$, where $\Gamma_{m}\left(B\right)$ is the restriction of $\sum_{k=1}^{m}I^{\otimes\left(k-1\right)}\otimes B\otimes I^{\otimes\left(m-k\right)}$ onto $\vee^{m}\mathcal{H}$, with $\Gamma_{0}\left(B\right)=0$. Then
$\mathrm{Tr}\;A_{F}\Gamma\left(B\right)=\frac{1}{\det\left(I-A\right)}\mathrm{Tr}\;\frac{A}{I-A}B.$ (A.39)
Proof. We have
$\mathrm{Tr}\;A_{F}\Gamma\left(B\right)=\sum_{m=0}^{\infty}\mathrm{Tr}\;\left(\vee^{m}A\right)\Gamma_{m}\left(B\right),$
$\mathrm{Tr}\;\left(\vee^{m}A\right)\Gamma_{m}\left(B\right)$ $=\sum_{\mathbf{m\in}D\left(m\right)}\left\langle e_{\mathbf{m}}|\left(\vee^{m}A\right)\Gamma_{m}\left(B\right)|e_{\mathbf{m}}\right\rangle$
$=\sum_{\mathbf{m\in}D\left(m\right)}\lambda_{\mathbf{m}}\left\langle e_{\mathbf{m}}|\Gamma_{m}\left(B\right)|e_{\mathbf{m}}\right\rangle.$
Set $\Gamma_{m,j}\left(B\right)=I^{\otimes\left(j-1\right)}\otimes B\otimes I^{\otimes\left(m-j\right)}$ and let $\check{\Gamma}_{m,j}\left(B\right)$ be the restriction to $\vee^{m}\mathcal{H}$ for $\mathcal{H}=\mathbb{C}^{d}$. We have
$\left\langle e_{\mathbf{m}}|\Gamma_{m}\left(B\right)|e_{\mathbf{m}}\right\rangle=\sum_{j=1}^{m}\left\langle e_{\mathbf{m}}|\check{\Gamma}_{m,j}\left(B\right)|e_{\mathbf{m}}\right\rangle=\sum_{j=1}^{m}\left\langle e_{\mathbf{m}}|\Gamma_{m,j}\left(B\right)|e_{\mathbf{m}}\right\rangle$ (A.40)
since $e_{\mathbf{m}}\in\vee^{m}\mathcal{H}$. Furthermore, using (A.38)
$\left\langle e_{\mathbf{m}}|\Gamma_{m,j}\left(B\right)|e_{\mathbf{m}}\right\rangle=\frac{1}{d_{\mathbf{m}}}\sum_{\nu,\mu\in\Pi\left(\mathbf{m},d\right)}\left\langle e_{\nu}|\Gamma_{m,j}\left(B\right)|e_{\mu}\right\rangle.$
We note that any term $\left\langle e_{\nu}|\Gamma_{m,j}\left(B\right)|e_{\mu}\right\rangle$ must be zero unless $\nu=\mu$. Indeed
$\left\langle e_{\nu}|\Gamma_{m,j}\left(B\right)|e_{\mu}\right\rangle=\left({\displaystyle\prod\limits_{k=1}^{j-1}}\left\langle e_{\nu\left(k\right)}|e_{\mu\left(k\right)}\right\rangle\right)\left\langle e_{\nu\left(j\right)}|B|e_{\mu\left(j\right)}\right\rangle\left({\displaystyle\prod\limits_{k=j+1}^{m}}\left\langle e_{\nu\left(k\right)}|e_{\mu\left(k\right)}\right\rangle\right).$ (A.41)
For two partitions $\nu\neq\mu$, there must be at least two indices $k\in\left\{1,\ldots,m\right\}$ such that $\nu\left(k\right)\neq\mu\left(k\right)$. Indeed if there is no such index then $\nu=\mu$, and if there is only one such index then this contradicts the assumption that both $\nu$ and $\mu$ are in $\Pi\left(\mathbf{m},d\right)$ (i.e. the $l$-th group has a given number of elements $m_{l}$, $l=1,\ldots,d$). This implies that on the r.h.s. of (A.41), either the first or the third factor (or both) are zero, unless $\nu=\mu$. Hence
$\left\langle e_{\mathbf{m}}|\Gamma_{m,j}\left(B\right)|e_{\mathbf{m}}\right\rangle$ $=\frac{1}{d_{\mathbf{m}}}\sum_{\nu\in\Pi\left(\mathbf{m},d\right)}\left\langle e_{\nu}|\Gamma_{m,j}\left(B\right)|e_{\nu}\right\rangle$
$=\frac{1}{d_{\mathbf{m}}}\sum_{\nu\in\Pi\left(\mathbf{m},d\right)}\left\langle e_{\nu\left(j\right)}|B|e_{\nu\left(j\right)}\right\rangle$
and with (A.40)
$\left\langle e_{\mathbf{m}}|\Gamma_{m}\left(B\right)|e_{\mathbf{m}}\right\rangle$ $=\frac{1}{d_{\mathbf{m}}}\sum_{\nu\in\Pi\left(\mathbf{m},d\right)}\sum_{j=1}^{m}\left\langle e_{\nu\left(j\right)}|B|e_{\nu\left(j\right)}\right\rangle$
$=\frac{1}{d_{\mathbf{m}}}\sum_{\nu\in\Pi\left(\mathbf{m},d\right)}\sum_{k=1}^{d}m_{k}\left\langle e_{k}|B|e_{k}\right\rangle$
$=\sum_{k=1}^{d}m_{k}\left\langle e_{k}|B|e_{k}\right\rangle.$
Hence
$\mathrm{Tr}\;\left(\vee^{m}A\right)\Gamma_{m}\left(B\right)$ $=\sum_{\mathbf{m\in}D\left(m\right)}\lambda_{\mathbf{m}}\left(\sum_{k=1}^{d}m_{k}\right)$
$=\sum_{k=1}^{d}\left\langle e_{k}|B|e_{k}\right\rangle\sum_{\mathbf{m\in}D\left(m\right)}\left({\displaystyle\prod\limits_{j=1,\ldots,d,j\neq k}}\lambda_{j}^{m_{j}}\right)m_{k}\lambda_{k}^{m_{k}}\text{.}$
By summing over $m\geq 0$, we obtain
$\mathrm{Tr}\;A_{F}\Gamma\left(B\right)=\sum_{k=1}^{d}\left\langle e_{k}|B|e_{k}\right\rangle\left(\sum_{m=0}^{\infty}m\lambda_{k}^{m}\right){\displaystyle\prod\limits_{j=1,\ldots,d,j\neq k}}\left(\sum_{m=0}^{\infty}\lambda_{j}^{m}\right)$
Using the elementary relation, for $0\leq x<1$
$\sum_{m=0}^{\infty}mx^{m}=\frac{x}{\left(1-x\right)^{2}}$ (A.42)
we obtain
$\mathrm{Tr}\;A_{F}\Gamma\left(B\right)$ $=\sum_{k=1}^{d}\left\langle e_{k}|B|e_{k}\right\rangle\left(\frac{\lambda_{k}}{\left(1-\lambda_{k}\right)^{2}}\right){\displaystyle\prod\limits_{j=1,\ldots,d,j\neq k}}\frac{1}{1-\lambda_{j}}$
$=\left({\displaystyle\prod\limits_{j=1,\ldots,d}}\frac{1}{1-\lambda_{j}}\right)\sum_{k=1}^{d}\frac{\lambda_{k}}{1-\lambda_{k}}\left\langle e_{k}|B|e_{k}\right\rangle$
$=\frac{1}{\det\left(I-A\right)}\mathrm{Tr}\;\frac{A}{I-A}B.$
To compute the relative entropy of Gaussian states, we need the logarithm of a Fock operator. This can be found with the help of the spectral decomposition of Lemma A.2.
Lemma A.4
Let $B$ be Hermitian on $\mathcal{H}=\mathbb{C}^{d}$ with spectral decomposition $B=\sum_{k=1}^{d}\lambda_{k}\left|e_{k}\right\rangle\left\langle e_{k}\right|$. Then
$\log\vee^{m}B=\Gamma_{m}\left(\log B\right)$
where $\Gamma_{m}\left(\cdot\right)$ has been defined in Lemma A.3.
Proof. From Lemma A.2 we obtain, if $B=$
$\log\vee^{m}B$ $=\sum_{\mathbf{m}\in D\left(m\right)}\left(\log\lambda_{\mathbf{m}}\right)\left|e_{\mathbf{m}}\right\rangle\left\langle e_{\mathbf{m}}\right|$
$=\sum_{\mathbf{m}\in D\left(m\right)}\left(\sum_{j=1}^{d}m_{j}\log\lambda_{j}\right)\left|e_{\mathbf{m}}\right\rangle\left\langle e_{\mathbf{m}}\right|.$
It now suffices to show that each $e_{\mathbf{m}}$ is an eigenvector of $\Gamma_{m}\left(\log B\right)$ for eigenvalue $\sum_{j=1}^{d}m_{j}\log\lambda_{j}$:
$\Gamma_{m}\left(\log B\right)e_{\mathbf{m}}=\left(\sum_{j=1}^{d}m_{j}\log\lambda_{j}\right)e_{\mathbf{m}}.$
Equivalently we can show that $\sum_{\nu\in\Pi\left(\mathbf{m},d\right)}\hat{e}_{\nu}$ is an eigenvector for the same eigenvalue, where $\hat{e}_{\nu}$, $\nu\in\Pi\left(\mathbf{m},d\right)$ have been defined in (A.37). Write
$\Gamma_{m}\left(\log B\right)=\sum_{k=1}^{m}\check{\Gamma}_{m,j}\left(\log B\right)$
where $\check{\Gamma}_{m,j}\left(\log B\right)$ is the rectriction to $\vee^{m}\mathcal{H}$ of
$\Gamma_{m,j}\left(\log B\right)=I^{\otimes\left(j-1\right)}\otimes\log B\otimes I^{\otimes\left(m-j\right)}\text{, }j=1,\ldots,m.$
Note that $\sum_{\nu\in\Pi\left(\mathbf{m},d\right)}\hat{e}_{\nu}$ is an element of $\vee^{m}\mathcal{H}$ while the $\hat{e}_{\nu}$ generally are not. But it suffices to show that for all $\nu\in\Pi\left(\mathbf{m},d\right)$
$\left(\sum_{j=1}^{m}\Gamma_{m,j}\left(\log B\right)\right)\hat{e}_{\nu}=\left(\sum_{j=1}^{d}m_{j}\lambda_{j}\right)\hat{e}_{\nu}.$ (A.43)
Consider the particular $\nu\in\Pi\left(\mathbf{m},d\right)$ for which
$\hat{e}_{\nu}=e_{1}^{\otimes m_{1}}\otimes\ldots\otimes e_{d}^{\otimes m_{d}}.$ (A.44)
In this case we have
$\Gamma_{m,j}\left(\log B\right)\hat{e}_{\nu}$ $=\lambda_{1}\hat{e}_{\nu}\text{, }j=1,\ldots,m_{1},$
$\Gamma_{m,j}\left(\log B\right)\hat{e}_{\nu}$ $=\lambda_{2}\hat{e}_{\nu}\text{, }j=m_{1}+1,\ldots,m_{1}+m_{2},$
$\ldots$
$\Gamma_{m,j}\left(\log B\right)\hat{e}_{\nu}$ $=\lambda_{d}\hat{e}_{\nu}\text{, }j=\sum_{j=1}^{d-1}m_{j}+1,\ldots,m.$
This implies (A.43) for $\hat{e}_{\nu}$ given by (A.43). Since all other $\hat{e}_{\nu}$, $\nu\in\Pi\left(\mathbf{m},d\right)$ arise from a permutation of the tensor factors, they also fulfill (A.43).
A.3 Uniform convergence in distribution
Let us define uniform convergence in distribution, following [IH81], Appendix I. Consider a sample space $\left(\mathbb{R}^{d},\mathfrak{B}^{d}\right)$; convergence in distribution of a sequence of probability measures $Q_{n}$ to some $Q$ is written $Q_{n}\Longrightarrow_{d}Q$. Assume on $\left(\mathbb{R}^{d},\mathfrak{B}^{d}\right)$ there is a sequence of families of probability measures $\mathcal{P}_{n}=\left\{P_{n,\theta}\text{, }\theta\in\Theta\right\}$, $n\in\mathbb{N}$ where $\Theta$ is an arbitrary set. The family $\mathcal{P}_{n}$ is said to uniformly converge in distribution to a family $\mathcal{P}=\left\{P_{\theta}\text{, }\theta\in\Theta\right\}$ if for every bounded continuous function $g$ on $\mathbb{R}^{d}$ we have
$\int_{\mathbb{R}^{d}}gdP_{n,\theta}\rightarrow\int_{\mathbb{R}^{d}}gdP_{\theta}$ (A.45)
uniformly in $\theta$.
Consider the bounded Lipschitz norm for real valued functions $f$ on $\mathbb{R}^{d}$
$\left\|f\right\|_{BL}:=\left\|f\right\|_{\infty}+\sup_{x\neq y}\frac{\left|f\left(x\right)-f\left(y\right)\right|}{\left|x-y\right|}\text{, }\left\|f\right\|_{\infty}:=\sup_{x}\left|f\left(x\right)\right|$ (A.46)
and the bounded Lipschitz metric for probability measures $P,Q$ on $\mathbb{R}^{d}$
$\beta\left(P,Q\right):=\sup\left\{\left|\int f\left(dP-dQ\right)\right|:\left\|f\right\|_{BL}\leq 1\right\}.$
It is well known that $P_{n}\Longrightarrow_{d}Q$ if and only if $\beta\left(P_{n},Q\right)\rightarrow 0$ ([Dud89], Theorem 11.3.3). Also consider the total variation metric:
$\left\|P-Q\right\|_{TV}=\sup_{A\in\mathfrak{B}_{\mathcal{X}}}\left|P\left(A\right)-Q\left(A\right)\right|.$
Recall that for $\nu=P+Q$ and $p=dP/d\nu$, $q=dQ/d\nu$ one has
$\left\|P-Q\right\|_{TV}$ $=\frac{1}{2}\int\left|p-q\right|d\mu=\frac{1}{2}\left\|P-Q\right\|_{1}\text{ where }$ (A.47a)
$\left\|P-Q\right\|_{1}$ $=\sup\left\{\left|\int f\left(dP-dQ\right)\right|:\left\|f\right\|_{\infty}\leq 1\text{, }f\text{ measurable}\right\}.$ (A.47b)
Also consider the Hellinger metric
$H\left(P,Q\right)=\left(\int\left(p^{1/2}-q^{1/2}\right)^{2}d\nu\right)^{1/2}.$
See [Tsy09], Sec. 2.4 for relations between these distances. In particular, by Le Cam’s inequality ([Tsy09], Lemma 2.3), one has
$\left\|P-Q\right\|_{TV}\leq H\left(P,Q\right).$ (A.48)
Lemma A.5
Assume $\Theta$ is a compact metric space with metric $\mu$ and the mapping $\theta\rightarrow P_{\theta}$, $\theta\in\Theta$ is continuous in total variation metric. Then the following statements are equivalent:
(i) $\mathcal{P}_{n}$ uniformly converges in distribution to $\mathcal{P}$
(ii) $\sup_{\theta}\beta\left(P_{n,\theta},P_{\theta}\right)\rightarrow 0$
(iii) For every sequence $\left\{\theta_{n}\right\}$ such that $\theta_{n}\rightarrow\theta$ for some $\theta\in\Theta$, one has $P_{n,\theta_{n}}\Longrightarrow_{d}P_{\theta}$.
Proof. (i) $\Longrightarrow$(ii). Assume that (ii) does not hold. Then there is a subsequence $\mathcal{N}_{1}\subset\mathbb{N}$ such that $P_{\theta_{n}}$ converges in total variation to some $P_{\theta}$ along $n\in\mathcal{N}_{1}$, but for some $\delta>0$
$\beta\left(P_{n,\theta_{n}},P_{\theta_{n}}\right)>\delta,\text{ }n\in\mathcal{N}_{1}\text{.}$
In view of (i), we have for every bounded continuous $g$
$\left|\int_{\mathbb{R}^{d}}gdP_{n,\theta_{n}}-\int_{\mathbb{R}^{d}}gdP_{\theta_{n}}\right|\rightarrow 0,\text{ }n\in\mathcal{N}_{1}\text{.}$
Recall that the total variation metric satisfies
$\left\|P-Q\right\|_{TV}=\frac{1}{2}\sup\left\{\left|\int f\left(dP-dQ\right)\right|:\left\|f\right\|_{\infty}\leq 1\text{, }f\text{ measurable}\right\},$ (A.49)
hence
$\beta\left(P,Q\right)\leq 2\left\|P-Q\right\|_{TV}.$
and for every bounded continuous $g$
$\left|\int_{\mathbb{R}^{d}}gdP-\int_{\mathbb{R}^{d}}gdQ\right|\leq\frac{2\left\|P-Q\right\|_{TV}}{\left\|g\right\|_{\infty}}.$
Now $\left\|P_{\theta_{n}}-P_{\theta}\right\|_{TV}\rightarrow 0$, $n\in\mathcal{N}_{1}$ implies
$\beta\left(P_{n,\theta_{n}},P_{\theta}\right)$ $\geq\beta\left(P_{n,\theta_{n}},P_{\theta_{n}}\right)-\beta\left(P_{\theta_{n}},P_{\theta}\right)$
$\geq\delta/2,\text{ }n\in\mathcal{N}_{1}\text{, }n\text{ sufficiently large }$ (A.50)
and for every bounded continuous $g$
$\left|\int_{\mathbb{R}^{d}}gdP_{n,\theta_{n}}-\int_{\mathbb{R}^{d}}gdP_{\theta}\right|$
$\leq\left|\int_{\mathbb{R}^{d}}gdP_{n,\theta_{n}}-\int_{\mathbb{R}^{d}}gdP_{\theta_{n}}\right|+\frac{2\left\|P_{\theta_{n}}-P_{\theta}\right\|_{TV}}{\left\|g\right\|_{\infty}}\rightarrow 0,\text{ }n\in\mathcal{N}_{1}.$
The latter relation means $P_{n,\theta_{n}}\rightarrow_{d}P_{\theta}$ along $n\in\mathcal{N}_{1}$, hence $\beta\left(P_{n,\theta_{n}},P_{\theta}\right)\rightarrow 0$, which contradicts (A.50).
[…]
$\beta\left(P_{n,\theta_{n}},P_{\theta}\right)$ $\leq\beta\left(P_{n,\theta_{n}},P_{\theta_{n}}\right)+\beta\left(P_{\theta_{n}},P_{\theta}\right)$
$\leq\beta\left(P_{n,\theta_{n}},P_{\theta_{n}}\right)+2\left\|P_{\theta_{n}}-P_{\theta}\right\|_{TV}\rightarrow 0.$
Hence $\beta\left(P_{n,\theta_{n}},P_{\theta}\right)\rightarrow 0$, implying $P_{n,\theta_{n}}\overset{d}{\implies}P_{\theta}$.
[…]
$\left|\int_{\mathbb{R}^{d}}gdP_{n,\theta_{n}}-\int_{\mathbb{R}^{d}}gdP_{\theta_{n}}\right|\geq\delta\text{, }n\in\mathcal{N}_{1}.$
Then there is a further subsequence $\mathcal{N}_{2}\subset\mathcal{N}_{1}$ such that for some $\theta\in\Theta$ one has $\theta_{n}\rightarrow\theta$ along $\mathcal{N}_{2}$ and hence
$\left\|P_{\theta_{n}}-P_{\theta}\right\|_{TV}\rightarrow 0,n\in\mathcal{N}_{2}.$
This implies
[…]
Define a sequence $\theta_{n}^{\ast}$, $n\in\mathbb{N}$ by $\theta_{n}^{\ast}=\theta_{n}$ for $n\in\mathcal{N}_{2}$, $\theta_{n}^{\ast}=\theta$ for $n\notin\mathcal{N}_{2}$. Then $\theta_{n}^{\ast}\rightarrow\theta$ and by (iii) we have $P_{\theta_{n}^{\ast}}\overset{d}{\implies}P_{\theta}$, which contradicts (A.51).
In the context of the CLT, consider a set $\mathcal{S}$ of family of $d\times d$ nonsingular covariance matrices and the Hilbert-Schmidt norm $\left\|\Sigma\right\|_{2}=\left(\mathrm{Tr}\;\Sigma^{2}\right)^{1/2}$.
Lemma A.6
(Lemma 2.1 of [GNZ10]) Suppose the set $\mathcal{S}$ satisfies
$s_{1}:=\inf_{\Sigma\in\mathcal{S}}\lambda_{\min}\left(\Sigma\right)>0,\;\;s_{2}:=\sup_{\Sigma\in\mathcal{S}}\lambda_{\max}\left(\Sigma\right)<\infty.$
Then there exists $C>0$ depending on $s_{1},s_{2}$ but not on $d$ such that for all $\Sigma_{1},\Sigma_{2}\in\mathcal{S}$
$H\left(N_{d}\left(0,\Sigma_{1}\right),N_{d}\left(0,\Sigma_{2}\right)\right)\leq C\;\left\|\Sigma_{1}-\Sigma_{2}\right\|_{2}.$ (A.52)
Lemma A.7
Consider a set of normal distributions $\mathcal{P}=\left\{N_{d}\left(0,\Sigma\right),\Sigma\in\mathcal{S}\right\}$ where $\mathcal{S}$ is compact in Hilbert-Schmidt metric and satisfies
$\inf_{\Sigma\in\mathcal{S}}\lambda_{\min}\left(\Sigma\right)>0.$ (A.53)
Then the mapping $\Sigma\rightarrow N_{d}\left(0,\Sigma\right)$ is continuous on $\mathcal{S}$ in total variation metric.
Proof. By (A.48) we obtain for $\Sigma_{1},\Sigma_{2}\in\mathcal{S}$
$\left\|N_{d}\left(0,\Sigma_{1}\right)-N_{d}\left(0,\Sigma_{2}\right)\right\|_{TV}\leq H\left(N_{d}\left(0,\Sigma_{1}\right),N_{d}\left(0,\Sigma_{2}\right)\right).$ (A.54)
Note that compactness of $\mathcal{S}$ implies
$s_{2}=\sup_{\Sigma\in\mathcal{S}}\lambda_{\max}\left(\Sigma\right)\leq\sup_{\Sigma\in\mathcal{S}}\left(\mathrm{Tr\;}\Sigma^{2}\right)^{1/2}=\sup_{\Sigma\in\mathcal{S}}\left\|\Sigma\right\|_{2}<\infty$
and it is easy to see that (A.53) implies
$s_{1}=\inf_{\Sigma\in\mathcal{S}}\lambda_{\min}\left(\Sigma\right)>0.$ (A.55)
Then Lemma A.6 implies the claim.
A.4 Geometric distribution
Let $X$ be a r.v. with geometric law $\mathrm{Geo}\left(p\right)$ for parameter $p\in\left(0,1\right)$, given by
$P\left(X=k\right)=\mathrm{Geo}\left(p\right)\left(k\right)=\left(1-p\right)p^{k}\text{, }k=0,1,\ldots$
As is well known, for a sequence of i.i.d. Bernoulli r.v’s with success probability q=1p,q=1-p, the r.v. $X$ is the number of failures before the first success occurs ($X=0$ if success occurs in the first trial). Since $\mathrm{Geo}\left(p\right)$, $p\in\left(0,1\right)$ forms an exponential family, we refer to section 2.1. of [GN98] for some basic properties of that family. Setting $x=k$, the probabilities can be written
$\left(1-p\right)p^{x}$ $=\exp\left(x\log p+\log\left(1-p\right)\right)$
$=\exp\left(x\tau-V\left(\tau\right)\right)=:q\left(x,\tau\right)$ (A.56)
for $\tau=\log p$, and
$V\left(\tau\right)=-\log\left(1-p\right)=-\log\left(1-\exp\tau\right).$
Thus (A.56) is the canonical form of the exponential family, and $\tau\in\left(-\infty,0\right)$ the relevant parameter. The moments are, noting that $V^{\prime\prime}\left(\tau\right)$ is also the Fisher information $I\left(\tau\right)$,
$E_{\tau}X$ $=V^{\prime}\left(\tau\right)=\frac{\exp\tau}{1-\exp\tau}=\frac{p}{1-p},$ (A.57)
$\mathrm{Var}\left(X\right)$ $=V^{\prime\prime}\left(\tau\right)=\frac{\exp\tau\left(1-\exp\tau\right)+\left(\exp\tau\right)^{2}}{\left(1-\exp\tau\right)^{2}}$
$=\frac{\exp\tau}{\left(1-\exp\tau\right)^{2}}=I\left(\tau\right)=\frac{p}{\left(1-p\right)^{2}}.$ (A.58)
In connection with the representation of the thermal state $\mathfrak{N}_{1}\left(0,a\right)$, $a>1$ (cf. (1.20) and (2.5)) we are interested in yet another parametrization of $\mathrm{Geo}\left(p\right)$: setting $\mathbb{\;}p=\left(a-1\right)/\left(a+1\right)$, we obtain
$a=\frac{2}{1-p}-1=\frac{1+p}{1-p}.$
The canonical parameter $\tau$ then can be expressed as
$\tau=\tau\left(a\right)=\log\left(\left(a-1\right)/\left(a+1\right)\right).$ (A.59)
We note
$\tau^{\prime}\left(a\right)$ $=\frac{2}{a^{2}-1}\text{, }\tau^{\prime\prime}\left(a\right)=-\frac{4a}{\left(a^{2}-1\right)^{2}},$ (A.60)
$V^{\prime}\left(\tau\left(a\right)\right)$ $=\frac{a-1}{2}\text{,\ }V^{\prime\prime}\left(\tau\left(a\right)\right)=\frac{a^{2}-1}{4}.$ (A.61)
The fourth central moment of $X$ is, for r=1pr=1-p [Wei]
$E\left(X-EX\right)^{4}=\frac{\left(1-r\right)\left(r^{2}-9r+9\right)}{r^{4}}\leq\frac{10}{\left(1-p\right)^{4}}.$
In terms of parameter $a$ this bound is
$E\left(X-EX\right)^{4}\leq\frac{5}{8}\left(a+1\right)^{4}.$ (A.62)
In accordance with (A.56) and (A.59) the geometric probability function, now parametrized by $a$, is
$q\left(x,\tau\left(a\right)\right)=\exp\left(x\tau\left(a\right)-V\left(\tau\left(a\right)\right)\right)\text{, }x=0,1,\ldots$ (A.63)
Then the score function in this parametrization is
$s\left(x,a\right)$ $:=\frac{\partial}{\partial a}\log q\left(x,\tau\left(a\right)\right)=\left(x-V^{\prime}\left(\tau\left(a\right)\right)\right)\tau^{\prime}\left(a\right)$ (A.64)
$=\left(x-\frac{a-1}{2}\right)\frac{2}{a^{2}-1},$
and Fisher information is
$J\left(a\right)$ $:=E_{a}s^{2}\left(X,a\right)=\frac{4}{\left(a^{2}-1\right)^{2}}\mathrm{Var}_{a}\left(X\right)$ (A.65)
$=\frac{4}{\left(a^{2}-1\right)^{2}}\cdot\frac{a^{2}-1}{4}=\frac{1}{a^{2}-1}.$ (A.66)
Lemma A.8
(i) If $1+C_{1}^{-1}\leq a\leq C_{1}$ for some $C_{1}>0$ then for some $C_{2}$
$E_{a}s^{4}\left(X,a\right)\leq C_{2}.$
(ii) We have
$\frac{\partial^{2}}{\partial a^{2}}q^{1/2}\left(x,\tau\left(a\right)\right)=q^{1/2}\left(x,\tau\left(a\right)\right)\cdot\rho\left(x,a\right)$ (A.67)
where $\rho\left(x,a\right)$ has the property: $1+C_{1}^{-1}\leq a\leq C_{1}$ implies that for some $C_{3}$
$\sum_{x=0}^{\infty}\left(\frac{\partial^{2}}{\partial a^{2}}q^{1/2}\left(x,\tau\left(a\right)\right)\right)^{2}=E_{a}\rho^{2}\left(X,a\right)\leq C_{3}.$ (A.68)
Proof. (i) By (A.64) and (A.62)
$E_{a}s^{4}\left(X,a\right)=E_{a}\left(X-E_{a}X\right)^{4}\frac{2}{a^{2}-1}\leq\frac{5\cdot 2^{4}}{8}\frac{\left(a+1\right)^{4}}{a^{2}-1}=\frac{10}{\left(a-1\right)^{4}}\leq C_{2}.$
(ii) We have
$\frac{\partial}{\partial a}q^{1/2}\left(x,\tau\left(a\right)\right)$ $=\frac{1}{2}q^{1/2}\left(x,\tau\left(a\right)\right)\frac{\partial}{\partial a}\log q\left(x,\tau\left(a\right)\right)$
$=\frac{1}{2}q^{1/2}\left(x,\tau\left(a\right)\right)\cdot\left(x-V^{\prime}\left(\tau\left(a\right)\right)\right)\tau^{\prime}\left(a\right)$
and hence
$\frac{\partial^{2}}{\partial a^{2}}q^{1/2}\left(x,\tau\left(a\right)\right)$ $=\frac{1}{4}q^{1/2}\left(x,\tau\left(a\right)\right)\cdot\left(x-V^{\prime}\left(\tau\left(a\right)\right)\right)^{2}\left(\tau^{\prime}\left(a\right)\right)^{2}$
$-\frac{1}{2}q^{1/2}\left(x,\tau\left(a\right)\right)\cdot V^{\prime\prime}\left(\tau\left(a\right)\right)\left(\tau^{\prime}\left(a\right)\right)^{2}$
$+\frac{1}{2}q^{1/2}\left(x,\tau\left(a\right)\right)\cdot\left(x-V^{\prime}\left(\tau\left(a\right)\right)\right)\tau^{\prime\prime}\left(a\right).$
Hence the l.h.s. of (A.68) is bounded by
$E_{a}\left(X-V^{\prime}\left(\tau\left(a\right)\right)\right)^{4}\left(\tau^{\prime}\left(a\right)\right)^{4}+\left(V^{\prime\prime}\left(\tau\left(a\right)\right)\right)^{2}\left(\tau^{\prime}\left(a\right)\right)^{4}$
$+E_{a}\left(X-V^{\prime}\left(\tau\left(a\right)\right)\right)^{2}\left(\tau^{\prime\prime}\left(a\right)\right)^{2}.$
By (A.60), (A.61), the terms $V^{\prime\prime}\left(\tau\left(a\right)\right)$, $\tau^{\prime}\left(a\right)$ and $\tau^{\prime\prime}\left(a\right)$ are all bounded when $1+C_{1}^{-1}\leq a\leq C_{1}$. It now suffices to prove that
$E_{a}\left(X-V^{\prime}\left(\tau\left(a\right)\right)\right)^{4}+E_{a}\left(x-V^{\prime}\left(\tau\left(a\right)\right)\right)^{2}$
is bounded. The first term above is the fourth central moment of $X$ which is bounded by (A.62). The second term is the variance of $X$, which is $V^{\prime\prime}\left(\tau\left(a\right)\right)$ by (A.58) and (A.61) and thus bounded as well.
Estimation of parameter $a$. From (A.57) and (A.61) we obtain
$E_{\theta}\left(2X+1\right)=a.$ (A.69)
Setting $\hat{a}=2X+1$, we thus obtain an unbiased estimator of $a$ based on one observation $X$. We also have by (A.58) and (A.61)
[…]
$=a^{2}-1$ (A.70)
which is the inverse Fisher information $1/J\left(a\right)$ from (A.65). Hence $\hat{a}$ is best unbiased estimator of $a$. If $\bar{X}_{n}$ is the mean of $n$ i.i.d. observations with law $\mathrm{Geo}\left(p\right)$ then $\hat{a}_{n}=2\bar{X}_{n}+1$ is best unbiased estimator of $a$, with variance $1/nJ\left(a\right)$.
Asymptotically equivalent family. The local approximating Gaussian shift model (according to LAN theory) is
$Y=a+n^{-1/2}\sqrt{\left(a_{0}^{2}-1\right)}\xi$ (A.71)
where $\xi\sim N\left(0,1\right)$ and $a_{0}$ is the center of the parametric neighborhood in $a$. The variance-stable form (cf. Section 3.3 of [GN98]) is
$Y=2\log\left(\left(a-1\right)^{1/2}+\left(a+1\right)^{1/2}\right)+n^{-1/2}\xi.$ (A.72)
We can check this claim in the following way: setting
$f\left(a\right)=2\log\left(\left(a-1\right)^{1/2}+\left(a+1\right)^{1/2}\right),$
we obtain by a computation
$f^{\prime}\left(a\right)$ $=\frac{1}{\left(a-1\right)^{1/2}\left(a+1\right)^{1/2}},$
$\left(f^{\prime}\left(a\right)\right)^{2}$ $=\frac{1}{a^{2}-1}=J\left(a\right).$
This means that at $n=1$ the geometric law $\mathrm{Geo}\left[\left(a-1\right)/\left(a+1\right)\right]$ and the Gaussian model $N\left(f\left(a\right),1\right)$ have the same Fisher information, which implies that the model (A.72) is locally asymptotically equivalent to the model of $n$ i.i.d. geometrics.
A.5 Negative binomial distribution
[…]
$\mathrm{NB}\left(r,p\right)\left(k\right)=P\left(X=k\right)=\frac{\Gamma\left(k+r\right)}{k!\Gamma\left(r\right)}(1-p)^{r}p^{k}.$
[…]
$\frac{\Gamma\left(x+r\right)}{x!\Gamma\left(r\right)}\left(1-p\right)^{r}p^{x}=\exp\left(x\log p\right)h_{r}\left(x\right)\left(1-p\right)^{r}$ (A.73)
for $h_{r}\left(x\right)=$ $\Gamma\left(x+r\right)/x!\Gamma\left(r\right)$. This shows that for fixed $r$, is $\mathrm{NB}\left(r,p\right)$ is an exponential family in the parameter $p$ (with natural parameter $\tau=\log p\in\left(-\infty,0\right)$). Expectation and variance are
$EX=\frac{rp}{1-p},\;\mathrm{Var}\left(X\right)=\frac{rp}{\left(1-p\right)^{2}}$
and the characteristic function is
$\phi\left(t\right)=\left(\frac{1-p}{1-p\exp\left(it\right)}\right)^{r}\text{, }t\in\mathbb{R}\text{.}$ (A.74)
[…]
$f_{s,r}\left(x\right)=\frac{x^{r-1}s^{r}}{\Gamma\left(r\right)}\exp\left(-xs\right)\text{, }x\geq 0,$
[…]
$\mathrm{NB}\left(r,p\right)\left(k\right)=\int_{0}^{\infty}\mathrm{Po}\left(\lambda\right)\left(k\right)f_{s,r}\left(\lambda\right)d\lambda\text{ for }s=\left(1-p\right)/p.$ (A.75)
Relation (A.74) implies that $\mathrm{NB}\left(r,p\right)$ is infinitely divisible; equivalently , if $X_{1},\ldots,X_{n}$ are i.i.d. $\mathrm{NB}\left(r,p\right)$ then
$\sum_{j=1}^{n}X_{i}\sim\mathrm{NB}\left(nr,p\right).$ (A.76)
Moreover, if $X_{1},\ldots,X_{n}$ follow a parametric model as i.i.d. $\mathrm{NB}\left(r,p\right)$, $p\in\left(0,1\right)$, then by the exponential family representation (A.73), $\sum_{j=1}^{n}X_{i}$ is a sufficient statistic.
Lemma A.9
(i) Let $a_{1},a_{2}>0$ and $p_{j}=\left(a_{j}-1\right)/\left(a_{j}+1\right)$, $j=1,2$. Then for any $r>0$ we have
$H^{2}\left(\mathrm{NB}\left(r,p_{1}\right),\mathrm{NB}\left(r,p_{2}\right)\right)\leq\frac{r\left(a_{1}-a_{2}\right)^{2}}{\left(a_{1}-1\right)\left(a_{2}-1\right)}.$
(ii) Let $r_{1},r_{2}>0$. Then for any $p\in\left(0,1\right)$ we have
$H^{2}\left(\mathrm{NB}\left(r_{1},p\right),\mathrm{NB}\left(r_{2},p\right)\right)\leq 1-\frac{\Gamma\left(\left(r_{1}+r_{2}\right)/2\right)}{\Gamma^{1/2}\left(r_{1}\right)\Gamma^{1/2}\left(r_{2}\right)}.$
Proof. (i) The mixture (A.75) represents the operation of a stochastic kernel on $\mathrm{Gam}\left(s,r\right)$. Then it suffices to prove, for $s_{j}=\left(1-p_{j}\right)/p_{j}=2/\left(a_{j}-1\right)$, $j=1,2,$ that
$H^{2}\left(\mathrm{Gam}\left(s_{1},r\right),\mathrm{Gam}\left(s_{2},r\right)\right)\leq\frac{r}{\left(a_{1}-1\right)\left(a_{2}-1\right)}\left(a_{1}-a_{2}\right)^{2}$
(cf. [LM08], Problem 1.72). The squared Hellinger distance can be bounded by the Kullback-Leiber relative entropy $K\left(\cdot.\cdot\right)$ (cf. [Tsy09] ):
$H^{2}\left(\mathrm{Gam}\left(s_{1},r\right),\mathrm{Gam}\left(s_{2},r\right)\right)$ $\leq K\left(\mathrm{Gam}\left(s_{1},r\right),\mathrm{Gam}\left(s_{2},r\right)\right)=\int_{0}^{\infty}f_{s_{1},r}\left(x\right)\log\frac{f_{s_{1},r}\left(x\right)}{f_{s_{2},r}\left(x\right)}dx$
$=\int_{0}^{\infty}\frac{x^{r-1}s_{1}^{r}}{\Gamma\left(r\right)}\left[\left(-x\left(s_{1}-s_{2}\right)\right)+\log\left(\frac{s_{1}}{s_{2}}\right)^{r}\right]dx$
$=-\left(s_{1}-s_{2}\right)\int_{0}^{\infty}\frac{x^{r}s_{1}^{r}}{\Gamma\left(r\right)}dx+r\log\frac{s_{1}}{s_{2}}$
$=-\frac{\left(s_{1}-s_{2}\right)\Gamma\left(r+1\right)}{s_{1}\Gamma\left(r\right)}\int_{0}^{\infty}\frac{x^{r}s_{1}^{r+1}}{\Gamma\left(r+1\right)}dx+r\log\frac{s_{1}}{s_{2}}$
$=-r\left(1-\frac{s_{2}}{s_{1}}\right)-r\log\frac{s_{2}}{s_{1}}.$
The well-known inequality
$\log x\geq\frac{x-1}{x}\text{, }x>0$
applied for $x=s_{2}/s_{1}=\left(a_{1}-1\right)/\left(a_{2}-1\right)$ implies
$H^{2}\left(f_{s_{1},r},f_{s_{2},r}\right)$ $\leq-r\left(\left(1-x\right)+\frac{x-1}{x}\right)=r\frac{\left(x-1\right)^{2}}{x}$
$=r\frac{\left(a_{1}-a_{2}\right)^{2}}{\left(a_{1}-1\right)\left(a_{2}-1\right)}.$
(ii) Again it suffices to prove the bound for the respective Gamma laws, i.e. for $s=\left(1-p\right)/p$
$H^{2}\left(\mathrm{Gam}\left(s,r_{1}\right),\mathrm{Gam}\left(s,r_{2}\right)\right)$ $=1-\int_{0}^{\infty}f_{s_{1},r}^{1/2}\left(x\right)f_{s_{2},r}^{1/2}\left(x\right)dx$
$=1-\frac{1}{\Gamma^{1/2}\left(r_{1}\right)\Gamma^{1/2}\left(r_{2}\right)}\int_{0}^{\infty}x^{\left(r_{1}+r_{2}\right)/2-1}s^{\left(r_{1}+r_{2}\right)/2}\exp\left(-xs\right)dx$
$=1-\frac{\Gamma\left(\left(r_{1}+r_{2}\right)/2\right)}{\Gamma^{1/2}\left(r_{1}\right)\Gamma^{1/2}\left(r_{2}\right)}.$
A.6 A covariance formula for Gaussians
Let $\left(X,Y\right)$ have a bivariate normal distribution
$\left(\begin{array}[c]{c}X\\ Y\end{array}\right)\sim N_{2}\left(0,\Sigma\right)\text{ where }\Sigma=\left(\begin{array}[c]{cc}\sigma_{x}^{2}&\sigma_{xy}\\ \sigma_{xy}&\sigma_{y}^{2}\end{array}\right).$
Then
$EX^{2}Y^{2}$ $=2\sigma_{xy}^{2}+\sigma_{x}^{2}\sigma_{y}^{2},$ (A.77)
$\mathrm{Cov}\left(X^{2},Y^{2}\right)$ $=2\sigma_{xy}^{2}.$ (A.78)
Proof. Consider the well-known regression representation
$Y=\beta X+\eta\text{ where }\eta\sim N\left(0,\sigma_{y}^{2}-\frac{\sigma_{xy}^{2}}{\sigma_{x}^{2}}\right)\text{, }\beta=\frac{\sigma_{xy}}{\sigma_{x}^{2}}$
and $\eta$ is independent of $X$. Then
$EX^{2}Y^{2}$ $=EX^{2}\left(\beta X+\eta\right)^{2}=EX^{2}\left(\beta^{2}X^{2}+2\beta X\eta+\eta^{2}\right)$
$=\beta^{2}EX^{4}+EX^{2}E\eta^{2}=\frac{\sigma_{xy}^{2}}{\sigma_{x}^{4}}3\sigma_{x}^{4}+\sigma_{x}^{2}\left(\sigma_{y}^{2}-\frac{\sigma_{xy}^{2}}{\sigma_{x}^{2}}\right)$
$=3\sigma_{xy}^{2}+\sigma_{x}^{2}\sigma_{y}^{2}-\sigma_{xy}^{2}=2\sigma_{xy}^{2}+\sigma_{x}^{2}\sigma_{y}^{2}.$
This proves (A.77). Then (A.78) follows immediately by
$\mathrm{Cov}\left(X^{2},Y^{2}\right)=EX^{2}Y^{2}-EX^{2}EY^{2}=EX^{2}Y^{2}-\sigma_{x}^{2}\sigma_{y}^{2}.$